### v1

# Nonstandard Universes

Abstract
These notes are concerned with the existence and the basic properties of the set-theoretic universes for nonstandard analysis, compiled by a beginner in the subject. It assumes a basic background in first-order logic, though the necessary material is revised in Appendix A. Needless to say, none of the material presented here is original, but has been adapted from the following sources: [Gol98], [Lin88],[CK90][AFHKL09],[LW15],[Vät07], [Lam15].
1 Universes
1.1 Definition of a Universe
We work in a set theory with atoms (individuals). In our universe $\mathbb{U}$, there will be two kinds of entities, namely individuals and sets. Individuals are entities that contain no members — at least, no members that belong to the universe. Thus from $\mathbb{U}$’s point of view, an element $a\in\mathbb{U}$ is an individual if and only if $a\cap\mathbb{U}=\varnothing$, yet $a\neq\varnothing$. Note that the empty set is not regarded as an individual, but as a set. Sets are entities that are sets in the usual sense, but have the property that each of their elements also belong to the universe.
Convention 1.1
We will use lower case letters $a,b,c\dots$ to range over both sets and individuals, and reserve upper case letters $A,B,C,\dots$ to range over sets.
$\Box$
Definition 1.2
(Transitive Set) A set $A\in\mathbb{U}$ is said to be transitive if and only if elements of elements of $A$ are elements of $A$:
$\forall a\in A\;\forall b\in a\;(b\in A).$
(This is vacuously true if $a\in A$ is an individual, as the quantifiers range over members of $\mathbb{U}$.) Equivalently, if $a\in A$ and $a$ is a set, then $a\subseteq A$.
$\Box$
Definition 1.3
  1. 1.
    A universe $\mathbb{U}$ is a set with the following properties:
    1. (a)
      $\mathbb{U}$ is strongly transitive, i.e. for every $A\in\mathbb{U}$ there is a set $B\in\mathbb{U}$ such that $B$ is transitive and $A\subseteq B\subseteq\mathbb{U}$.
    2. (b)
      If $a,b\in\mathbb{U}$, then $\{a,b\}\in\mathbb{U}$.
    3. (c)
      If $A,B\in\mathbb{U}$, then $A\cup B\in\mathbb{U}$.
    4. (d)
      If $A\in\mathbb{U}$, then $\mathcal{P}(A)\in\mathbb{U}$, where $\mathcal{P}(A)$ is the powerset of $A$.
$\Box$
It follows directly from strong transitivity that a universe is transitive with respect to its set-members: If $A\in\mathbb{U}$ is a set, then $A\subseteq\mathbb{U}$. We say that $\mathbb{U}$ is transitive over sets.
A set $T$ is called supertransitive if and only if whenever $A\in T$, then $A\cup\mathcal{P}(A)\subseteq T$. Clearly, a supertransitive set is transitive.
Lemma 1.4
If  $\mathbb{U}$ is a universe, then every $A\in\mathbb{U}$ is an element of some supertransitive set $T\in\mathbb{U}$.
Proof: Suppose that $S\in\mathbb{U}$ is transitive, and define $T:=S\cup\mathcal{P}(S)$. Then $T\in\mathbb{U}$, by the definition of a universe. We claim that $T$ is supertransitive. Indeed if $A\in T$, then either $A\in S$ or $A\in\mathcal{P}(S)$. Either way, we see that $A\subseteq S$, and hence $A\subseteq T$. Now if $B\in\mathcal{P}(A)$, then $B\subseteq A\subseteq S$, so $B\in\mathcal{P}(S)\subseteq T$. Hence $\mathcal{P}(A)\subseteq T$. It follows that if $A\in T$, then $A\cup\mathcal{P}(A)\subseteq T$, so that $T$ is supertransitive.
Now suppose $A\in\mathbb{U}$. As $\mathbb{U}$ is strongly transitive, there is a transitive set $S\in\mathbb{U}$ such that $A\subseteq S$. Let $T=S\cup\mathcal{P}(S)$. Then $T$ is supertransitive and $A\in T$.
$\dashv$
Definition 1.5
(Universe over $X$) If $X$ is a set, then $\mathbb{U}$ is said to be a universe over $X$ if and only if $X\in\mathbb{U}$, $\varnothing\not\in X$, and no element of a member of $X$ belongs to $\mathbb{U}$, i.e. $\bigcup X\cap\mathbb{U}=\varnothing$. Thus, from the point of view of $\mathbb{U}$, no member of $X$ has elements, yet none are the empty set — they are individuals.
$\Box$
Note that if $\mathbb{U}$ is a universe over $X$, and $Y\subseteq X$, then $\mathbb{U}$ is also a universe over $Y$. Similarly, if $\mathbb{U}$ is a universe over $X$ for each $X\in\mathcal{X}\in\mathbb{U}$, then $\mathbb{U}$ is a universe over $\bigcup\mathcal{X}$.
Example 1.6
Superstructures are the most common universes in practice.
Suppose that $X$ is a set. The superstructure over $X$, denoted $V(X)$ is defined inductively as follows:
$V_{0}(X)$ $:=X$
$V_{n+1}(X)$ $:=V_{n}(X)\cup\mathcal{P}(V_{n}(X))$
$V(X)$ $:=\bigcup_{n<\omega}V_{n}(X)$
Clearly $V_{n}(X)\subseteq V_{n+1}(X)$ for all $n<\omega$, and $a\in V_{n+1}(X)$ if and only if $a\in V_{n}(X)$ or $a\subseteq V_{n}(X)$.
It is easy to show by induction that $V_{n+1}(X)=X\cup\mathcal{P}(V_{n}(X))$: This is obvious in the case $n=0$. Next, suppose that $V_{n}(X)=X\cup\mathcal{P}(V_{n-1}(X))$. If $a\in V_{n+1}(X)$, then either (i) $a\in V_{n}(X)$ or (ii) $a\subseteq V_{n}(X)$ (or both). Hence either (i) $a\in X$ or $a\subseteq V_{n-1}(X)$, or (ii) $a\subseteq V_{n}(X)$. Since $V_{n-1}(X)\subseteq V_{n}(X)$ it follows that either (i) $a\in X$, or (ii) $a\subseteq V_{n}(X)$ (or both). It follows that $V_{n+1}(X)\subseteq X\cup\mathcal{P}(V_{n}(X))$. The reverse inclusion is obvious.
In order for $V(X)$ to be a universe over $X$, one requirement is that the members of $X$ act like individuals, i.e. that $x\cap V(X)=\varnothing$ for all $x\in X$. The set $X$ is said to be a base set if $\varnothing\not\in X$, and $\forall x\in X(x\cap V(X)=\varnothing)$. Note that it is always possible to replace a set $X$ by a base set of the same size: For example, given an infinite ordinal $\alpha$, pick a set $Y$ with the same cardinality as $X$ such that every element of an element of $Y$ has rank $\alpha$. One can always choose $\alpha$ sufficiently large so that this is possible. It is then easy to see by induction that each element of $V_{n}(Y)$ has a rank $\beta$ where either $\beta<n$ or $\alpha<\beta\leq\alpha+n+1$. Now if $z\in y\in Y$, then $\text{rank}(z)=\alpha$, so $z\not\in V(Y)$, i.e. $y\cap V(Y)=\varnothing$.
Now assume that $X$ is a base set. It is easy to see that each by induction $V_{n}(X)$ is transitive over sets: Certainly if $A\in V_{n+1}(X)=X\cap\mathcal{P}(V_{n}(X))$, then $A\subseteq V_{n}(X)$, since $A$ is a set, i.e. $A\not\in X$. It follows that $A\subseteq V_{n+1}(X)$. In particular, it follows that $V(X)$ is strongly transitive.
It is also easy to see that if $a,b\in V_{n}(X)$, then $\{a,b\}\in V_{n+1}(X)$. Next, if $A,B\in V_{n}(X)$, then $A\cup B\subseteq V_{n-1}(X)$, so $A\cup B\in V_{n}(X)$. Further, if $A\in V_{n}(X)$, then $A\subseteq V_{n-1}(X)$, so $\mathcal{P}(A)\subseteq\mathcal{P}(V_{n-1}(X))\subseteq V_{n}(X)$, and hence $\mathcal{P}(A)\in V_{n+1}(X)$.
Hence if $X$ is a base set, then $V(X)$ is a universe over $X$.
Observe that the sets $V_{n}(X)$ that make up a superstructure $V(X)$ over $X$ are supertransitive.
$\Box$
Remarks 1.7
Suppose that $\mathbb{U}$ is a universe over $X$. In that case it easy to prove by induction on $n$ that each $V_{n}(X)\subseteq\mathbb{U}$, from which it follows that $V(X)\subseteq\mathbb{U}$. Thus $V(X)$ is the smallest universe over $X$, assuming that one exists.
Not every universe is of the form $V(X)$, however. For example, consider $\mathbb{U}:=V_{\omega+\omega}$ in the usual cumulative hierarchy of sets. This is a universe over $\varnothing$, and there are no individuals, since if $y\in x\in\mathbb{U}$, then $y\in\mathbb{U}$. Yet $\mathbb{U}\neq V(\varnothing)$, since $V(\varnothing)=V_{\omega}$.
$\Box$
1.2 Closure Properties of a Universe
Observe the following closure properties of a universe, which are easy consequences of the definition of universe. (Recall again our convention concerning upper and lower case letters.)
  1. 1.
    $a\in\mathbb{U}$ implies $\{a\}\in\mathbb{U}$.
  2. 2.
    $A_{1},\dots,A_{m}\in\mathbb{U}$ implies $A_{1}\cup\dots\cup A_{m}\in\mathbb{U}$.
  3. 3.
    If $A\subseteq\mathbb{U}$ is finite, then $A\in\mathbb{U}$
    (By 1., 2.)
  4. 4.
    If $A\subseteq B$ and $B\in\mathbb{U}$, then $A\in\mathbb{U}$.
    (For $A\in\mathcal{P}(B)\in\mathbb{U}$, and $\mathbb{U}$ is transitive.)
  5. 5.
    If $\{A_{i}:i\in I\}\subseteq B\in\mathbb{U}$, then $\bigcup_{i\in I}A_{i}\in\mathbb{U}$.
    (For there is transitive $T\in\mathbb{U}$ such that $B\subseteq T$. Then each $A_{i}\in T$, so each $A_{i}\subseteq T$, and hence $\bigcup_{i\in I}A_{i}\subseteq T$. Now apply 4.)
  6. 6.
    If $B=\{A_{i}:i\in I\}\in\mathbb{U}$, then $\bigcup B=\bigcup_{i\in I}A_{i}\in\mathbb{U}$.
    (Follows directly from 5.)
  7. 7.
    If $\{A_{i}:i\in I\}\subseteq\mathbb{U}$ is a non-empty family of sets, then $\bigcap_{i\in I}A_{i}\in\mathbb{U}$.
    (The intersection is a subset of $A_{i_{0}}$, for any $i_{0}\in I$. Now apply 4.)
  8. 8.
    $a,b\in\mathbb{U}$ implies $(a,b)\in\mathbb{U}$. More generally, if $a_{1},\dots,a_{m}\in\mathbb{U}$, then $(a_{1},\dots,a_{m})\in\mathbb{U}$.
    (Because $(a,b):=\{\{a\},\{a,b\}\}$. Then $(a,b,c):=((a,b),c)$, etc.)
  9. 9.
    If $A,B\in\mathbb{U}$ and $R\subseteq A\times B$, then $R\in\mathbb{U}$. More generally, if $A_{1},\dots,A_{m}\in\mathbb{U}$ and $R\subseteq A_{1}\times\dots\times A_{m}$, then $R\in\mathbb{U}$.
    (Since $R\subseteq\mathcal{P}\mathcal{P}(A\cup B)$, the result follows from 2. and 4.)
  10. 10.
    If $R\in\mathbb{U}$ is a binary relation, then $\text{\rm dom}(R),\text{\rm ran}(R),R^{-1}\in\mathbb{U}$. Furthermore, if $C\subseteq\text{\rm dom}(R)$, then $R[C]\in\mathbb{U}$.
    (For $\text{\rm dom}(R),\text{\rm ran}(R)\subseteq\bigcup\bigcup R\in\mathbb{U}$, so by 4. we have that $\text{\rm dom}(R),\text{\rm ran}(R)\in\mathbb{U}$. Also $R^{-1}\subseteq\text{\rm ran}(R)\times\text{\rm dom}(R)$, so that $R^{-1}\in\mathbb{U}$ follows by 9. Finally $R[C]\subseteq\text{\rm ran}(R)\in\mathbb{U}$.)
  11. 11.
    If $A,B\in\mathbb{U}$ and $f:A\to B$, then $f\in\mathbb{U}$. Furthermore, if $A^{\prime}\subseteq A,B^{\prime}\subseteq B$, then $f[A^{\prime}],f^{-1}[B^{\prime}]\in\mathbb{U}$.
    (First, $f\subseteq A\times B\in\mathbb{U}$. Then $f[A^{\prime}]\subseteq B\in\mathbb{U}$ and $f^{-1}[B^{\prime}]\subseteq A\in\mathbb{U}$.)
  12. 12.
    If $A,B\in\mathbb{U}$, then $B^{A}\in\mathbb{U}$, where $B^{A}$ is the set of all functions from $A$ to $B$..
    (For $B^{A}\subseteq\mathcal{P}(A\times B)$.)
  13. 13.
    If $\{A_{i}:i\in\mathbb{I}\}\in\mathbb{U}$,and $I\in\mathbb{U}$, then $\prod_{i\in I}A_{i}\in\mathbb{U}$.
    (Because $\prod_{I}A_{i}\subseteq\;(\bigcup_{I}A_{i})^{I}$.)
1.3 Nonstandard Embeddings
We assume some familiarity with basic first-order logic, including the basics of model theory. Refer to Appendix A for a quick reminder of the basic notions used below.
Let $\mathcal{L}_{\in}$ denote a first-order language with equality $=$ and a single binary relation symbol $\in$. We assume that there is a countable collection of variables, and take as basic propositional connectives the connectives $\lnot$ (not) and $\land$ (and), and as basic quantifier $\forall$ (for all). The other connectives $\lor$ (or), $\to$ (then), $\leftrightarrow$ (if and only if) are defined in terms of ,$\land$,$\lnot$ , ¬ in the usual way, and the existential quantifier is defined in terms of ¬$\lnot$,$\forall$ , ∀ in the usual way.
In nonstandard analysis, the following types of $\mathcal{L}_{\in}$-formula play a central role:
Definition 1.8
(Bounded Formula) A bounded $\mathcal{L}_{\in}$-formula is an $\mathcal{L}_{\in}$-formula all of whose quantifiers are bounded, i.e. of the form $\forall x\in y$ or $\exists x\in y$, where $\forall x\in y\;\varphi(x,y)$ is an abbreviation of $\forall x\;(x\in y\to\varphi(x,y))$, and $\exists x\in y\;\varphi(x,y)$ abbreviates $\exists x\;(x\in y\land\varphi(x,y))$.
$\Box$
Definition 1.9
(Transfer Map)
  1. (a)
    A transfer map for a set $X$ is a function ${}^{*}:\mathbb{U}\to\mathbb{V}$ between two universes $\mathbb{U},\mathbb{V}$ with the properties that:
    1. 1.
      $\mathbb{U}$ is a universe over $X$.
    2. 2.
      ${}^{*}x=x$ for every $x\in X$, and ${}^{*}\varnothing=\varnothing$
    3. 3.
      Transfer: * is a bounded elementary embedding, i.e. if $\varphi(x_{1},\dots,x_{n})$ is an $\mathcal{L}_{\in}$-formula, and $u_{1},\dots,u_{n}\in\mathbb{U}$, then
      $\mathbb{U}\vDash\varphi[u_{1},\dots,u_{n}]\quad\text{iff}\quad\mathbb{V}\vDash\varphi[{}^{*}u_{1},\dots,{}^{*}u_{n}].$
    If $\mathbb{U}=V(X),\mathbb{V}=V(Y)$ are superstructures over $X,Y$ respectively, then it is usually also required that ${}^{*}X=Y$.
  2. (b)
    A transfer map $*:\mathbb{U}\to\mathbb{V}$ for $X$ is a nonstandard framework, or a nonstandard embedding, if there is a countable $C\in\mathbb{U}$ such that ${}^{\sigma}C$ is a proper subset of ${}^{*}C$, i.e. ${}^{\sigma}C\subsetneqq{}^{*}C$.
$\Box$
The transfer property in the preceding definition states that the *-map transfers properties that are definable by bounded formulas from $\mathbb{U}$ to $\mathbb{V}$.
Notation: For $A\in\mathbb{U}$, define:
${}^{\sigma}A:=*[A]:=\{{}^{*}a:a\in A\}$
to be the image of the set $A$ under the *-map.
Note that if * is a transfer map, then always ${}^{\sigma}A\subseteq{}^{*}A$, as $V(X)\vDash a\in A$ implies $V(Y)\vDash{}^{*}a\in{}^{*}A$. If $A$ is finite, then we shall see that ${}^{\sigma}A={}^{*}A$. If, however, the *-map is a nonstandard framework — so that $C$ is a proper subset of ${}^{*}C$ for some countable $C\in\mathbb{U}$— then it will transpire that ${}^{\sigma}A\subsetneqq{}^{*}A$ whenever $A\in\mathbb{U}$ is infinite. In that case, we can think of ${}^{*}A$ as a version of $A\in\mathbb{U}$ that lives in $\mathbb{V}$ — in that $A$ and ${}^{*}A$ satisfy the same bounded sentences — but where ${}^{*}A$ has additional elements that do not correspond to members of $A$.
Lemma 1.10
Suppose that $*:\mathbb{U}\to\mathbb{V}$ is a transfer map for $X$. Then $\mathbb{V}$ is a universe over ${}^{*}X$.
Proof: We need only show that $\varnothing\not\in{}^{*}X$ and that $\bigcup{}^{*}X\cap\mathbb{V}=\varnothing$. Since ${}^{*}\varnothing=\varnothing$, we see that $\varnothing\not\in X$ transfers to $\varnothing\not\in{}^{*}X$. Next, if $\bigcup{}^{*}X\cap\mathbb{V}\neq\varnothing$, then $\mathbb{V}\vDash\exists x\in{}^{*}X\;\exists y\in x\;(y=y)$ (which simply says that there is an $x\in{}^{*}X$ which has an element $y\in\mathbb{V}$). Then transfer implies that $\bigcup X\cap\mathbb{U}\neq\varnothing$, which contradicts the fact that $\mathbb{U}$ is a universe over $X$.
$\dashv$
Convention 1.11
The language $\mathcal{L}_{\in}$ in which we work has no constant symbols. However, in the interests of brevity we will often write formulas as if there are constant symbols for every member $a\in\mathbb{U}$. For example, when we write
$\mathbb{U}\vDash\exists x\in A\;\forall y\in B\;(c\in y\land\psi(x,y,d)),\qquad(\text{where }A,B,c,d\in\mathbb{U})$
this should be taken to mean
$\mathbb{U}\vDash\varphi[A,B,c,d],\quad\text{where }\varphi(u,v,w,t)\equiv\exists x\in u\;\exists y\in v\;(z\in y\land\psi(x,y,t)).$
By transfer, we then have
$\mathbb{V}\vDash\varphi[{}^{*}A,{}^{*}B,{}^{*}c,{}^{*}d]\quad\text{i.e.}\quad\mathbb{V}\vDash\exists x\in{}^{*}A\;\forall y\in{}^{*}B\;({}^{*}c\in y\land\psi(x,y,{}^{*}d)).$
Thus without loss of generality, we may assume that, when working with a transfer map $*:\mathbb{U}\to\mathbb{V}$, the language $\mathcal{L}_{\in}$ is expanded to a language — denoted $\mathcal{L}_{\mathbb{U}}$ — which has a constant symbol $c_{a}$ for every entity $a\in\mathbb{U}$. Naturally, the constant symbol $c_{a}$ is to be interpreted as the entity $a$ in the model $\mathbb{U}$. If $*:\mathbb{U}\to\mathbb{V}$ is a transfer map, then $c_{a}$ will be interpreted as ${}^{*}a$ in $\mathbb{V}$. When we replace all occurrences of these constants in a formula $\varphi$ by their *-value, we obtain the *-transform ${}^{*}\varphi$ of the formula. Thus, for example
${}^{*}\Big{(}\exists x\in A\;\forall y\in B\;(c\in y\land\psi(x,y,d))\Big{)}\equiv\exists x\in{}^{*}A\;\forall y\in{}^{*}B\;({}^{*}c\in y\land\psi(x,y,{}^{*}d)).$
It is not hard to see how to define ${}^{*}\varphi$ for formulas $\varphi$ by induction on the complexity of $\varphi$.
The transfer property is then easily seen to be equivalent to the following: If $\varphi$ is a bounded sentence of $\mathcal{L}_{\mathbb{U}}$, then $\mathbb{U}\vDash\varphi$ if and only if $\mathbb{V}\vDash{}^{*}\varphi$.
$\Box$
In the next section, we will discuss some of the properties of nonstandard frameworks, should they exist. The question of existence is dealt with in Section 6.1, but, assuming familiarity with the model-theoretic concepts in Appendix A, this can be read now. We’ll have more to say about part (b) of Definition 1.9 later on. For now, note that this property is essential for nonstandard analysis to have any real power via the introduction of nonstandard objects.
2 Properties of the *-map
In this section, we assume that $*:\mathbb{U}\to\mathbb{V}$ is a transfer map between two universes $\mathbb{U},\mathbb{V}$.
Note that the *-map is injective: For if $\mathbb{V}\vDash{}^{*}a={}^{*}b$, then $\mathbb{U}\vDash a=b$, by the transfer property.
Further observe that * maps individuals in $\mathbb{U}$ to individuals in $\mathbb{V}$: For suppose that $a\in\mathbb{U}$ is an individual. Then $\mathbb{U}\vDash\lnot\exists x\in a\;(x=x)$, i.e. $a$ has no elements (in common with $\mathbb{U}$). By transfer, ${}^{*}a$ has no elements (in common with $\mathbb{V}$). Since is injective, ${}^{*}a\neq{}^{*}\varnothing$. Now ${}^{*}\varnothing=\varnothing$, by definition of the *-map. Thus ${}^{*}a\in\mathbb{V}$ is an element which has no members (in common with $\mathbb{V}$), yet is not the empty set, i.e. ${}^{*}a$ is an individual in $\mathbb{V}$.
The following lemma shows that certain basic operations can be expressed by bounded $\mathcal{L}_{\in}$-formulas $\varphi_{n}$. In what follows, we will be able to improve readability by using these abbreviations instead of the $\varphi_{n}$ inside bounded $\mathcal{L}_{\in}$-formulas.
Lemma 2.1
Let $\mathbb{U}$ be a universe. There are bounded formulas $\varphi_{0},\dots,\varphi_{7}$ such that for all elements $a_{n}\in\mathbb{U}$ and all sets $A_{n}\in\mathbb{U}$, the following hold:
  1. (a)
    $A_{1}=\varnothing$ if and only if $\mathbb{U}\vDash\varphi_{0}[A_{1}]$.
  2. (b)
    $A_{1}=\{a_{1},\dots,a_{n}\}$ if and only if $\mathbb{U}\vDash\varphi_{1,n}[A_{1},a_{1},\dots,a_{n}]$.
  3. (c)
    $A_{1}=(a_{1},\dots,a_{n})$ if and only if $\mathbb{U}\vDash\varphi_{2,n}[A_{1},a_{1},\dots,a_{n}]$.
  4. (d)
    $A_{1}\subseteq A_{2}$ if and only if $\mathbb{U}\vDash\varphi_{3}[A_{1},A_{2}]$.
  5. (e)
    $A_{1}=A_{2}\times A_{3}$ if and only if $\mathbb{U}\vDash\varphi_{4}[A_{1},A_{2},A_{3}]$.
  6. (f)
    $A_{1}:A_{2}\to A_{3}$ if and only if $\mathbb{U}\vDash\varphi_{5}[A_{1},A_{2},A_{3}]$.
  7. (g)
    If $\mathbb{U}$ is a universe over $X$, so that $V(X)\subseteq\mathbb{U}$, there is $\varphi_{6,n}$ such that $a_{1}\in V_{n}(X)$ if and only if $\mathbb{U}\vDash\varphi_{6,n}[X,a_{1}]$.
  8. (h)
    If $\mathbb{U}$ is a universe over $X$, so that $V(X)\subseteq\mathbb{U}$, there is $\varphi_{7,n}$ such that $A_{1}$ is a set in $V_{n}(X)$ if and only if $\mathbb{U}\vDash\varphi_{7,n}[X,A_{1}]$.
Proof: (a) Take $\varphi_{0}(x)\equiv\forall y\in x\;(y\neq y)$. (Recall that $A_{1}$ is required to be a set.)
(b) Take $\varphi_{1,n}(x,y_{1},\dots,y_{n})$ to be the formula
$y_{1}\in x\land y_{2}\in x\land\dots\land y_{n}\in x\land\forall y\in x\;(y=y_{1}\lor y=y_{2}\lor\dots\lor y=y_{n}).$
(c) For $n=2$, we see that $A_{1}=(a_{1},a_{2})=\{\{a_{1}\},\{a_{1},a_{2}\}\}$ if and only if
$\mathbb{U}\vDash\exists x\in A_{1}\;\exists y\in A_{1}\;(A_{1}=\{x,y\}\land x=\{a_{1}\}\land y=\{a_{1},a_{2}\}).$
The formulas inside the brackets are abbreviations of $\varphi_{1}$. This defines a bounded formula $\varphi_{2,2}$ for the case $n=2$. Then for $n=3$, we can proceed in a similar way to define $\varphi_{2,3}$, using the just defined $\varphi_{2,2}$, and the fact that $(a_{1},a_{2},a_{3}):=((a_{1},a_{2}),a_{3})$, etc.
(d) Take $\forall u\in x\;(u\in y)$ for $\varphi_{3}(x,y)$.
(e) Note that $A_{1}=A_{2}\times A_{3}$ if and only if
$\forall u\in A_{1}\;\exists v\in A_{2}\;\exists w\in A_{3}\;(u=(v,w))\land\forall v\in A_{2}\;\forall w\in A_{3}\;\exists u\in A_{1}\;(u=(v,w)),$
where statements of the form $x=(y,z)$ are to be replaced by versions of $\varphi_{2}(x,y,z)$. From here, $\varphi_{4}(x,y,z)$ is apparent.
(f) Note that $A_{1}$ is a function from $A_{2}$ to $A_{3}$ if and only if
$A_{1}\subseteq A_{2}\times A_{3}\land\forall v\in A_{2}\;\exists!w\in A_{3}\;\exists u\in A_{1}\;(u=(v,w)),$
where $\exists!y\in x\;\psi(x,y)$ abbreviates the bounded formula $\exists y\in x\;(\psi(x,y)\land\forall z\in x\;(\psi(x,z)\to z=y))$, which states that $y$ is the unique member of $x$ for which $\psi(x,y)$ holds. The required formula $\varphi_{5}$ can now be constructed easily with the aid of $\varphi_{3},\varphi_{4}$.
(g) This is proved by induction on $n$. For the case $n=0$, note that $V_{0}(X)=X$, so take $\varphi_{6,0}(x,y)\equiv y\in x$. Then if $a_{1}\in V_{1}(X)=X\cup\mathcal{P}(X)$, necessarily $a_{1}\in X$ or $a_{1}\subseteq X$, so take $\varphi_{6,1}(x,y)\equiv y\in x\lor\forall z\in y\;(z\in x)$. Similarly, note that if $a_{1}\in V_{n+1}(X)=X\cup\mathcal{P}(V_{n}(X))$, then $a_{1}\in X$ or $a_{1}\subseteq V_{n}(X)$, so define $\varphi_{6,n+1}(x,y)\equiv y\in x\lor\forall z\in y\;\varphi_{6,n}(x,z)$.
(h) Take $\varphi_{7,n}(x,y)$ to be $\varphi_{6,n}(x,y)\land y\not\in x$. Then $\varphi_{7,n}(X,A_{1})$ holds if and only if $A_{1}$ is a set in $V_{n}(X)$.
$\dashv$
Theorem 2.2
(*-Comprehension) Let $\varphi(y,x_{1},\dots,x_{n})$ be a bounded $\mathcal{L}_{\in}$-formula. For all $u_{1},\dots,u_{n},a\in\mathbb{U}$,
{${}^{*}\{y\in a:\mathbb{U}\vDash\varphi(y,u_{1},\dots,u_{n})\}=\{y\in{}^{*}a:\mathbb{V}\vDash\varphi(y,{}^{*}u_{1},\dots,{}^{*}u_{n})\}.$
Proof: By definition of the *-map, ${}^{*}\varnothing=\varnothing$, so the result is true if $a\in\mathbb{U}$ is an individual. Assume therefore that $a$ is a set in $\mathbb{U}$. Define
$B:=\{y\in a:\mathbb{U}\vDash\varphi(y,u_{1},\dots,u_{n})\}.$
Then $B\in\mathcal{P}(a)\subseteq\mathbb{U}$, and
$\mathbb{U}\vDash\forall y\in a\;(y\in B\leftrightarrow\varphi(y,u_{1},\dots,u_{n})).$
By transfer,
$\mathbb{V}\vDash\forall y\in{}^{*}a\;(y\in{}^{*}B\leftrightarrow\varphi(y,{}^{*}u_{1},\dots,{}^{*}u_{n})),$
from which it follows that ${}^{*}B\cap{}^{*}a=\{y\in{}^{*}a:\varphi(y,{}^{*}u_{1},\dots,{}^{*}u_{n})\}$. But $B\subseteq a$, so ${}^{*}B\subseteq{}^{*}a$, again by transfer, since the formula $B\subseteq a$ is clearly equivalent to a bounded formula of $\mathcal{L}_{\in}$. Thus the result follows.
$\dashv$
Observe that if $*:\mathbb{U}\to\mathbb{V}$ is a transfer map for a set $X$, then:
  • $a\in B$ if and only if ${}^{*}a\in{}^{*}B$ and $a=b$ if and only if ${}^{*}a={}^{*}b$.
  • $A\subseteq B$ if and only if ${}^{*}A\subseteq{}^{*}B$.
  • If $A\subseteq X$, then $A\subseteq{}^{*}A\subseteq{}^{*}X$. In particular, $X\subseteq{}^{*}X$: This is because ${}^{*}x=x$ for $x\in X$.
  • ${}^{*}(A\cup B)={}^{*}A\cup{}^{*}B$, ${}^{*}(A\cap B)={}^{*}A\cap{}^{*}B$, ${}^{*}(A-B)={}^{*}A-{}^{*}B$: for example, by the preceding Lemma, ${}^{*}(A\cup B)={}^{*}\{x\in A\cup B:x\in A\lor x\in B\}=\{x\in{}^{*}(A\cup B):x\in{}^{*}A\lor x\in{}^{*}B\}={}^{*}(A\cup B)\cap({}^{*}A\cup{}^{*}B)$. Since ${}^{*}A,{}^{*}B\subseteq{}^{*}(A\cup B)$, it follows that ${}^{*}(A\cup B)={}^{*}A\cup{}^{*}B$.
  • If $A=\{A_{1},\dots,a_{n}\}$ is a finite set, then ${}^{*}A={}^{\sigma}A=\{{}^{*}a_{1},\dots,{}^{*}a_{n}\}$: This is because $A=\{x\in A:x=a_{1}\lor\dots\lor x=a_{n}\}$.
  • If $A$ is a transitive set, then ${}^{*}A$ is a transitive set: For $A$ is transitive if and only if the bounded formula $\forall x\in A\;\forall y\in x\;(y\in A)$ holds in $\mathbb{U}$.
  • ${}^{*}\mathcal{P}(A)\subseteq\mathcal{P}({}^{*}A)$: This follows by transfer of the formula $\forall X\in\mathcal{P}(A)\;(X\subseteq A)$. Hence every element $X$ of ${}^{*}\mathcal{P}(A)$ is a subset of ${}^{*}A$ and thus in $\mathcal{P}({}^{*}A)$.
  • ${}^{*}(a_{1},\dots,a_{n})=({}^{*}a_{1},\dots,{}^{*}a_{n})$: This follows from Lemma 2.1(c).
  • ${}^{*}(A_{1}\times\dots\times A_{n})={}^{*}A_{1}\times\dots\times{}^{*}A_{n}$: This follows from Lemma 2.1(e).
  • If $R$ is an $n$-ary relation, so is ${}^{*}R$: For then $R\subseteq A_{1}\times\dots\times A_{n}$ for some sets $A_{1},\dots,A_{n}\in\mathbb{U}$.
  • If $R$ is a binary relation, then
    • ${}^{*}\text{\rm dom}(R)=\text{\rm dom}({}^{*}R)$: For if $x\in\text{\rm dom}(R)$, then $(x,y)=\{\{x\},\{x,y\}\}$ belongs to $R$ for some $y$. Hence, with $a:=\{\{x\},\{x,y\}\}$, $b:=\{x\},c:=\{x,y\}$, we have $x\in\text{\rm dom}(R)$ iff
      $\exists a\in R\;\exists b\in a\;\exists c\in a\;(a=\{b,c\}\land b=\{x\}\land\exists y\in c\;(c=\{x,y\})).$
      The result now follows by Lemma 2.1(b). (Or, write it all out:
      $\exists a\in R\;\exists b\in a\;\exists c\in a$ $(\forall d\in a\;(d=b\lor d=c)\land x\in b\land\forall w\in b\;(w=x)$
      $\land x\in c\land\exists y\in c\;\forall w\in c(w=x\lor w=y)),$
      a bounded formula.)
    • ${}^{*}\text{\rm ran}(R)=\text{\rm ran}(^{*}R)$: The proof is very similar to the preceding.
    • ${}^{*}(R^{-1})=({}^{*}R)^{-1}$: Choose $T\in\mathbb{U}$ transitive so that $R\cup R^{-1}\subseteq T$. Then if $(x,y)\in R$, it follows that $x,y\in T$. Then
      $\mathbb{U}\vDash\forall x\in T\;\forall y\in T\;((x,y)\in R^{-1}\leftrightarrow(y,x)\in R),$
      and hence
      $\mathbb{V}\vDash\forall x\in{}^{*}T\;\forall y\in{}^{*}T\;((x,y)\in{}^{*}(R^{-1})\leftrightarrow(y,x)\in{}^{*}R),$
      using Lemma 2.1(c). As ${}^{*}R\cup{}^{*}(R^{-1})\subseteq{}^{*}T$, it follows that ${}^{*}(R^{-1})=({}^{*}R)^{-1}$.
    • If $C\subseteq\text{\rm dom}(R)$, then ${}^{*}(R[C])={}^{*}R[{}^{*}C]$: Consider the bounded formula
      $\forall y\in\text{\rm ran}(R)\;(y\in R[C]\leftrightarrow\exists x\in C\;((x,y)\in R)$. if and only if
    • If $D\subseteq\text{\rm ran}(R)$, then ${}^{*}(R^{-1}[D])=({}^{*}(R^{-1}[{}^{*}D])$: For $\text{\rm ran}(R)=\text{\rm dom}(R^{-1})$.
    • If $R,S$ are binary relations, then ${}^{*}(R\circ S)={}^{*}R\circ{}^{*}S$: Choose a transitive $T\in\mathbb{U}$ such that $R\cup S\subseteq T$. Then use transfer on the bounded formula
      $\forall x\in T\;\forall z\in T\;((x,z)\in R\circ S\leftrightarrow\exists y\in T\;((x,y)\in S\land(y,z)\in R)).$
  • If $f:A\to B$ is a function in $\mathbb{U}$, then ${}^{*}f:{}^{*}A\to{}^{*}B$ is a function in $\mathbb{V}$. Moreover, ${}^{*}(f(a))={}^{*}f({}^{*}a)$ for $a\in A$, and ${}^{*}f$ is injective/surjective if and only if $f$ is injective/surjective: Since $f$ is a binary relation with $\text{\rm dom}(f)=A,\text{\rm ran}(f)\subseteq B$, we immediately see that ${}^{*}f$ is a binary relation with $\text{\rm dom}({}^{*}f)={}^{*}A,\text{\rm ran}({}^{*}f)\subseteq{}^{*}B$. Since $\forall a\in A\exists!b\in B\;((a,b)\in f)$, it follows by transfer that ${}^{*}f$ is a function. Furthermore, transfer of $\forall a\in A\forall b\in B\;((a,b)\in f\leftrightarrow b=f(a))$ now leads to $({}^{*}f)({}^{*}a)={}^{*}f(a)$. Next, $f$ is injective if and only if the bounded formula $\forall a_{1}\in A\;\forall a_{2}\in A\;(\exists b\in B\;((a_{1},b)\in f\land(a_{2},b)\in f)\rightarrow a_{1}=a_{2})$ is satisfied, from which it follows by transfer ${}^{*}f$ is injective if and only if $f$ is. Finally $f$ is surjective if and only if $\forall b\in B\;\exists a\in A\;((a,b)\in f)$.
Lemma 2.3
If $*:\mathbb{U}\to\mathbb{V}$ is a transfer map for a set $X$ and $V_{n}(X)\subseteq B\in\mathbb{U}$, then ${}^{*}V_{n}(X)=V_{n}({}^{*}X)\cap{}^{*}B$.
Proof: By Lemma 2.1(g), we have
$V_{n}(X)=V_{n}(X)\cap B=\{x\in B:\mathbb{U}\vDash\varphi_{6,n}(X,x)\},$
and thus
${}^{*}V_{n}(X)=\{x\in{}^{*}B:\mathbb{V}\vDash\varphi_{6,n}({}^{*}X,x)\}=V_{n}({}^{*}X)\cap{}^{*}B.$
$\dashv$
The following result is crucial, as we wish to be able to talk about structures in our universes. For example, suppose that $*:V(\mathbb{R})\to V({}^{*}\mathbb{R})$ is a transfer map over the set $\mathbb{R}$ between two superstructures. When talking about the reals, we wish to take into account the operations and relations +,$+,\cdot,-,{}^{-1},\leq$ as well, so that we want to be able to talk about the model (,+$(\mathbb{R},+,\cdot,-,{}^{-1},\leq)$. This will transfer to a model (,+$({}^{*}\mathbb{R},{}^{*}+,{}^{*}\cdot,{}^{*}-,{}^{*}{}^{-1},{}^{*}\leq)$. The following result shows that these two models are elementarily equivalent, i.e. that they satisfy the same first-order sentences. Indeed, the *-map induces an elementary embedding from the one into the other:
Theorem 2.4
[…]
Proof: To simplify notation, suppose that $\mathfrak{A}=(A,R^{\mathfrak{A}})$ where $R\in\mathcal{L}$ is a single binary relation symbol. Then $R^{\mathfrak{A}}\subseteq A\times A$, and thus $R^{\mathfrak{A}}\in\mathbb{U}$. Hence $\mathfrak{A}\in\mathbb{U}$, and ${}^{*}\mathfrak{A}=({}^{*}A,{}^{*}R^{\mathfrak{A}})$. Choose $T\in\mathbb{U}$ transitive so that $\mathfrak{A}\in T$. We will show by induction on complexity that for every $\mathcal{L}$-formula $\varphi(x_{1},\dots,x_{n})$ there is a bounded $\mathcal{L}_{\in}$-formula $\bar{\varphi}$ such that
$\mathfrak{A}\vDash\varphi[a_{1},\dots,a_{n}]\qquad\text{if and only if}\qquad\mathbb{U}\vDash\bar{\varphi}[a_{1},\dots,a_{n},\mathfrak{A},T]\qquad\text{(for $a_{1},\dots,a_{n}\in A$)}.$
Suppose first that $\varphi(x,y)\equiv R(x,y)$ is atomic. Define the bounded $\mathcal{L}_{\in}$ formula $\bar{\varphi}$ by
$\bar{\varphi}(x,y,z,t)\equiv\exists u\in t\;\exists v\in t\;\exists w\in t\;\Big{(}w=(x,y)\land z=(u,v)\land w\in v\Big{)}.$
Then $\bar{\varphi}[a_{1},a_{2},\mathfrak{A},T]$ asserts that there are $u,v,w\in T$ such that $w=(a_{1},a_{2}),z=(A,R^{\mathfrak{A}})$ and $w\in R^{\mathfrak{A}}$. Now clearly $\mathfrak{A}\vDash R[a_{1},a_{2}]$ if and only if $\mathbb{U}\vDash\exists u\in T\;\exists v\in T\;\exists w\in T\;\Big{(}w=(a_{1},a_{2})\land(u,v)=\mathfrak{A}\land w\in v\Big{)}$, i.e. if and only if $\mathbb{U}\vDash\bar{\varphi}[a_{1},a_{2},\mathfrak{A},T]$. This deals with the case where $\varphi$ is an atomic $\mathcal{L}$ formula. The propositional connectives are easily handled. If $\varphi(x_{1},\dots,x_{n})\equiv\exists z\psi(z,x_{1},\dots,x_{n})$, then
$\mathfrak{A}\vDash\varphi[a_{1},\dots,a_{n}]$ $\Leftrightarrow\mathfrak{A}\vDash\psi[b,a_{1},\dots,a_{n}]\quad\text{some $b\in A$},$
$\Leftrightarrow\mathbb{U}\vDash\bar{\psi}[b,a_{1},\dots,a_{n},\mathfrak{A},T]\quad\text{some $b\in A$, by induction hypothesis},$
$\Leftrightarrow\mathbb{U}\vDash\underbrace{\exists u\in T\;\exists v\in T\;\exists z\in T\;((u,v)=\mathfrak{A}\land z\in u\land\bar{\psi}(z,a_{1},\dots,a_{n},\mathfrak{A},T))}_{\bar{\varphi}[a_{1}.\dots,a_{n},\mathfrak{A},T]}$
This completes the induction.
Since in $\mathbb{V}$ we have that ${}^{*}\mathfrak{A}\in{}^{*}T$ and ${}^{*}T$ is transitive, we obtain similarly that
${}^{*}\mathfrak{A}\vDash\varphi[a_{1},\dots,a_{n}]\qquad\text{if and only if}\qquad\mathbb{V}\vDash\bar{\varphi}[a_{1},\dots,a_{n},{}^{*}\mathfrak{A},{}^{*}T]\qquad\text{(for $a_{1},\dots,a_{n}\in{}^{*}A$)}$
for every $\mathcal{L}$ formula $\varphi$. But since $\bar{\varphi}$ is then a bounded $\mathcal{L}_{\in}$-formula and * is a transfer map, we have
$\mathbb{U}\vDash\bar{\varphi}[a_{1},\dots,a_{n},\mathfrak{A},T]\qquad\text{if and only if}\qquad\mathbb{V}\vDash\bar{\varphi}[{}^{*}a_{1},\dots,{}^{*}a_{n},{}^{*}\mathfrak{A},{}^{*}T],$
from which it follows that
$\mathfrak{A}\vDash\varphi[a_{1},\dots,a_{n}]\qquad\text{if and only if}\qquad{}^{*}\mathfrak{A}\vDash\varphi[{}^{*}a_{1},\dots,{}^{*}a_{n}].$
Thus $*:\mathfrak{A}\to{}^{*}\mathfrak{A}$ is an elementary embedding.
$\dashv$
3 Standard, Internal and External Objects
In this section, we suppose that we’re working with a transfer map $*:\mathbb{U}\to\mathbb{V}$ for a base set $X\in\mathbb{U}$.
3.1 Standard and Internal Objects
Definition 3.1
  1. (a)
    An element $v\in\mathbb{V}$ is said to be standard if and only if there is $u\in\mathbb{U}$ such that ${}^{*}u=v$. Thus the standard objects in $\mathbb{V}$ are those in the range of the *-map.
    We denote the set of standard objects in $\mathbb{V}$ by ${}^{\sigma}\mathbb{U}$.
  2. (b)
    An element $v\in\mathbb{V}$ is said to be internal if and only if there is $A\in\mathbb{U}$ such that $v\in{}^{*}A$. Thus the internal objects in $\mathbb{V}$ are those that belong to a standard set.
    We denote the set of internal objects in $\mathbb{V}$ by ${}^{*}\mathbb{U}$.
  3. (c)
    Sets in $\mathbb{V}$ which are not internal are said to be external.
$\Box$
Here is the reason that internal objects play an important role: A transfer map $*:\mathbb{U}\to\mathbb{V}$ transfers the truth of bounded $\mathcal{L}_{\in}$-formulas from $\mathbb{U}$ to $\mathbb{V}$. Thus, for example if $a,b\in\mathbb{U}$, and $\varphi\equiv\forall x\in a\;\exists y\in b\;\exists z\in y\;\varphi(x,y,z,a,b)$, then with ${}^{*}\varphi\equiv\forall x\in{}^{*}a\;\exists y\in{}^{*}b\;\exists z\in y\;\varphi(x,y,z,{}^{*}a,{}^{*}b)$ we have $\mathbb{U}\vDash\varphi$ if and only if $\mathbb{V}\vDash{}^{*}\varphi$. Now to check if $\mathbb{V}\vDash{}^{*}\varphi$, we need merely check ${}^{*}\varphi$ over elements $x\in{}^{*}a,y\in{}^{*}b,z\in y$, i.e. we need only consider internal $x,y,z$. Then if ${}^{*}\varphi$ is true in $\mathbb{V}$, $\varphi$ is true in $\mathbb{U}$.
Remarks 3.2
  1. (a)
    ${}^{*}\mathbb{U}\subseteq\mathbb{V}$ is transitive, i.e. if $A\in\mathbb{V}$ is internal, and $a\in A$, then $a$ is internal: Since $A$ is internal, there is $U\in\mathbb{U}$ such that $A\in{}^{*}U$. By definition of universe, there is a transitive $T\in\mathbb{U}$ such that $U\subseteq T$, and hence $A\in{}^{*}T$. Since ${}^{*}T$ is also transitive, we have $a\in{}^{*}T$, which shows that $A$ is internal.
  2. (b)
    Note that since ${}^{*}x=x$ for every member $x\in X$ of the base set, each $x\in X$ is standard.
  3. (c)
    Note that every standard object is internal. Indeed, ${}^{*}u\in\{{}^{*}u\}={}^{*}\{u\}$ for every standard object ${}^{*}u\in\mathbb{V}$. hence ${}^{\sigma}\mathbb{U}\subseteq{}^{*}\mathbb{U}$.
  4. (d)
    Note that $v\in\mathbb{V}$ is internal if and only if there is transitive $T\in\mathbb{U}$ such that $v\in{}^{*}T$: Indeed, if $v$ is internal, there is $U\in\mathbb{U}$ such that $v\in{}^{*}U$. But since $\mathbb{U}$ is a universe, there is a transitive $T\in\mathbb{U}$ such that $U\subseteq T$, and so $v\in{}^{*}U\subseteq{}^{*}T$.
  5. (e)
    Better yet, given any internal objects $v_{1},\dots,v_{n}$ in $\mathbb{V}$, there is a transitive $T\in\mathbb{U}$ such that $v_{1},\dots,v_{n}\in{}^{*}T$: For there are $A_{1},\dots,A_{n}\in\mathbb{U}$ such that $v_{i}\in{}^{*}A_{i}$. But then there is transitive $T\in\mathbb{U}$ such that $A_{1}\cup\dots\cup A_{n}\subseteq T$, from which it follows that $v_{i}\in{}^{*}A_{i}\subseteq{}^{*}T$.
$\Box$
Theorem 3.3
(Internal/Standard Definition Principle) Let $\varphi(y,x_{1},\dots,x_{n})$ be a bounded $\mathcal{L}_{\in}$-formula, and let $B,A_{1},\dots,A_{n}\in\mathbb{V}$ be internal (resp. standard). Then the set
$\{y\in B:\mathbb{V}\vDash\varphi[y,A_{1},\dots,A_{n}]\}$
is internal (resp. standard).
Proof: The case where $B,A_{1},\dots,A_{n}\in\mathbb{V}$ are standard follows directly from Proposition 2.2.
Now suppose that $B,A_{1},\dots,A_{n}\in\mathbb{V}$ are internal. Let $T\in\mathbb{U}$ be a transitive set such that $B,A_{1},\dots,A_{n}\in{}^{*}T$. Consider the bounded $\mathcal{L}_{\in}$-formula
$\psi(t,p)\equiv\forall b\in t\;\forall x_{1}\in t\;\dots\forall x_{n}\in t\;\exists u\in p\;\forall y\in t\;(y\in u\leftrightarrow(y\in b\land\varphi(y,x_{1},\dots,x_{n})).$
Then $\mathbb{U}\vDash\psi[T,\mathcal{P}(T)]$, since if $b,x_{1},\dots,x_{n}\in T$, then the set $\{y\in b:\varphi[y,x_{1},\dots,x_{n}]\}$ is a subset of $b$, thus of $T$, and hence a member of $\mathcal{P}(T)$. By transfer $\mathbb{V}\vDash\psi[{}^{*}T,{}^{*}\mathcal{P}(T)]$. In particular, it follows that
$\mathbb{V}\vDash\exists u\in{}^{*}\mathcal{P}(T)\;\forall y\in{}^{*}T\;(y\in u\leftrightarrow(y\in B\land\varphi(y,A_{1},\dots,A_{n})).$
Thus there is $u\in{}^{*}\mathcal{P}(T)$ such that
$u\cap{}^{*}T=\{y\in B:\varphi(y,A_{1},\dots,A_{n})\}\cap{}^{*}T.$
But since $u\in{}^{*}\mathcal{P}(T)$, we have $u\subseteq{}^{*}T$, so that $u\cap{}^{*}T=u$. On the other hand, since $B\subseteq{}^{*}T$ (by transitivity of ${}^{*}T$), we have that $\{y\in B:\varphi(y,A_{1},\dots,A_{n})\}\cap{}^{*}T=\{y\in B:\varphi(y,A_{1},\dots,A_{n})\}$, and hence $u=\{y\in B:\varphi(y,A_{1},\dots,A_{n})\}$. Finally, since $u\in{}^{*}\mathcal{P}(T)$, $u$ is internal.
$\dashv$
Lemma 3.4
Suppose that $U,V$ are sets, that $U$ is transitive and that $U\subseteq V$. Then $(U,\in)$ is a bounded elementary submodel of $(V,\in)$.
Proof: The proof that $U\vDash\varphi$ if and only if $V\vDash\varphi$, for any bounded $\mathcal{L}_{\in}$-formula $\varphi$ is by induction on formula complexity. The only troublesome case in the induction step for formulas of the form $\varphi(x,y)\equiv\exists z\in x\;\psi(x,y,z)$, where $\psi$ is a bounded $\mathcal{L}_{\in}$-formula. Suppose that $a,b\in U$ and that $V\vDash\varphi(a,b)$. Then there is $c\in V$ such that $c\in a$ and $V\vDash\psi(a,b,c)$. As $U$ is transitive and $c\in a$, we also have $c\in U$. By induction hypothesis, therefore, we have $U\vDash\psi(a,b,c)$, from which it follows that $U\vDash\varphi(a,b)$.
$\dashv$
Proposition 3.5
The map $*:(\mathbb{U},\in)\hookrightarrow({}^{*}\mathbb{U},\in)$ and the inclusion $({}^{*}\mathbb{U},\in)\hookrightarrow(\mathbb{V},\in)$ are bounded elementary embeddings.
Proof: We already know that $({}^{*}\mathbb{U},\in)$ is a transitive submodel of $(\mathbb{V},\in)$. Hence by the previous Lemma, it follows that $({}^{*}\mathbb{U},\in)\hookrightarrow(\mathbb{V},\in)$ is a bounded elementary embedding.
Now if $\varphi(x,y)$ is a bounded $\mathcal{L}_{\in}$-formula and $a,b\in\mathbb{U}$, then we have $\mathbb{U}\vDash\varphi(a,b)$ if and only if $\mathbb{V}\vDash\varphi({}^{*}a,{}^{*}b)$ if and only if ${}^{*}\mathbb{U}\vDash\varphi({}^{*}a,{}^{*}b)$, which shows that $*:(\mathbb{U},\in)\hookrightarrow({}^{*}\mathbb{U},\in)$ is a bounded elementary embedding.
$\dashv$
3.2 Examples of Standard and Internal Objects
In this section, assume that $*:\mathbb{U}\hookrightarrow\mathbb{V}$ is a transfer map over $X$.
Lemma 3.6
  1. (a)
    If $a_{1},\dots,a_{n}$ are internal, so are $\{a_{1},\dots,a_{n}\}$ and $(a_{1},\dots,a_{n})$.
  2. (b)
    If $A,B\in\mathbb{V}$ are internal, so are $A\cup B,A\cap B,A-B,A\times B$.
  3. (c)
    If $\mathcal{A}$ is internal, then $\bigcup\mathcal{A}$ and $\bigcap\mathcal{A}$ are internal.
  4. (d)
    If a binary relation $R$ is internal, then so are $\text{\rm dom}(R),\text{\rm ran}(R)$, and $R^{-1}$. If $C\subseteq\text{\rm dom}(R)$ is internal, then $R[C]\subseteq\text{\rm ran}(R)$ is internal.
  5. (e)
    If $R,S$ are internal binary relations, their composition $R\circ S$ is internal.
  6. (f)
    If a function $f$ is internal, and $a\in\text{\rm dom}(f)$, then $f(a)$ is internal.
Proof: (a) For example, there is transitive $T\in\mathbb{U}$ such that $a_{1},\dots,a_{n}\in{}^{*}T$. Then
$\{a_{1},\dots,a_{n}\}=\{x\in{}^{*}T:x=a_{1}\lor\dots\lor x=a_{n}\}.$
This set is internal, by the internal definition principle. Now if $a,b$ are internal, then so are $\{a\},\{a,b\}$ and hence so is $\{\{a\},\{a,b\}\}=(a,b)$, etc.
(b) For example, if $A,B$ are internal, then there is transitive $T\in\mathbb{U}$ such that $A,B\subseteq{}^{*}T$. Since ${}^{*}T\times{}^{*}T={}^{*}(T\times T)$, we see that
$A\times B=\{x\in{}^{*}(T\times T):\exists y\in A\;\exists z\in B\;(x=(y,z))\},$
which is internal, by the internal definition principle and the fact that the formula $x=(y,z)$ is bounded, according to Lemma 2.1(c).
(c) Choose transitive $T\in\mathbb{U}$ such that $\mathcal{A}\in{}^{*}T$. Then $\bigcup\mathcal{A}\subseteq{}^{*}T$, and hence
$\bigcup\mathcal{A}=\{x\in{}^{*}T:\exists A\in\mathcal{A}\;(x\in A)\},$
which is internal by the internal definition principle. For $\bigcap\mathcal{A}\subseteq{}^{*}T$, just replace $\exists A\in\mathcal{A}\;(x\in A)$ by $\forall A\in\mathcal{A}\;(x\in A)$.
(d) For example, if $T\in\mathbb{U}$ is transitive so that $R\in\mathbb{}^{*}T$, then $\text{\rm dom}(R),\text{\rm ran}(R)\subseteq{}^{*}T$, and apply the internal definition principle to
$\text{\rm dom}(R)$ =$=\{x\in T:\exists y\in T\;(x,y)\in R\},$
$\text{\rm ran}(R)$ =$=\{y\in{}^{*}T:\exists x\in T\;(x,y)\in R\},$
$R^{-1}$ =$=\{w\in{}^{*}(T\times T):\exists x\in{}^{*}T\;\exists y\in{}^{*}T\;\exists v\in R\;(v=(x,y)\land w=(y,x))\},$
$R[C]$ =$=\{y\in{}^{*}T:\exists x\in C\;((c,y)\in R)\}.$
(e) Choose $T\in\mathbb{U}$ transitive so that $R,S\in{}^{*}T$. Then
$R\circ S=\{w\in{}^{*}(T\times T):\exists x\in{}^{*}T\;\exists y\in{}^{*}T\;\exists z\in{}^{*}T\;((x,y)\in S\land(y,z)\in R\land w=(x,z))\},$
where $(x,y)\in S$ is short for $\exists s\in S\;(s=(x,y))$, and $s=(x,y)$ is a bounded formula, by Lemma 2.1(c).
(d) $f$ is a binary relation, so $\text{\rm ran}(f)$ is internal. Since $f(a)\in\text{\rm ran}(f)$, $f(a)$ is internal, by transitivity of ${}^{*}\mathbb{U}$.
$\dashv$
Lemma 3.7
Let $*:\mathbb{U}\hookrightarrow\mathbb{V}$ be a transfer map for $X$. Then
${}^{*}V_{n}(X)=V_{n}({}^{*}X)\cap{}^{*}\mathbb{U}\quad\text{ i.e.}\quad{}^{*}V_{n}(X)=\{x\in V_{n}({}^{*}X):x\text{ is internal}\}.$
Proof: Recall that if $\mathbb{U}$ is a universe over $X$, then $V(X)\subseteq\mathbb{U}$. Since $\mathbb{V}$ is a universe over ${}^{*}X$ (by Lemma 1.10), we have $V({}^{*}X)\subseteq\mathbb{V}$. Now by Lemma 2.1,
$\mathbb{U}\vDash\forall x\in V_{n}(X)\;\varphi_{6,n}(X,x),\quad\text{so}\quad\mathbb{V}\vDash\forall x\in{}^{*}V_{n}(X)\;\varphi_{6,n}({}^{*}X,x),$
from which it follows that ${}^{*}V_{n}(X)\subseteq\{x\in V_{n}({}^{*}X):x\text{ is internal}\}$. Now suppose that ${}^{*}V_{n}(X)\subsetneqq\{x\in V_{n}({}^{*}X):x\text{ is internal}\}$. Then there must be an internal $a\in V_{n}({}^{*}X)$ which is not in ${}^{*}V_{n}(X)$. Since $a$ is internal, there is $B\in\mathbb{U}$ such that $a\in{}^{*}B$. Hence
$\mathbb{V}\vDash\exists a\in{}^{*}B[\varphi_{6,n}({}^{*}X,a)\land a\not\in{}^{*}V_{n}(X)],$
from which we obtain
$\mathbb{U}\vDash\exists a\in B[\varphi_{6,n}(X,a)\land a\not\in V_{n}(X)],$
which is impossible.
$\dashv$
Lemma 3.8
Let $A\in\mathbb{U}$ be a set. Then ${}^{*}\mathcal{P}(A)=\mathcal{P}({}^{*}A)\cap{}^{*}\mathbb{U}$ is the set of all internal subsets of ${}^{*}A$.
Proof: Transfer of the true bounded sentence $\forall B\in\mathcal{P}(A)\;(B\subseteq A)$ shows that every $B\in{}^{*}\mathcal{P}(A)$ is a subset of ${}^{*}A$, and hence in $\mathcal{P}({}^{*}A)$. Since every $B\in{}^{*}\mathcal{P}(A)$ is obviously internal, we have ${}^{*}\mathcal{P}(A)\subseteq\mathcal{P}({}^{*}A)\cap{}^{*}\mathbb{U}$.
Now suppose that $B$ is an arbitrary internal subset of ${}^{*}A$. Then there is $C\in\mathbb{U}$ such that $B\in{}^{*}C$. Now transfer of the true sentence $\forall B\in C\;(B\subseteq A\to B\in\mathcal{P}(A))$ shows that $B\in{}^{*}\mathcal{P}(A)$, from which ${}^{*}\mathcal{P}(A)\supseteq\mathcal{P}({}^{*}A)\cap{}^{*}\mathbb{U}$.
$\dashv$
Lemma 3.9
If $A\in\mathbb{V}$ is internal, then so is the set $\mathcal{P}(A)\cap{}^{*}\mathbb{U}$ of internal subsets of $A$.
Proof: Choose $T\in\mathbb{U}$ transitive such that $A\subseteq{}^{*}T$. If $B\subseteq A$ is internal, then $B$ is an internal subset of ${}^{*}T$, and hence $B\in{}^{*}\mathcal{P}(T)$. So
$\mathcal{P}(A)\cap{}^{*}\mathbb{U}=\{B\in{}^{*}\mathcal{P}(T):B\subseteq A\}.$
Since $A,B,{}^{*}\mathcal{P}(T)$ are internal, it follows by the internal definition principle that $\mathcal{P}(A)\cap{}^{*}\mathbb{U}$ is internal.
$\dashv$
Lemma 3.10
Let $\mathcal{A}\in\mathbb{U}$ be a family of sets. Then
{${}^{*}\{\mathcal{P}(A):A\in\mathcal{A}\}=\{P_{A}:A\in{}^{*}\mathcal{A}\},$
where $P_{A}:=\mathcal{P}(A)\cap{}^{*}\mathbb{U}$ is the set of internal subsets of $A$.
Proof: Let $\mathcal{X}:=\{\mathcal{P}(A):A\in\mathcal{A}\}$, and let $T\in\mathbb{U}$ be transitive such that $\mathcal{X}\subseteq T$. Observe that
$\mathcal{X}=\{P\in T:\exists A\in\mathcal{A}\;\forall B\in T\;(B\in P\leftrightarrow B\subseteq A)\},$
so that
${}^{*}\mathcal{X}=\{P\in{}^{*}T:\exists A\in{}^{*}\mathcal{A}\;\forall B\in{}^{*}T\;(B\in P\leftrightarrow B\subseteq A)\}.$
Thus $P\in{}^{*}\mathcal{X}$ if and only if $P\in{}^{*}T$ and there is $A\in{}^{*}\mathcal{A}$ such that $P\cap{}^{*}T=\mathcal{P}(A)\cap{}^{*}T$. Since $P\in{}^{*}T$ and ${}^{*}T$ is transitive, we have $P=\mathcal{P}(A)\cap{}^{*}T$ for some $A\in{}^{*}\mathcal{A}$. It follows that $P\subseteq P_{A}$. Now if $B\in P_{A}$, then $B\subseteq A$ and there is $S\in\mathbb{U}$ such that $B\in{}^{*}S$. Since
$\mathbb{U}\vDash\forall B\in S\;\forall A\in\mathcal{A}\;(B\subseteq A\to B\in T),$
we have that
$\mathbb{V}\vDash\forall B\in{}^{*}S\;\forall A\in{}^{*}\mathcal{A}\;(B\subseteq A\to B\in{}^{*}T),$
from which it follows that $B\in{}^{*}T$, and thus that $B\in T$. Hence $P=\mathcal{P}(A)\cap{}^{*}T$ is the set $P_{A}=\mathcal{P}(A)\cap{}^{*}\mathbb{U}$ of all internal subsets of $A$. It therefore follows that $P\in{}^{*}\mathcal{X}$ if and only if $P=P_{A}$ for some $A\in{}^{*}\mathcal{A}$.
$\dashv$
Lemma 3.11
Let $\mathcal{A}\in\mathbb{U}$ be a family of sets. Then ${}^{*}(\bigcup\mathcal{A})=\bigcup{}^{*}\mathcal{A}$.
Proof: Just transfer the true bounded formula
$[\forall a\in\bigcup\mathcal{A}\;\exists A\in\mathcal{A}\;(a\in A)]\land[\forall A\in\mathcal{A}\;\forall a\in A\;(a\in\bigcup\mathcal{A})].$
$\dashv$
(Note that every element of an element of ${}^{*}\mathcal{A}$ is internal, since ${}^{*}\mathbb{U}$ is transitive. Thus automatically $\bigcup{}^{*}\mathcal{A}\subseteq{}^{*}\mathbb{U}$, and we do not need to write ${}^{*}(\bigcup\mathcal{A})=(\bigcup{}^{*}\mathcal{A})\cap{}^{*}\mathbb{U}$, as for some of the other operations in this section.)
We often deal with unions, intersections and products of indexed families of sets. Suppose, for example, that $\mathcal{A}:=\{A_{i}:i\in I\}\in\mathbb{U}$ is an indexed family of sets in $\mathbb{U}$, and let $f:I\to\mathcal{A}:i\mapsto A_{i}$ be the indexing function, where $f\in\mathbb{U}$. Then ${}^{*}f:{}^{*}I\to{}^{*}\mathcal{A}$ is a function, with ${}^{*}f({}^{*}i)={}^{*}(f(i))={}^{*}A_{i}$ when $i\in I$. For $i\in{}^{*}I-I$, define ${}^{*}A_{i}:={}^{*}f(i)$ — but note that if $i\in{}^{*}I-I$, then ${}^{*}A_{i}$ is not necessarily the *-value of a member of $\mathbb{U}$. Now since $f$ is surjective, so is ${}^{*}f$, and hence ${}^{*}\mathcal{A}=\{{}^{*}A_{i}:i\in{}^{*}I\}$. Hence by the previous result,
${}^{*}\left(\bigcup_{i\in I}A_{i}\right)={}^{*}\bigcup\mathcal{A}=\bigcup{}^{*}\mathcal{A}=\bigcup_{i\in{}^{*}I}{}^{*}A_{i}.$
Lemma 3.12
Let $A,B\in\mathbb{U}$ be sets. Then ${}^{*}(B^{A})={}^{*}B^{{}^{*}A}\cap{}^{*}\mathbb{U}$ is the set of all internal functions ${}^{*}A\to{}^{*}B$.
Proof: Consider the bounded formula $\varphi_{5}(f,A,B)$ of Lemma 2.1, which asserts that $f:A\to B$, i.e. that $f\in B^{A}$. Transfer of the bounded formula $\forall f\in B^{A}\;\varphi_{5}(f,A,B)$ shows that every member of ${}^{*}(B^{A})$ is an internal function ${}^{*}A\to{}^{*}B$. Conversely, suppose that $f\in{}^{*}B^{{}^{*}A}$ is internal. Let $C\in\mathbb{U}$ so that $f\in C$. Transfer of $\forall f\in C\;(\varphi_{5}(f,A,B)\to f\in B^{A})$ shows that $f\in{}^{*}(B^{A})$.
$\dashv$
Lemma 3.13
Suppose that $A,B\in\mathbb{V}$ are internal. Then so is the set $B^{A}\cap{}^{*}\mathbb{U}$ of internal functions from $A$ to $B$.
Proof: We have seen that $A\times B$ is internal, i.e. there is a transitive $T\in\mathbb{U}$ such that $A\times B\in{}^{*}T$. Let $P:=\{X\subseteq A\times B:X\text{ is }internal\}$, which is internal by Lemma 3.9. Clearly,
$B^{A}\cap{}^{*}\mathbb{U}=\{f\in P:\varphi_{5}(f,A,B)\},$
where $\varphi_{5}(f,A,B)$ is the bounded formula of Lemma 2.1 which asserts that $f:A\to B$. By the internal definition principle, $B^{A}\cap{}^{*}\mathbb{U}$ is internal.
$\dashv$
Lemma 3.14
Let $\mathcal{A}\in\mathbb{U}$ be a family of sets. Then ${}^{*}(\prod\mathcal{A})=(\prod{}^{*}\mathcal{A})\cap{}^{*}\mathbb{U}$.
Proof: Recall that $\prod\mathcal{A}$ is the set of all choice functions, i.e. that $f\in\prod\mathcal{A}$ if and only if $f:\mathcal{A}\to\bigcup\mathcal{A}$ is such that $f(A)\in A$ for all $A\in\mathcal{A}$. Thus transfer of
$\forall f\in\prod\mathcal{A}\;\Big{(}\varphi_{5}(f,\mathcal{A},\bigcup\mathcal{A})\land\forall A\in\mathcal{A}\;(f(A)\in A)\Big{)}$
shows that every member of ${}^{*}(\prod\mathcal{A})$ is an internal choice function ${}^{*}\mathcal{A}\to{}^{*}(\bigcup\mathcal{A})=\bigcup{}^{*}\mathcal{A}$, and thus a member of $\prod{}^{*}\mathcal{A}\cap{}^{*}\mathbb{U}$.
Conversely if $f\in\prod{}^{*}\mathcal{A}\cap{}^{*}\mathbb{U}$, then there is $C\in\mathbb{U}$ such that $f\in{}^{*}C$. Then transfer of
$\forall f\in C\;\Big{(}\big{[}\varphi_{5}(f,\mathcal{A},\bigcup\mathcal{A})\land\forall A\in\mathcal{A}\;(f(A)\in A)\big{]}\to f\in\prod\mathcal{A}\Big{)}$
shows that $f\in{}^{*}(\prod\mathcal{A})$.
$\dashv$
We can deal with indexed products in a manner very similar to the way we handled indexed unions: Let $\mathcal{A}=\{A_{i}:i\in I\}\in\mathbb{U}$ be an indexed family of sets. Then
${}^{*}\Big{(}\prod_{i\in I}A_{i}\Big{)}={}^{*}(\prod\mathcal{A})=\Big{(}\prod{}^{*}\mathcal{A}\Big{)}\cap{}^{*}\mathbb{U}=\Big{(}\prod_{i\in{}^{*}I}{}^{*}A_{i}\Big{)}\cap{}^{*}\mathbb{U}.$
Lemma 3.15
Let $\mathcal{A},\mathcal{B}\in\mathbb{U}$ be families of sets. Then
{${}^{*}\{A\times B:A\in\mathcal{A},B\in\mathcal{B}\}=\{A\times B:A\in{}^{*}\mathcal{A},B\in{}^{*}\mathcal{B}\}.$
Proof: Let $\mathcal{X}:=\{A\times B:A\in\mathcal{A},B\in\mathcal{B}\}$, and choose a transitive $T\in\mathbb{U}$ such that $\mathcal{X}\subseteq T$. The bounded formula $\varphi_{4}$ of Lemma 2.1 asserts that $\varphi_{4}(P,A,B)$ holds if and only if $P=A\times B$. Thus
$\mathcal{X}=\{P\in T:\exists A\in\mathcal{A}\;\exists B\in\mathcal{B}\;\varphi_{4}(P,A,B)\}.$
Transfer then yields that ${}^{*}\mathcal{X}=\{A\times B:A\in{}^{*}\mathcal{A},B\in{}^{*}\mathcal{B}\}\cap{}^{*}T$. But if $A\in{}^{*}\mathcal{A},B\in{}^{*}\mathcal{B}$, then $A\times B$ is internal, and so there is $S\in\mathbb{U}$ such that $A\times B\in{}^{*}S$. But
$\mathbb{U}\vDash\forall P\in S\;\Big{(}\exists A\in\mathcal{A}\;\exists B\in\mathcal{B}\;\varphi_{4}(P,A,B)\to P\in T)\Big{)},$
so
$\mathbb{V}\vDash\forall P\in{}^{*}S\;\Big{(}\exists A\in{}^{*}\mathcal{A}\;\exists B\in{}^{*}\mathcal{B}\;\varphi_{4}(P,A,B)\to P\in{}^{*}T)\Big{)}.$
Hence $A\times B\in{}^{*}T$ for all $A\in{}^{*}\mathcal{A},B\in{}^{*}\mathcal{B}$. Thus ${}^{*}\mathcal{X}=\{A\times B:A\in{}^{*}\mathcal{A},B\in{}^{*}\mathcal{B}\}$.
$\dashv$
3.3 ${}^{\sigma}A$ is External if $A$ is Infinite
In this subsection, we assume that $*:\mathbb{U}\to\mathbb{V}$ is a nonstandard framework for a set $X$, i.e. a transfer map with the property that there is a countable set $C\in\mathbb{U}$ such that ${}^{\sigma}C\subsetneqq{}^{*}C$.
Theorem 3.16
If $*:\mathbb{U}\to\mathbb{V}$ is a nonstandard framework, then ${}^{\sigma}A$ is external whenever $A\in\mathbb{U}$ is infinite. Hence ${}^{\sigma}A\subsetneqq{}^{*}A$ whenever $A\in\mathbb{U}$ is infinite.
Proof: By definition of nonstandard framework, there is a countable set $C\in\mathbb{U}$ such that ${}^{\sigma}C\subsetneqq{}^{*}C$. We will begin by showing that the difference $D:={}^{*}C-{}^{\sigma}C$ is external. Suppose that $\{c_{m}:n\in\mathbb{N}\}$ enumerates $C$. This induces a well-ordering $\preceq$ on $C$ by $c_{n}\preceq c_{m}$ if and only if $n\leq m$. Observe that $\preceq\;\in\mathcal{P}(C\times C)\in\mathbb{U}$, so that $\preceq\;\in\mathbb{U}$. Now the assertion that $\preceq$ is a well-ordering is a bounded sentence, where the fact that every non-empty subset of $C$ has a $\preceq$-least element is given by the bounded sentence.
$\mathbb{U}\vDash\forall X\in\mathcal{P}(C)\;[x\neq\varnothing\to\exists x_{0}\in X\;\forall x\in X\;(x_{0}\preceq x)].$
Thus ${}^{*}\!\!\preceq\;\in\mathbb{V}$ is a linear ordering on ${}^{*}C$, and .
$\mathbb{V}\vDash\forall X\in{}^{*}\mathcal{P}(C)\;[x\neq\varnothing\to\exists x_{0}\in X\;\forall x\in X\;(x_{0}{}^{*}\!\!\preceq x)].$
This does not assert that ${}^{*}\!\!\preceq$ is a well-ordering on ${}^{*}C$, however, because it may not be the case that ${}^{*}\mathcal{P}(C)=\mathcal{P}({}^{*}C)$. What it does assert is that every non-empty internal subset of ${}^{*}C$ has a ${}^{*}\!\!\preceq$-least element, since we know that ${}^{*}\mathcal{P}(C)=\mathcal{P}({}^{*}C)\cap{}^{*}\mathbb{U}$ is the set of internal subsets of ${}^{*}C$. In particular, if $D$ is internal, there is a ${}^{*}\!\!\preceq$-least element $d_{0}$ of $D$.
Now for every $n\in\mathbb{N}$,
$\mathbb{U}\vDash\forall x\in C\;(x=c_{0}\lor x=c_{1}\lor\dots\lor x=c_{n}\lor x\succ c_{n}),$
and thus
$\mathbb{V}\vDash\forall x\in{}^{*}C\;(x={}^{*}c_{0}\lor x={}^{*}c_{1}\lor\dots\lor x={}^{*}c_{n}\lor x\>{}^{*}\!\!\succ{}^{*}c_{n})$
Now since $d_{0}\not\in{}^{\sigma}C$, we see that $d_{0}\neq{}^{*}c_{n}$ for any $n\in\mathbb{N}$, and thus that $d_{0}\;{}^{*}\!\!\succ{}^{*}c_{n}$ for all $n$. Furthermore, in $\mathbb{U}$ every element of $C$ has an immediate $\preceq$-predecessor, excepting of course the least element $c_{0}$:
$\mathbb{U}\vDash\forall x\in C\;(x\neq c_{0}\to\exists y\in C\;(y\prec x\land\forall z\in C\;(z\prec x\to z\preceq y))).$
Hence
$\mathbb{V}\vDash\forall x\in{}^{*}C\;(x\neq{}^{*}c_{0}\to\exists y\in{}^{*}C\;(y\;{}^{*}\!\!\prec x\land\forall z\in{}^{*}C\;(z\;{}^{*}\!\!\prec x\to z\;{}^{*}\!\!\preceq y))).$
So if $D$ is internal, its least element $d_{0}$ has an immediate predecessor $d_{-1}\in{}^{*}C$. Now clearly we cannot have $d_{-1}={}^{*}c_{n}$ for any $n\in\mathbb{N}$, for else necessarily $d_{0}={}^{*}c_{n+1}\in{}^{\sigma}C$. So $d_{-1}\in D$ — contradicting the fact that $d_{0}$ is the least element of $D$. Thus the assumption that $D$ is internal leads to contradiction, i.e. $D$ is external.
Now since ${}^{*}C$ is internal, it follows that ${}^{\sigma}C$ is external, because the difference of two internal sets is internal.
Now suppose that $A\in\mathbb{U}$ is infinite, and that $f\in\mathbb{U}$ is a surjection $f:A\twoheadrightarrow C$. It follows easily that ${}^{*}f:{}^{*}A\twoheadrightarrow{}^{*}C$. Since ${}^{*}(f(a))={}^{*}f({}^{*}a)$ for all $a\in A$, we have ${}^{*}f[{}^{\sigma}A]={}^{\sigma}C$. If ${}^{\sigma}A$ is internal, then by Lemma 3.6 it would follows that ${}^{\sigma}C$ is internal as well. As this is false, ${}^{\sigma}A$ is external.
Finally, it is always the case that ${}^{\sigma}A\subseteq{}^{*}A$. If ${}^{\sigma}A={}^{*}A$, then ${}^{\sigma}A$ would be internal. If $A$ is infinite, this is not the case. Hence ${}^{\sigma}A\subsetneqq{}^{*}A$ if $A\in\mathbb{U}$ is infinite.
$\dashv$
4 Hyperfinite Sets
4.1 The Set ${}^{*}\mathbb{N}$ of Hypernatural Numbers
Throughout this section, assume that $*:\mathbb{U}\to\mathbb{V}$ is a nonstandard embedding for an infinite set $X$. Without loss of generality (e.g. by renaming elements) we may assume that $\mathbb{N}\subseteq X$. Then as $\mathcal{P}(X)\in\mathbb{U}$ and $\mathbb{U}$ is transitive, also $\mathbb{N}\in\mathbb{U}$. Since ${}^{*}x=x$ for all $x\in X$, it follows in particular that ${}^{\sigma}\mathbb{N}=\mathbb{N}$. Since * is nonstandard, $\mathbb{N}\subsetneqq{}^{*}\mathbb{N}$.
Given a fixed but arbitrary $N\in\mathbb{N}$, transfer of the true bounded sentence
$\forall n\in\mathbb{N}\;(n=1\lor n=2\lor\dots\lor n=N\lor n>N)$
shows that any member of ${}^{*}\mathbb{N}-\mathbb{N}$ is $>N$. As this is true for all $N\in\mathbb{N}$, it follows that every member of ${}^{*}\mathbb{N}-\mathbb{N}$ is greater than any natural number, and thus said to be infinite. We thus define $\mathbb{N}_{\infty}:={}^{*}\mathbb{N}-\mathbb{N}$ to be the set of infinite natural numbers, so that we have the disjoint union ${}^{*}\mathbb{N}=\mathbb{N}\cup\mathbb{N}_{\infty}$. As $\mathbb{N}={}^{\sigma}\mathbb{N}$ is external, so is $\mathbb{N}_{\infty}$.
Now the structure $(\mathbb{N},+,\dotdiv,\cdot,0,1)\in\mathbb{U}$ is elementarily equivalent to the corresponding structure $({}^{*}\mathbb{N},+,\dotdiv,\cdot,0,1)\in{}^{*}\mathbb{U}$ (where $n\dotdiv m:=n-m$ if $n\geq m$ and $:=0$ else). We should really have written $({}^{*}\mathbb{N},{}^{*}+,{}^{*}\dotdiv,{}^{*}\cdot,0,1)$, but we drop the stars on the arithmetic operations for easier reading. It is clear that if $n\in\mathbb{N}_{\infty}$ then $n-1\in\mathbb{N}_{\infty}$ also, and thus $\mathbb{N}_{\infty}$ has no least element, i.e. ${}^{*}\mathbb{N}$ is not well-ordered. However:
Theorem 4.1
Every non-empty internal subset of ${}^{*}\mathbb{N}$ has a least element.
Proof: Let $A\subseteq{}^{*}\mathbb{N}$ be internal and non-empty. Then $A\in\mathcal{P}({}^{*}\mathbb{N})\cap{}^{*}\mathbb{U}={}^{*}\mathcal{P}(\mathbb{N})$. Transfer of the true bounded sentence
$\forall A\in\mathcal{P}(\mathbb{N})\;[A\neq\varnothing\to\exists a_{0}\in A\;\forall a\in A\;(a_{0}\leq a)]$
now shows that $A$ has a least element.
$\dashv$
Theorem 4.2
(Overflow and Underflow)
  1. (a)
    (Overflow) Let $N\in\mathbb{N}$, and suppose that $X\in{}^{*}\mathcal{P}(\mathbb{N})$ is an internal subset of ${}^{*}\mathbb{N}$ with the property that whenever $n\in\mathbb{N}$ satisfies $n\geq N$, then $n\in X$. Then there is $M\in\mathbb{N}_{\infty}$ such that whenever $n\in{}^{*}\mathbb{N}$ satisfies $N\leq n\leq M$, then $n\in X$
  2. (b)
    (Underflow) Let $M\in\mathbb{N}_{\infty}$, and suppose that $X\in{}^{*}\mathcal{P}(\mathbb{N})$ is an internal subset of ${}^{*}\mathbb{N}$ with the property that whenever $n\in{}^{*}\mathbb{N}_{\infty}$ satisfies $n\leq M$, then $n\in X$. Then there is $N\in\mathbb{N}$ such that whenever $n\in{}^{*}\mathbb{N}$ satisfies $N\leq n\leq M$, then $n\in X$
Proof: (a) By the internal definition principle, the set $Y:=\{n\in{}^{*}\mathbb{N}:n>N\land n\in{}^{*}\mathbb{N}-X\}$ is internal. If $Y=\varnothing$, then every $M\in\mathbb{N}_{\infty}$ has the desired property. If $Y$ is non-empty, then it must have a least element $K$, i.e. $K$ is least such that $K>N$ and $K\not\in X$. By the assumption on $X$ we must have $K\in\mathbb{N}_{\infty}$. It follows that every infinite natural number which is $<K$ belongs to $X$, so let $M:=K-1$.
(b) By the internal definition principle, the set $Z:=\{k\in{}^{*}\mathbb{N}:\forall n\in{}^{*}\mathbb{N}\;(k\leq n\leq M\to n\in X)\}$ is internal. Then $Z\neq\varnothing$, since $M\in\mathbb{Z}$, and so $Z$ has a least member $N$. As $N\leq M$, the assumption on $X$ implies that we must have $N\in\mathbb{N}$.
$\dashv$
4.2 Hyperfinite sets
In this section, we suppose that $*:\mathbb{U}\to\mathbb{V}$ is a transfer map over a set $X$, where $\mathbb{N}\subseteq X$. Recall that the $<$-relation on $\mathbb{N}$ is just a subset of $<\;\subseteq\mathbb{N}\times\mathbb{N}$, so to say that $m<n$ is equivalent to saying $(m,n)\in\;<$.
Definition 4.3
A set $A\in\mathbb{V}$ is said to be hyperfinite (or *-finite) if and only if there is an internal bijection $f:\{0,1,\dots,n-1\}\to A$ for some $n\in{}^{*}\mathbb{N}$.
In that case, we denote #$\#A=n$.
$\Box$
Observe that if $f:\{0,1,\dots,n-1\}\to A$ is an internal bijection, then $A=\text{\rm ran}(f)$ is internal, i.e. every hyperfinite set is internal.
Observe also that there is a bounded formula $\psi(A,f,n,U)$ which asserts that $f:U\to A$ is a bijection, and that $U=\{0,\dots,n-1\}$: Indeed, $\psi$ is the conjunction of
  • $U\subseteq\mathbb{N}\land\forall m\in\mathbb{N}\;(m<n\leftrightarrow m\in U)$ — i.e. $U=\{0,1,\dots n-1\}$.
  • $\forall x\in f\;\exists u\in U\;\exists a\in A\;(x=(u,a))$ — i.e. $f$ is a binary relation, with $\text{\rm dom}(f)\subseteq U$ and $\text{\rm ran}(f)\subseteq A$.
    (Recall that the bounded formula $\varphi_{2,2}(c,a,b)$ of Lemma 2.1 asserts that $c=(a,b)$.)
  • $\forall u\in U\;\exists a\in A\;((u,a)\in f)$ — i.e. $\text{\rm dom}(f)\supseteq U$.
  • $\forall a\in A\;\exists u\in U\;((u,a)\in f)$ — i.e. $A\subseteq\text{\rm ran}(f)$.
  • $\forall u\in U\;\forall a\in A\;\forall b\in A\;(u,a)\in f\land(u,b)\in f\to a=b)$ — i.e $f$ is a function.
  • $\forall u\in U\;\forall v\in U\;(\exists a\in A\;((u,a)\in f\land(v,a)\in f)\to u=v)$ — i.e. $f$ is 1-1.
Now define the formula   $\Psi(A,f,n)\equiv\exists U\in\mathcal{P}(\mathbb{N})\;\psi(A,f,n,U)$.
Then $\Psi(A,f,n)$ is a bounded formula which asserts that $f$ is a bijection $f:\{0,\dots,n-1\}\to A$. Transfer guarantees that ${}^{*}\Psi(A,f,n)$ asserts the same for $A,f\in{}^{*}\mathbb{U}$ and $n\in{}^{*}\mathbb{N}$. We thus see that, by definition, $A\in\mathbb{V}$ is hyperfinite if and only if there are an internal $f$ and an $n\in{}^{*}\mathbb{N}$ such that ${}^{*}\Psi(A,f,n)$. Suppose that $B\in\mathbb{U}$, and recall that ${}^{*}\mathcal{P}(B)=\mathcal{P}({}^{*}B)\cap{}^{*}\mathbb{U}$ is the set of all internal subsets of $B$. Let
$\mathcal{P}^{<\omega}(B):=\{C\subseteq B:C\text{ is finite}\}.$
Then ${}^{*}\mathcal{P}^{<\omega}(B)\subseteq{}^{*}\mathcal{P}(B)$. The next theorem characterizes the hyperfinite sets as members of some ${}^{*}\mathcal{P}^{<\omega}(B)$.
Theorem 4.4
A set $A\in\mathbb{V}$ is hyperfinite if and only if there is $B\in\mathbb{U}$ such that $A\in{}^{*}\mathcal{P}^{<\omega}(B)$.
Proof: ()($\Rightarrow$)( ⇒ ): Suppose that $A\in\mathbb{V}$ is hyperfinite, and let $f:\{0,1,\dots,n-1\}\to A$ be an internal bijection, for some $n\in{}^{*}\mathbb{N}$. Let $U:=\{0,1,\dots,n-1\}=\text{\rm dom}(f)$ and $A:=\text{\rm ran}(f)$. Then $U,A$ are internal so $U\in{}^{*}\mathcal{P}(\mathbb{N})$, and there is a transitive set $B\in\mathbb{U}$ such that $A\in{}^{*}B$. As ${}^{*}B$ is transitive also, we have $A\subseteq{}^{*}B$, and since $f$ is internal, we must have $f\in\mathcal{P}({}^{*}\mathbb{N}\times{}^{*}B)\cap{}^{*}\mathbb{U}={}^{*}\mathcal{P}(\mathbb{N}\times B)$. It follows that ${}^{*}\Psi(A,f,n)$ holds. Now observe that
$\mathbb{U}\vDash\forall X\in B\;\Big{(}\exists n\in\mathbb{N}\;\exists f\in\mathcal{P}(\mathbb{N}\times B)\;\Psi(X,f,n)\to X\in\mathcal{P}^{<\omega}(B)\Big{)}$
Applying transfer, with $X=A$, it follows that $A\in{}^{*}\mathcal{P}^{<\omega}(B)$.
()($\Leftarrow$)( ⇐ ): Conversely, suppose that $A\in{}^{*}\mathcal{P}^{<\omega}(B)$ for some $B\in\mathbb{U}$. Observe that
$\mathbb{U}\vDash\forall X\in\mathcal{P}^{<\omega}(B)\;\exists n\in\mathbb{N}\;\exists f\in\mathcal{P}(\mathbb{N}\times B)\;\Psi(X,f,n),$
and hence by transfer that there exist $n\in{}^{*}\mathbb{N}$ and $f\in{}^{*}\mathcal{P}(\mathbb{N}\times B)$ such that ${}^{*}\Psi(A,f,n)$. Then $f$ is an internal bijection from $\{0,1,\dots,n-1\}$ onto $A$.
$\dashv$
5 Enlargements and Saturation
5.1 Definitions and Basic Properties
Throughout this section, assume that $*:\mathbb{U}\to\mathbb{V}$ is a transfer map for a set $X$.
Recall Convention 1.11. Let $\mathcal{L}_{\mathbb{U}}$ denote the expansion of the langue $\mathcal{L}_{\in}$ (or some expansion thereof) with additional constant symbols for element of $\mathbb{U}$. Similarly, let $\mathcal{L}_{{}^{*}\mathbb{U}}$ denote the expansion with constant symbols for every internal set.
Recall also that a family $\mathcal{A}$ of sets has the finite intersection property (f.i.p.) if and only if the intersection of any finitely many members of $\mathcal{A}$ is non-empty.
Definition 5.1
Suppose that $\kappa$ is an infinite cardinal.
  1. (i)
    $*:\mathbb{U}\to\mathbb{V}$ is a $\kappa$-enlargement if and only if for every set $\Sigma(x)$ of $<\kappa$-many bounded formulas of $\mathcal{L}_{\mathbb{U}}$, if $\Sigma(x)$ is finitely satisfiable in $\mathbb{U}$ by elements of some set $T\in\mathbb{U}$, then $\Sigma(x)$ is satisfiable in $\mathbb{V}$ by an element of ${}^{*}\mathbb{U}$.
    * is an enlargement if it is a $|\mathbb{U}|^{+}$-enlargement, or, what is equivalent, if it is a $\kappa$-enlargement for any cardinal $\kappa$.
  2. (ii)
    $*:\mathbb{U}\to\mathbb{V}$ is $\kappa$-saturated if and only if for every set $\Sigma(x)$ of $<\kappa$-many bounded formulas of $\mathcal{L}_{{}^{*}\mathbb{U}}$, if $\Sigma(x)$ is finitely satisfiable in $\mathbb{V}$ by elements of some set $T\in{}^{*}\mathbb{U}$, then $\Sigma(x)$ is satisfiable in $\mathbb{V}$ by an element of ${}^{*}\mathbb{U}$.
    * is polysaturated if it is $|\mathbb{U}|^{+}$-saturated.
$\Box$
Theorem 5.2
  1. (a)
    $*:\mathbb{U}\to\mathbb{V}$ is a $\kappa$-enlargement if and only if whenever $\mathcal{A}\subseteq\mathbb{U}$ is a family of sets of cardinality $<\kappa$ such that $\mathcal{A}$ has the f.i.p., then $\bigcap{}^{\sigma}\mathcal{A}=\bigcap\{{}^{*}A:A\in\mathcal{A}\}$ is non-empty.
  2. (b)
    $*:\mathbb{U}\to\mathbb{V}$ is $\kappa$-saturated if and only if whenever $\mathcal{A}\subseteq{}^{*}\mathbb{U}$ is a family of internal sets of cardinality $<\kappa$ such that $\mathcal{A}$ has the f.i.p., then $\bigcap\mathcal{A}$ is non-empty.
Proof: (a) Suppose first that $*:\mathbb{U}\to\mathbb{V}$ is a $\kappa$-enlargement, and that $\mathcal{A}:=\{A_{\beta}:\beta<\alpha\}\subseteq\mathbb{U}$ is a family of sets of cardinality $<\kappa$ such that $\mathcal{A}$ has the f.i.p. By replacing each $A_{\beta}$ by $A_{\beta}\cap A_{0}$ we may without loss of generality assume that each $A_{\beta}\subseteq A_{0}$ — This does not affect the f.i.p. nor the intersection of all the $A_{\beta}$. Let $\Sigma(x):=\{\sigma_{\beta}(x):\beta<\alpha\}$, where $\sigma_{\beta}(x)\equiv x\in A_{\beta}$. Then $\Sigma(x)$ is a set of $<\kappa$-many bounded formulas of $\mathcal{L}_{\mathbb{U}}$ which is finitely satisfiable by elements of $A_{0}$. Hence $\Sigma(x)$ is satisfiable in $\mathbb{V}$ by an element of ${}^{*}\mathbb{U}$. Thus there is $a\in{}^{*}\mathbb{U}$ such that $a\in{}^{*}A_{\beta}$ for all $\beta<\alpha$, so that $\bigcap{}^{\sigma}\mathcal{A}\neq\varnothing$.
For the reverse direction, suppose we have a set $\Sigma(x)=\{\sigma_{\beta}(x):\beta<\alpha\}$ of $<\kappa$-many bounded formulas of $\mathcal{L}_{\mathbb{U}}$, and that $\Sigma(x)$ is finitely satisfiable in $\mathbb{U}$ by elements of some set $T\in\mathbb{U}$. Let $A_{\beta}:=\{t\in T:\mathbb{U}\vDash\sigma_{\beta}(t)\}$. As $\Sigma(x)$ is finitely satisfiable, the family $\mathcal{A}:=\{A_{\beta}:\beta<\alpha\}$ has the f.i.p. If $s\in\bigcap{}^{\sigma}\mathcal{A}$, then $s\in{}^{*}\mathbb{U}$. As $\mathbb{U}\vDash\forall x\in A_{\beta}(\sigma_{\beta}(x))$, we have $\mathbb{V}\vDash\forall x\in{}^{*}A_{\beta}(\sigma_{\beta}(x))$ so that $\mathbb{V}\vDash\Sigma(s)$.
(b) Suppose first that $*:\mathbb{U}\to\mathbb{V}$ is $\kappa$-saturated, and that $\mathcal{A}:=\{A_{\beta}:\beta<\alpha\}\subseteq{}^{*}\mathbb{U}$ is a family of internal sets of cardinality $<\kappa$ such that $\mathcal{A}$ has the f.i.p. By replacing each $A_{\beta}$ by $A_{\beta}\cap A_{0}$ we may without loss of generality assume that each $A_{\beta}\subseteq A_{0}$ — This does not affect the f.i.p. nor the intersection of all the $A_{\beta}$. Let $\Sigma(x):=\{\sigma_{\beta}(x):\beta<\alpha\}$, where $\sigma_{\beta}(x)\equiv x\in A_{\beta}$. Then $\Sigma(x)$ is a set of $<\kappa$-many bounded formulas of $\mathcal{L}_{{}^{*}\mathbb{U}}$ which is finitely satisfiable by elements of $A_{0}$. Hence $\Sigma(x)$ is satisfiable in $\mathbb{V}$ by an element of ${}^{*}\mathbb{U}$. Thus there is $a\in{}^{*}\mathbb{U}$ such that $a\in A_{\beta}$ for all $\beta<\alpha$, so that $\bigcap\mathcal{A}\neq\varnothing$.
For the reverse direction, suppose we have a set $\Sigma(x)=\{\sigma_{\beta}(x):\beta<\alpha\}$ of $<\kappa$-many bounded formulas of $\mathcal{L}_{{}^{*}\mathbb{U}}$, and that $\Sigma(x)$ is finitely satisfiable in $\mathbb{V}$ by elements of some set $T\in{}^{*}\mathbb{U}$. Let $A_{\beta}:=\{t\in T:\sigma_{\beta}(t)\}$. By the internal definition principle, each $A_{\beta}$ is internal. As $\Sigma(x)$ is finitely satisfiable by members of $T$, the family $\mathcal{A}:=\{A_{\beta}:\beta<\alpha\}$ has the f.i.p. If $s\in\bigcap\mathcal{A}$, then $s\in{}^{*}\mathbb{U}$, and $\mathbb{V}\vDash\Sigma(s)$.
$\dashv$
Note that we do not demand that $\mathcal{A}\in\mathbb{U}$ in (a), or that $\mathcal{A}\in\mathbb{V}$ in (b) of Theorem 5.2. However:
Theorem 5.3
To verify that a transfer map $*:\mathbb{U}\to\mathbb{V}$ for a set $X$ is a $\kappa$-enlargement, it suffices to consider sets $\mathcal{A}\in\mathbb{U}$ in (a) of of Theorem 5.2.
To verify that a transfer map $*:\mathbb{U}\to\mathbb{V}$ for a set $X$ is $\kappa$-saturated, it suffices to consider sets $\mathcal{A}$ which are subsets of standard sets — and thus in $\mathbb{V}$ — in (b) of of Theorem 5.2.
Proof: Suppose that * satisfies (i) of Definition 5.1 for sets $\mathcal{A}\in\mathbb{U}$. Let $\mathcal{A}^{\prime}\subseteq\mathbb{U}$ be a set of cardinality $<\kappa$ which satisfies the f.i.p. Choose $A_{0}\in\mathcal{A}^{\prime}$, and let $\mathcal{A}:=\{A\cap A_{0}:A\in\mathcal{A}^{\prime}\}$. Observe that $|\mathcal{A}|<\kappa$, that $\mathcal{A}$ satisfies the f.i.p. as well, and that since $\mathcal{A}\subseteq\mathcal{P}(A_{0})$ we have $\mathcal{A}\in\mathbb{U}$. By assumption, $\bigcap{}^{\sigma}\mathcal{A}\neq\varnothing$. Now as ${}^{*}(A\cap A_{0})\subseteq{}^{*}A$ for all $A\in\mathcal{A}^{\prime}$, we have $\varnothing\neq\bigcap{}^{\sigma}\mathcal{A}=\bigcap{}^{\sigma}\mathcal{A}^{\prime}$.
Next, suppose that * satisfies (ii) of Definition 5.1 for $\mathcal{B}\subseteq{}^{*}\mathcal{A}$, where $\mathcal{A}\in\mathbb{U}$. Let $\mathcal{B}^{\prime}\subseteq{}^{*}\mathbb{U}$ be a family of internal sets of cardinality $<\kappa$ which satisfies the f.i.p. Choose $B_{0}\in\mathcal{B}^{\prime}$, and let :=$\mathcal{B}:=\{B\cap B_{0}:B\in\mathcal{B}^{\prime}\}$. Observe that $|\mathcal{B}|<\kappa$, that $\mathcal{B}$ has the f.i.p. also, and that $\mathcal{B}\subseteq{}^{*}\mathbb{U}\cap\mathcal{P}(B_{0})$. By Lemma 3.8, $P:={}^{*}\mathbb{U}\cap\mathcal{P}(B_{0})$ is internal, and hence there is some transitive $\mathcal{A}\in\mathbb{U}$ such that $P\in{}^{*}\mathcal{A}$. Then also $\mathcal{B}\subseteq{}^{*}\mathcal{A}$, and hence by assumption we have that $\bigcap\mathcal{B}\neq 0$. Hence $\varnothing\neq\bigcap\mathcal{B}=\bigcap\mathcal{B}^{\prime}$.
$\dashv$
Remark 5.4
It follows from Theorem 5.10 that $*:\mathbb{U}\to\mathbb{V}$ is an enlargement if and only if it is a $|\mathbb{U}|$-enlargement, as $\mathcal{A}\in\mathbb{U}$ implies $|\mathcal{A}|<|\mathbb{U}|$.
$\Box$
Suppose that $B\in{}^{*}\mathbb{U}$ is an infinite internal set. Then for $b\in B$, the set $B_{b}:=\{c\in B:c\neq b\}$ $=B-\{b\}$ is internal, by the internal definition principle and the transitivity of ${}^{*}\mathbb{U}$. Clearly :=$\mathcal{B}:=\{B_{b}:b\in B\}$ has the f.i.p., yet $\bigcap\mathcal{B}=\varnothing$. Hence * cannot be a $|B|^{+}$-saturated extension. In particular, if $*:\mathbb{U}\to\mathbb{V}$ is a $\kappa$-saturated transfer map for an infinite set $X$ of atoms, then necessarily $|{}^{*}X|\geq\kappa$. It is therefore impossible to find an extension which is $\kappa$-saturated for every cardinal $\kappa$.
The above argument also shows that:
Proposition 5.5
If $*:\mathbb{U}\to\mathbb{V}$ is $\kappa$-saturated and $B\in\mathbb{V}$ has cardinality $|B|<\kappa$, then $B$ is external.
$\Box$
Lemma 5.6
If $*:\mathbb{U}\to\mathbb{V}$ is $\kappa$-saturated, it is a $\kappa$-enlargement. Hence a polysaturated extension is an enlargement.
Proof: If $\mathcal{A}\subseteq\mathcal{U}$ is a family of sets with the f.i.p., then $\mathcal{B}:={}^{\sigma}\mathcal{A}$ is a family of internal (indeed, standard) sets with the f.i.p. For, given $A_{1},\dots,A_{n}\in\mathcal{A}$, transfer of the bounded sentence $\exists x\in A_{1}(x\in A_{2}\land\dots\land x\in A_{n})$ shows that ${}^{*}A_{1}\cap\dots\cap{}^{*}A_{n}\neq\varnothing$. Clearly $|\mathcal{B}|=|\mathcal{A}|$.
$\dashv$
5.2 Enlargements, Concurrency and Hyperfinite Approximation
Theorem 5.7
A transfer map $*:\mathbb{U}\to\mathbb{V}$ for an infinite set $X$ is $\omega_{1}$-enlargement if and only if it is a nonstandard embedding.
Proof: ()($\Rightarrow$)( ⇒ ): Suppose that * is an $\omega_{1}$-enlargement. To show that * is a nonstandard embedding, it suffices to show that there is a countable $C\in\mathbb{U}$ such that ${}^{\sigma}C\subsetneqq{}^{*}C$. So let $C\subseteq X$ be countable. Since $\mathcal{P}(X)\in\mathbb{U}$ and $\mathbb{U}$ is transitive, we have $C\in\mathbb{U}$. Let $\mathcal{C}:=\{C-\{c\}:c\in C\}$. Then $\mathcal{C}$ has the f.i.p. and $|\mathcal{C}|<\omega_{1}$, so $\bigcap{}^{\sigma}\mathcal{C}\neq\varnothing$. Pick $c_{0}\in\bigcap{}^{\sigma}\mathcal{C}$. Then $c_{0}\in{}^{*}(C-\{c\})={}^{*}C-\{{}^{*}c\}$ for every $c\in C$. Hence $c_{0}\in{}^{*}C-{}^{\sigma}C$.
()($\Leftarrow$)( ⇐ ): Now assume that * is a nonstandard embedding. Let $\mathcal{A}\subseteq\mathbb{U}$ be a family of sets with the f.i.p. such that $|\mathcal{A}|<\omega_{1}$, i.e. $|\mathcal{A}|$ is countable. Let $\mathcal{A}=\{A_{1},A_{2},\dots\}$ be an enumeration of $\mathcal{A}$. Without loss of generality (by renaming elements if necessary), we may assume that $\mathbb{N}\subseteq X$. Let $f:\mathbb{N}\twoheadrightarrow\mathcal{P}(A_{1}):n\mapsto\bigcap_{k\leq n}A_{k}$. Then $f\in\mathbb{U}$, and $\varnothing\not\in\text{ran}f$. By transfer, $\varnothing\not\in\text{ran}\,{}^{*}f$, since ${}^{*}\varnothing=\varnothing$ and ${}^{*}(\text{ran}f)=\text{ran}\,{}^{*}f$. In particular, for all $n\in{}^{*}\mathbb{N}$, ${}^{*}f(n)\neq\varnothing$. Now fix an arbitrary $n\in\mathbb{N}$, and note that
$\mathbb{U}\vDash\forall m\in\mathbb{N}\;(m>n\to f(m)\subseteq A_{n}).$
By transfer
$\mathbb{V}\vDash\forall m\in{}^{*}\mathbb{N}\;(m>n\to{}^{*}f(m)\subseteq{}^{*}A_{n}).$
(Recall ${}^{*}x=x$ for all $x\in X$, and thus ${}^{*}n=n$.) Since * is a nonstandard embedding, it follows from Theorem 3.16 that there exists $m_{0}\in{}^{*}\mathbb{N}-{}^{\sigma}\mathbb{N}$ for which we have both ${}^{*}f(m_{0})\neq\varnothing$ and $m_{0}\geq n$. It follows that ${}^{*}f(m_{0})\subseteq{}^{*}A_{n}$. Since $n\in\mathbb{N}$ was arbitrary, we have that $\varnothing\neq f(m_{0})\subseteq\bigcap_{n\in\mathbb{N}}{}^{*}A_{n}=\bigcap{}^{\sigma}\mathcal{A}$. This shows that * is an $\omega_{1}$-enlargement.
$\dashv$
As a corollary we immediately see that:
Corollary 5.8
Every enlargement, and hence every polysaturated extension, is a nonstandard embedding.
$\Box$
Definition 5.9
A binary relation $R$ is concurrent if and only if for every finite $\{x_{1},\dots,x_{n}\}\subseteq\text{\rm dom}(R)$ there is $y\in\text{\rm ran}(R)$ such that $x_{i}Ry$ for all $i=1,\dots,n$.
$\Box$
For a set $A$, and important example of a concurrent relation is the set
$R:=\{(a,F):a\in F\in\mathcal{P}^{<\omega}(A)\}.$
Indeed, given $a_{1},\dots,a_{n}\in A=\text{\rm dom}(R)$, we can define $F:=\{a_{1},\dots,a_{n}\}$, and then observe that $a_{i}RF$ for all $i=1,\dots,n$.
Theorem 5.10
Suppose that $*:\mathbb{U}\to\mathbb{V}$ is a transfer map for a set $X$. Then the following are equivalent:
  1. (i)
    $*:\mathbb{U}\to\mathbb{V}$ is a $\kappa$-enlargement.
  2. (ii)
    For every concurrent binary relation $R\in\mathbb{U}$ with $|\text{\rm dom}(R)|<\kappa$ there is $y_{0}\in\mathbb{V}$ such that ${}^{*}x({}^{*}R)y_{0}$ for all $x\in\text{\rm dom}(R)$.
  3. (iii)
    (Hyperfinite Approximation) For each set $A\in\mathbb{U}$ with $|A|<\kappa$ there is a hyperfinite subset $B$ of ${}^{*}A$ which contains all the standard members of ${}^{*}A$, i.e.
    ${}^{\sigma}A\subseteq B\in{}^{*}\mathcal{P}^{<\omega}(A).$
Proof: (i) $\Rightarrow$ (ii): Suppose $|\text{\rm dom}(R)|<\kappa$. For $x\in\text{\rm dom}(R)$, let $A_{x}:=\{y\in\text{\rm ran}(R):xRy\}$, and let $\mathcal{A}:=\{A_{x}:x\in\text{\rm dom}(R)\}$. Then $\mathcal{A}\subseteq\mathcal{P}(\text{\rm ran}(R))$, so $\mathcal{A}\in\mathbb{U}$. Since $R$ is concurrent, $\mathcal{A}$ has the f.i.p. In addition $|\mathcal{A}|<\kappa$. Since * is a $\kappa$-enlargement, there is $y_{0}\in\bigcap{}^{\sigma}\mathcal{A}$. Then $y_{0}\in{}^{*}A_{x}$ for all $x\in\text{\rm dom}(R)$. Now $\mathbb{U}\vDash\forall y\in A_{x}\;(xRy)$, so by transfer we have $\mathbb{V}\vDash\forall y\in{}^{*}A_{x}\;({}^{*}x({}^{*}R)y)$. It follows that ${}^{*}x({}^{*}R)y_{0}$ for all $x\in\text{\rm dom}(R)$.
(ii)$\Rightarrow$ (iii): Given a set $A\in\mathbb{U}$ with $|A|<\kappa$, define the concurrent relation $R$ by
$R:=\{(a,F):a\in F\in\mathcal{P}^{<\omega}(A)\}.$
Then $|\text{\rm dom}(R)|<\kappa$, so by assumption, there is $B\in\mathbb{V}$ such that ${}^{*}a({}^{*}R)B$ — i.e. such that ${}^{*}a\in B$ — for all $a\in A$. Thus ${}^{\sigma}A\subseteq B$. Now as $B\in\text{\rm ran}({}^{*}R)$, and since $\text{\rm ran}(R)\subseteq\mathcal{P}^{<\omega}(A)$, we have that $\text{\rm ran}({}^{*}R)\subseteq{}^{*}\mathcal{P}^{<\omega}(A)$, so that $B\in{}^{*}\mathcal{P}^{<\omega}(A)$.
(iii) $\Rightarrow$ (i): Let $\mathcal{A}\in\mathbb{U}$ have the f.i.p. with $|\mathcal{A}|<\kappa$, and let $T\in\mathbb{U}$ be transitive so that $\mathcal{A}\subseteq T$. Then $A\subseteq T$ for all $A\in\mathcal{A}$, and hence
$\mathbb{U}\vDash\forall\mathcal{F}\in\mathcal{P}^{<\omega}(\mathcal{A})\;\exists x\in T\;\forall A\in\mathcal{F}\;(x\in A).$
By assumption, there is a hyperfinite $\mathcal{B}\subseteq{}^{*}\mathcal{A}$ such that ${}^{\sigma}\mathcal{A}\subseteq\mathcal{B}$. Then by transfer, with $\mathcal{F}=\mathcal{B}\in{}^{*}\mathcal{P}^{<\omega}(\mathcal{A})$, there is $x\in{}^{*}T$ such that $x\in A$ for all $A\in\mathcal{B}$. In particular, $x\in{}^{*}A$ for all $A\in\mathcal{A}$.
$\dashv$
5.3 Saturation and Concurrency
Theorem 5.11
Suppose that $*:\mathbb{U}\to\mathbb{V}$ is a nonstandard transfer map. Then the following are equivalent:
  1. (i)
    $*:\mathbb{U}\to\mathbb{V}$ is $\kappa$-saturated.
  2. (ii)
    For every internal concurrent binary relation $R\in\mathbb{V}$ with and every (internal or external) $A\subseteq\text{\rm dom}(R)$ with $|A|<\kappa$, there is $y_{0}\in\mathbb{V}$ such that $xRy_{0}$ for all $x\in A$.
Proof: (i) $\Rightarrow$ (ii): Suppose that $R\in\mathbb{V}$ is an internal concurrent relation, and that $A\subseteq\text{\rm dom}(R)$ is such that $|A|<\kappa$. For each $x\in A$, let $A_{x}:=\{y\in\text{\rm ran}(R):xRy\}$. By the internal definition principle, each $A_{x}$ is internal. Put $\mathcal{A}:=\{A_{x}:x\in A\}$, so that $\mathcal{A}$ is a family of internal sets with $|\mathcal{A}|<\kappa$. Since $R$ is concurrent, $\mathcal{A}$ has the f.i.p. By $\kappa$-saturation, there is $y_{0}\in\bigcap\mathcal{A}$. Then $y_{0}\in A_{x}$ for all $x\in A$, i.e. $xRy_{0}$ holds for all $x\in A$.
(ii) $\Rightarrow$ (i): Suppose that $\mathcal{A}$ is a family of internal sets with the f.i.p., where $|\mathcal{A}|<\kappa$. We must show that $\bigcap\mathcal{A}\neq\varnothing$. Fix $A_{0}\in\mathcal{A}$. By replacing each $A\in\mathcal{A}$ by $A\cap A_{0}$, we may assume that $A\subseteq A_{0}$ for each $A\in\mathcal{A}$. Thus $\mathcal{A}\subseteq\mathcal{P}(A_{0})\cap{}^{*}\mathbb{U}$. By Lemma 3.10, $P(A_{0})\cap{}^{*}\mathbb{U}$ is an internal set. Consider the relation
$R:=\{(A,a)\in(P(A_{0})\cap{}^{*}\mathbb{U})\times A_{0}:a\in A\}.$
By the internal definition principle, $R$ is internal. Now $\mathcal{A}\subseteq\text{\rm dom}(R)$, as each $A\in\mathcal{A}$ is non-empty (by the f.i.p.). Hence there is $a_{0}\in\mathbb{V}$ such that $(A,a_{0})\in R$ for all $A\in\mathcal{A}$. But then $a_{0}\in\bigcap\mathcal{A}$.
$\dashv$
5.4 Comprehensiveness
Definition 5.12
Let $\kappa$ be an infinite cardinal. A transfer map $*:\mathbb{U}\to\mathbb{V}$ is said to be $\kappa$-comprehensive if and only if for any sets $A,B\in\mathbb{U}$ such that $|A|<\kappa$, and any map $f:A\to{}^{*}B$, there is an internal function ${}^{+}\!f:{}^{*}A\to{}^{*}B$ with the property that ${}^{+}\!f({}^{*}a)=f(a)$ for all $a\in A$.
$*:\mathbb{U}\to\mathbb{V}$ is said to be comprehensive if and only if it is $\kappa$-comprehensive for every cardinal $\kappa$, or equivalently, if it is $|\mathbb{U}|^{+}$–comprehensive.
$*:\mathbb{U}\to\mathbb{V}$ is said to be countably comprehensive if and only if it is $\omega_{1}$-comprehensive.
$\Box$
Here is a small improvement:
Proposition 5.13
If a transfer map $*:\mathbb{U}\to\mathbb{V}$ is $\kappa$-comprehensive, then for any set $A\in\mathbb{U}$ of cardinality $|A|<\kappa$, and any internal set $B\in\mathbb{V}$, if $f:A\to B$, then there is an internal function ${}^{+}\!f:{}^{*}A\to B$ with the property that ${}^{+}\!f({}^{*}a)=f(a)$ for all $a\in A$.
Proof: Suppose that $f:A\to B$, where $A\in\mathbb{U}$ and that $B\in\mathbb{V}$ is internal. Let $T\in\mathbb{U}$ be transitive such that $B\in{}^{*}T$. Since ${}^{*}T$ is transitive, we have that $B\subseteq{}^{*}T$, so that $f:A\to{}^{*}T$. By definition of comprehensiveness, there is an internal function ${}^{+}\!g:{}^{*}A\to{}^{*}T$ such that ${}^{+}\!g({}^{*}a)=f(a)$ for all $a\in A$. Choose an arbitrary $b_{0}\in B$, and define
${}^{+}\!f:=\Big{\{}(a,b)\in{}^{*}A\times B:\exists c\in{}^{*}T\;\Big{(}(a,c)\in{}^{+}\!g\land(c\in B\to b=c)\land(c\in{}^{*}T-B\to b=b_{0})\Big{)}\Big{\}}.$
By Lemmas 2.1, 3.6 and the internal definition principle, ${}^{+}\!f$ is internal. Moreover, if $a\in{}^{*}A$, then if ${}^{+}\!g(a)\in B$, we have $(a,{}^{+}\!g(a))\in{}^{+}\!f$, whereas if ${}^{+}\!g(a)\in{}^{*}T-B$, then $(a,b_{0})\in{}^{+}\!f$. Thus ${}^{+}\!f$ is an internal function ${}^{+}\!f:{}^{*}A\to B$. Finally, if $a\in A$, then ${}^{+}\!g({}^{*}a)=f(a)\in B$, and hence $({}^{*}a,{}^{+}g({}^{*}a))\in{}^{+}\!f$, i.e. ${}^{+}\!f({}^{*}a)=f(a)$ when $a\in A$.
$\dashv$
The most useful application of comprehensiveness is in the countable case:
Corollary 5.14
Suppose that $*:\mathbb{U}\to\mathbb{V}$ is a countably comprehensive transfer map, where $\mathbb{U}$ is a universe over a set $X$ that contains $\mathbb{N}$. Suppose also that $B$ is internal and that $(b_{n}:n\in\mathbb{N})$ is a sequence of members of $B$. Then there is an internal sequence $(b_{n}:n\in{}^{*}\mathbb{N})$ which extends the sequence $(b_{n}:n\in\mathbb{N})$.
Proof: Define $f:\mathbb{N}\to B:n\mapsto b_{n}$, and let ${}^{+}\!f:{}^{*}\mathbb{N}\to B$ be the map provided by Proposition 5.13. Then since ${}^{*}n=n$ for all $n\in\mathbb{N}$, we have ${}^{+}\!f(n)={}^{+}\!f({}^{*}n)=f(n)=b_{n}$ for $n\in\mathbb{N}$. For $n\in{}^{*}\mathbb{N}-\mathbb{N}$, define $b_{n}:={}^{+}\!f(n)$.
$\dashv$
Theorem 5.15
Suppose that $*:\mathbb{U}\to\mathbb{V}$ is a $\kappa$-saturated nonstandard embedding. Then * is $\kappa$-comprehensive.
Proof: Suppose that * is $\kappa$-saturated. Let $A,B\in\mathbb{U}$, where $|A|<\kappa$, and suppose that $f:A\to{}^{*}B$.
$B_{a}:=\{g\in{}^{*}\mathcal{P}(A\times B):\text{$g$ is a function}\land\text{\rm dom}(g)={}^{*}A\land({}^{*}a,f(a))\in g\}.$
By the internal definition principle, each $B_{a}$ is internal.
Let :=$\mathcal{B}:=\{B_{A}:a\in A\}$, so that $|\mathcal{B}|<\kappa$. To apply $\kappa$-saturation, we must show that $\mathcal{B}$ has the f.i.p. Now given for each $n\in\mathbb{N}$, we have that
$\mathbb{U}\vDash\forall x_{1}\dots x_{n}\in B\;\exists g\in\mathcal{P}(A\times B)\;\Big{(}\text{$g$ is a function}\land\text{\rm dom}(g)=A\land(a_{1},x_{1})\in g\land\dots\land(a_{n},x_{n})\in g\Big{)}.$
Thus by transfer, setting $x_{i}:=f(a_{i})$ for $a_{1},\dots,a_{n}\in A$, there is $g\in{}^{*}\mathcal{P}(A\times B)$ such that
[…]
and then clearly $g\in\bigcap_{i\leq n}B_{a_{i}}$. As $n\in\mathbb{N}$ is arbitrary, it follows that $\mathcal{B}$ has the f.i.p.
By $\kappa$-saturation, there is an element ${}^{+}\!f\in\bigcap\mathcal{B}$. Then as ${}^{+}\!f\in{}^{*}\mathcal{P}(A\times B)$, it is an internal function ${}^{*}A\to{}^{*}B$ with the property that ${}^{+}\!f({}^{*}a)=f(a)$ for all $a\in A$.
$\dashv$
In the case of $\omega_{1}$-saturation, we also have the converse:
Theorem 5.16
Suppose that $*:\mathbb{U}\to\mathbb{V}$ is a nonstandard embedding, where $\mathbb{U}$ is a universe over a set $X$ that contains $\mathbb{N}$. The following are equivalent:
  1. (i)
    * is $\omega_{1}$-saturated, i.e. every countable family $\mathcal{A}$ of internal sets with the f.i.p. has non-empty intersection.
  2. (ii)
    * is countably comprehensive.
Proof: This follows directly from Theorem 5.15.
(ii) $\Rightarrow$ (i): Suppose that * is countably comprehensive, and that $\mathcal{A}:=\{A_{n}:n\in\mathbb{N}\}$ is a countable family of internal sets with the f.i.p. By replacing $A_{n}$ with $\bigcap_{m\leq n}A_{n}$, we may assume that the $A_{n}$ form a decreasing sequence of non-empty internal sets. As $A_{0}$ is internal, there is a transitive $T$ such that $A_{0}\in{}^{*}T$. Then as ${}^{*}T$ is transitive, we see that each $A_{n}$ is an internal subset of ${}^{*}T$, and thus $A_{n}\in{}^{*}\mathcal{P}(T)$. Thus we have a map $f:\mathbb{N}\to{}^{*}\mathcal{P}(T):n\to A_{n}$. By countable comprehensiveness, there is an internal map ${}^{+}\!f:{}^{*}\mathbb{N}\to{}^{*}\mathcal{P}(T)$ such that ${}^{+}\!f(n)=A_{n}$ for every $n\in\mathbb{N}$. For $n\in{}^{*}\mathbb{N}-\mathbb{N}$, define $A_{n}:={}^{+}\!f(n)$. Then $(A_{n}:n\in{}^{*}\mathbb{N})$ is an internal hypersequence of internal subsets of ${}^{*}T$ that extends the sequence $(A_{n}:n\in\mathbb{N})$. Let
$X:=\{n\in{}^{*}\mathbb{N}:\forall k\in{}^{*}\mathbb{N}\;(k\leq n\to{}^{+}\!f(k)\supseteq{}^{+}\!f(n))\land{}^{+}\!f(n)\neq\varnothing\}$
Then $X$ is internal, by the internal definition principle. By assumption, $\mathbb{N}\subseteq X\subseteq{}^{*}\mathbb{N}$. As $\mathbb{N}$ is not internal, there is $N\in{}^{*}\mathbb{N}-\mathbb{N}$ such that $N\in X$. Then $\bigcap_{n\in\mathbb{N}}A_{n}\supseteq A_{N}\neq\varnothing$. (One could also apply overflow to the set $X$ to deduce that $X$ has an infinite member.)
$\dashv$
6 Questions of Existence
This section is concerned with the existence of nonstandard embeddings, enlargements, comprehensive extensions, and polysaturated extensions. Since every polysaturated extension is both an enlargement (Lemma 5.6) and comprehensive (Theorem 5.15), the reader — if there is one — may want to read only Section 6.1.1 and either Section 6.4 (which in turn depends on the existence of good ultrafilters — cf. Appendix B) or Section 6.5.
6.1 Existence of Nonstandard Frameworks
Theorem 6.1
Let $V(X)$ be a superstructure over an infinite base set $X$. Then there exists a transfer map (i.e. a bounded elementary embedding) $V(X)\buildrel*\over{\to}V(Y)$ of $V(X)$ into some superstructure $V(Y)$ with base set $Y$ such that
  1. (i)
    ${}^{*}X=Y$.
  2. (ii)
    ${}^{*}\varnothing=\varnothing$.
  3. (iii)
    There is a countable $A\subseteq X$ such that ${}^{\sigma}A:=\{^{*}a:a\in A\}$ is a proper subset of ${}^{*}A$.
We will provide two proofs of this result, the first via ultrapowers, and the second by use of the Compactness Theorem of first-order logic.
6.1.1 Ultrapower Proof
We start from a superstructure $V(X)$ over a base set $X$. Let $\mathcal{U}$ be a countably incomplete ultrafilter over some index set $I$. The construction of the nonstandard framework proceeds over 9 steps. The first 8 steps will define a transfer map $*:V(X)\to V(Y)$ from the superstructure $V(X)$ to a superstructure $V(Y)$ over base set $Y$. These steps do not require $\mathcal{U}$ to be countably incomplete. In the $9^{\text{th}}$ step, the countable incompleteness of $\mathcal{U}$ is used to show that * induces a nonstandard framework.
Step 1: First we construct the base set $Y$: Define an equivalence relation $\sim_{\mathcal{U}}$ on $X^{I}$ by
$f\sim_{\mathcal{U}}g\qquad\text{ if and only if }\qquad\{i\in I:f(i)=g(i)\}\in\mathcal{U}.$
Define $Y$ to be the family of all equivalence classes
$Y:=X^{I}/\sim_{\mathcal{U}},$
i.e. $Y$ is just the ultrapower $X^{I}/\mathcal{U}$. We shall assume that $I$ is chosen so that $Y$ is a base set for $V(Y)$. This can be done as follows: Suppose that $X$ is a set of rank $\beta$, and let $I$ be a set of rank $\gamma\geq\beta+\omega$. Let $\mathcal{U}$ be an ultrafilter over $I$, and let $f:I\to X$. If $\gamma$ is a successor ordinal, $\gamma=\delta+1$, then $I$ has an element $i_{0}$ of rank $\delta$, but no elements of higher rank. It follows easily that the rank of $f$ is $\delta+3=\gamma+2$. On the other hand, if $\gamma$ is a limit ordinal, then $\sup_{i\in I}\text{rank}(i)=\gamma$, but the supremum is not attained. Hence $f$ has rank $\gamma$. It therefore follows that every $f:I\to X$ has the same rank, namely either $\gamma+2$ if $\gamma$ is a successor ordinal, or $\gamma$ if it is limit. Now an element of an element of $Y=X^{I}/\mathcal{U}$ is precisely a function $f:I\to X$. Thus all elements of elements of $Y$ have the same rank. The argument in Example 1.6 now shows that $Y$ is a base set, i.e. that $y\cap V(Y)=\varnothing$ for all $y\in Y$. Step 2: Now we define a structure $(W,=_{\mathcal{U}},\in_{\mathcal{U}})$ which forms part of an intermediate step in the definition of the transfer map $*:V(X)\to V(Y)$: Define binary relations $=_{\mathcal{U}}$ and $\in_{\mathcal{U}}$ on $V(X)^{I}$ as follows:
$f$ $=_{\mathcal{U}}g\quad\Longleftrightarrow\quad\{i\in I:f(i)=g(i)\}\in\mathcal{U},$
$f$ $\in_{\mathcal{U}}g\quad\Longleftrightarrow\quad\{i\in I:f(i)\in g(i)\}\in\mathcal{U}.$
If $f=_{\mathcal{U}}g$, we say that $f=g$ almost everywhere, and if $f\in_{\mathcal{U}}g$, we say that $f\in g$ almost everywhere.
We can associate with each $a\in V(X)$ the constant mapping $c_{a}:I\to V(X)$ which takes the value constant $a$. For each $n\in\mathbb{N}$, let
$W_{n}:=\{f\in V(X)^{I}:f\in_{\mathcal{U}}c_{V_{n}(X)}\},$
i.e. a function $f:I\to V(X)$ belongs to $W_{n}$ if and only if $f(i)\in V_{n}(X)$ for almost all $i$. Equivalently, $f\in W_{n}$ if and only if there is $g:I\to V_{n}(X)$ such that $f=_{\mathcal{U}}g$. Observe that since $V_{0}(X)=X$, we have that $W_{0}$ is essentially just $X^{I}$, i.e.
$W_{0}=\{f\in V(X)^{I}:f(i)\in X\text{ for almost all $i\in I$}\}.$
Clearly
$W_{0}\subseteq W_{1}\subseteq\dots\subseteq W_{n}\subseteq\dots$
Define
$W=\bigcup_{n}W_{n}.$
Step 3: We now show that there is a unique map $\cdot/\mathcal{U}:W\to V(Y)$ such that
  1. (i)
    $f/\mathcal{U}=\{g\in X^{I}:f=_{\mathcal{U}}g\}$ if $f\in W_{0}$, and
  2. (ii)
    $f/\mathcal{U}=\{g/\mathcal{U}:g\in W\land g\in_{\mathcal{U}}f\}$ if $f\in W-W_{0}$.
  3. (iii)
    $f/\mathcal{U}\in V_{n}(Y)$ whenever $f\in W_{n}$.
To begin with, define $f/\mathcal{U}:=\{g\in X^{I}:f=_{\mathcal{U}}g\}$ when $f\in W_{0}$. Then $f/\mathcal{U}=g/\sim_{\mathcal{U}}$ for any $g\in X^{I}$ such that $f=_{\mathcal{U}}g$. Since $g/\sim_{\mathcal{U}}\in X^{I}/\mathcal{U}=Y$, it follows that $f/\mathcal{U}\in V_{0}(Y)=Y$ when $f\in W_{0}$.
Now proceed by induction. Suppose we have shown that, for each $m\leq n$, there is a unique map $h_{m}:W_{m}\to V_{m}(Y)$ such that
$h_{m}(f)=f/\mathcal{U}\text{ if }f\in W_{0},\qquad h_{m}(f)=\{h_{m}(g):g\in W\land g\in_{\mathcal{U}}f\}\text{ if }f\in W_{m}-W_{0}.$
(Note that this condition makes sense, since if $f\in W_{m}$ and $g\in W$ is such that $g\in_{\mathcal{U}}f$, then
$\{g\in_{\mathcal{U}}c_{V_{m}(X)}\}\supseteq\{g\in_{\mathcal{U}}f\}\cap\{f\in_{\mathcal{U}}c_{V_{m}(X)}\}\in\mathcal{U},$
by transitivity of $V_{m}(X)$, so that $g\in W_{m}$ also.)
Define a map $h_{n+1}:W_{n+1}\to V_{n+1}(Y)$ as follows: First, for $f\in W_{n}$, define $h_{n+1}(f)=h_{n}(f)$, so that $h_{n+1}\restriction W_{n}=h_{n}$. Next, suppose that $f\in W_{n+1}-W_{n}$, and that $g\in W$ is such that $g\in_{\mathcal{U}}f$. Since also $f\in_{\mathcal{U}}c_{V_{n+1}(X)}$ (by definition of $W_{n+1}$), it is easy to see that $g\in_{\mathcal{U}}c_{V_{n}(X)}$, and hence that $g\in W_{n}$. Hence $h_{n+1}(g)=h_{n}(g)$ has already been defined, and so we may define $h_{n+1}(f):=\{h_{n+1}(g):g\in W,g\in_{\mathcal{U}}f\}$. Note that then $h_{n+1}(f)\subseteq V_{n}(Y)$, so indeed $h_{n+1}(f)\in V_{n+1}(Y)$.
Clearly, $h_{0}\subseteq h_{1}\subseteq h_{2}\subseteq\dots$. We therefore define the required map $\cdot/\mathcal{U}$ by $\cdot/\mathcal{U}:=\bigcup_{n}h_{n}$. Then clearly $\cdot/\mathcal{U}$ satisfies statements (i),(ii), (iii). Uniqueness is easily established as well, as any two maps satisfying (i), (ii) must agree on $W_{0}$, and then, by induction, on all $W_{n}$.
Step 4: Suppose that $f,g\in W$. We show that
  1. (iv)
    $g\in_{\mathcal{U}}f$ if and only if $g/\mathcal{U}\in f/\mathcal{U}$, and
  2. (v)
    $g=_{\mathcal{U}}f$ if and only if $g/\mathcal{U}=f/\mathcal{U}$
The definition of the map $\cdot/\mathcal{U}$ ensures that (i) holds (even for $W_{0}$, since $Y$ is a base set for $V(Y)$). Thus we need only prove (ii).
Since $f,g\in W$, there is $n\in\mathbb{N}$ such that $f,g\in W_{n}$. If $f,g\in W_{0}$, the statement of the Lemma is obviously true. Next, suppose that the statement holds for members of $W_{m}$ whenever $m\leq n$, and that $f,g\in W_{n+1}$.
If $f/\mathcal{U}=g/\mathcal{U}$ are such that $f\neq_{\mathcal{U}}g$, then $\{i\in I:f(i)=g(i)\}\not\in\mathcal{U}$, and so one of the sets $\{i\in I:f(i)-g(i)\neq\varnothing\},\{i\in I:g(i)-f(i)\neq\varnothing\}$ belongs to $\mathcal{U}$. Suppose the former. Define $h:I\to V(X)$ by letting $h(i)\in f(i)-g(i)$ if this set is non-empty, and setting $h(i)=\varnothing$ otherwise. Then $h\in W_{n}$. Clearly then $h/\mathcal{U}\in f/\mathcal{U}-g/\mathcal{U}$, so that $f/\mathcal{U}\neq g/\mathcal{U}$ — contradiction.
Conversely, suppose that $f,g\in W_{n+1}$ are such that $f=_{\mathcal{U}}g$. Then if $h/\mathcal{U}\in f/\mathcal{U}$, it follows that $h\in_{\mathcal{U}}f$, i.e. that $\{i\in I:h(i)\in f(i)\}\in\mathcal{U}$, and thus that $\{i\in I:h(i)\in g(i)\}\supseteq\{i\in I:h(i)\in f(i)\}\cap\{i\in I:f(i)=g(i)\}\in\mathcal{U}$. Hence also $h\in_{\mathcal{U}}g$, so $h/\mathcal{U}\in g/\mathcal{U}$. It follows that $f/\mathcal{U}\subseteq g/\mathcal{U}$. By symmetry $f/\mathcal{U}=g/\mathcal{U}$.
Step 5: Now we define the embedding ${}^{*}:V(X)\to V(Y)$ by
${}^{*}a:=c_{a}/\mathcal{U}.$
To decompose this definition, define $\iota:V(X)\to W:a\mapsto c_{a}$. Then if $a\in V_{n}(X)$, $\iota(a)=c_{a}\in W_{n}$. Then $*:V(X)\to V(Y)$ is just the composition
[…]
It is clear that * is an embedding, i.e that if $a\neq b$ belong to $V(X)$, then ${}^{*}a\neq{}^{*}b$.
Step 6: We show that ${}^{*}\varnothing=\varnothing$ and ${}^{*}X=Y$. The first statement is obvious. Observe that
${}^{*}X=c_{X}/\mathcal{U}=\{f/\mathcal{U}:f\in W\land f\in_{\mathcal{U}}c_{V_{0}(X)}\}=\{f/\mathcal{U}:f\in W_{0}\}=X^{I}/\mathcal{U}=Y.$
Step 7: We show that if $a\in V_{n}(X)$, then ${}^{*}a\in V_{n}(Y)$
This is clear if $n=0$, and thus holds for all individuals. We must therefore prove it for sets. Suppose now that it holds for $n$. If $A\in V_{n+1}(X)$ is a set, then ${}^{*}A=\{f/\mathcal{U}:f\in W\land f\in_{\mathcal{U}}c_{A}\}$. Thus if $f/\mathcal{U}\in{}^{*}A$, then $\{i\in I:f(i)\in A\}\in\mathcal{U}$, and thus $\{i\in I:f(i)\in V_{n}(X)\}\in\mathcal{U}$. It follows that $f\in W_{n}$, and thus that $f/\mathcal{U}\in V_{n}(Y)$ (by (iii) of Step 3). We therefore see that ${}^{*}A\subseteq V_{n}(Y)$, and thus that ${}^{*}A\in V_{n+1}(Y)$.
Step 8: Next, we prove that is a transfer map, i.e. a bounded elementary embedding. We must show that for every bounded $\mathcal{L}_{\in}$–formula $\varphi(x_{1},\dots,x_{n})$ and every $a_{1},\dots,a_{n}\in V(X)$, we have
$V(X)\vDash\varphi[a_{1},\dots,a_{n}]\quad\Longleftrightarrow\quad V(Y)\vDash\varphi[{}^{*}a_{1},\dots,{}^{*}a_{n}].$
We will first show that $f_{1},\dots,f_{n}\in W$, then
$V(Y)\vDash\varphi[f_{1}/\mathcal{U},\dots,f_{n}/\mathcal{U}]\quad\Longleftrightarrow\quad\{i\in I:V(X)\vDash\varphi[f_{1}(i),\dots,f_{n}(i)]\}\in\mathcal{U}.$
The proof is similar to that of Łoś’ Theorem, and proceeds by induction on the complexity of $\varphi$. If $\varphi$ is an atomic formula, i.e. of the form $\varphi(x_{1},x_{2})\equiv x_{1}\in x_{2}$ or $\varphi(x_{1},x_{2})\equiv x_{1}=x_{2}$, the result is an easy consequence of Step 4. In the former case, for example, $V(Y)\vDash f/\mathcal{U}\in g/\mathcal{U}$ if and only if $f\in_{\mathcal{U}}g$ if and only if $\{i\in I:V(X)\vDash f(i)\in g(i)\}\in\mathcal{U}$.
Now suppose that $\varphi\equiv\psi\land\chi$, and that the result has been proved for $\psi,\chi$. Then since $\mathcal{U}$ is closed under intersections and supersets, we have
$\phantom{\Longleftrightarrow}V(Y)\vDash\varphi[f_{1}/\mathcal{U},\dots,f_{n}/\mathcal{U}]$
$\Longleftrightarrow V(Y)\vDash\psi[f_{1}/\mathcal{U},\dots,f_{n}/\mathcal{U}]\quad\text{and}\quad V(Y)\vDash\chi[f_{1}/\mathcal{U},\dots,f_{n}/\mathcal{U}]$
$\Longleftrightarrow\{i\in I:V(X)\vDash\psi[f_{1}(i),\dots,f_{n}(i)]\}\in\mathcal{U}\quad\text{and}\quad\{i\in I:V(X)\vDash\chi[f_{1}(i),\dots,f_{n}(i)]\}\in\mathcal{U}$
$\Longleftrightarrow\{i\in I:V(X)\vDash\psi[f_{1}(i),\dots,f_{n}(i)]\}\cap\{i\in I:V(X)\vDash\chi[f_{1}(i),\dots,f_{n}(i)]\}\in\mathcal{U}$
$\Longleftrightarrow\{i\in I:V(X)\vDash\varphi[f_{1}(i),\dots,f_{n}(i)]\}\in\mathcal{U}$
Next, suppose that $\varphi\equiv\lnot\psi$, and that the result has been proved for $\psi$. Then since $\mathcal{U}$ is an ultrafilter, we have that, for every $A\subseteq I$, either $A\in\mathcal{U}$ or $A^{c}\in\mathcal{U}$. Hence
$\phantom{\Longleftrightarrow}V(Y)\vDash\varphi[f_{1}/\mathcal{U},\dots,f_{n}/\mathcal{U}]$
$\Longleftrightarrow V(Y)\not\vDash\psi[f_{1}/\mathcal{U},\dots,f_{n}/\mathcal{U}]$
$\Longleftrightarrow\{i\in I:V(X)\vDash\psi[f_{1}(i),\dots,f_{n}(i)]\}\not\in\mathcal{U}$
$\Longleftrightarrow\{i\in I:V(X)\vDash\lnot\psi[f_{1}(i),\dots,f_{n}(i)]\}\in\mathcal{U}$
Finally, if $\varphi\equiv(\forall y\in x_{1})\psi(y,x_{1},\dots,x_{n})$, where $y,x_{1},\dots,x_{n}$ are variables, then
$\phantom{\Longleftrightarrow}V(Y)\vDash\varphi[f_{1}/\mathcal{U},\dots,f_{n}/\mathcal{U}]$
$\Longleftrightarrow V(Y)\vDash\psi[g/\mathcal{U},f_{1}/\mathcal{U},\dots,f_{n}/\mathcal{U}]\quad\text{ for some }g/\mathcal{U}\in f_{1}/\mathcal{U}$
$\Longleftrightarrow\{i\in I:V(X)\vDash\psi[g(i),f_{1}(i),\dots,f_{n}(i)]\}\in\mathcal{U}\quad\text{ for some }g\in_{\mathcal{U}}f_{1}$
$\Longleftrightarrow\{i\in I:V(X)\vDash\varphi[f_{1}(i),\dots,f_{n}(i)]\}\in\mathcal{U}$
We have now proved ()($\ddagger$)( ‡ ). In particular, we have
$V(Y)\vDash\varphi[c_{a_{1}}/\mathcal{U},\dots,c_{a_{n}}/\mathcal{U}]\quad\Longleftrightarrow\quad\{i\in I:V(X)\vDash\varphi[c_{a_{1}}(i),\dots,c_{a_{n}}(i)]\}\in\mathcal{U},$
But since $c_{a}(i)=a$ for all $i\in I$, the set $\{i\in I:V(X)\vDash\varphi[c_{a_{1}}(i),\dots,c_{a_{n}}(i)]\}$ is either all of $I$, in which case $V(X)\vDash\varphi[a_{1},\dots,a_{n}]$, or it is $\varnothing$, in which case $V(X)\vDash\lnot\varphi[a_{1},\dots,a_{n}]$. This proves that ()($\dagger$)( † ), i.e. that $\cdot/\mathcal{U}$ is a transfer map for the language $\mathcal{L}_{\in}$.
Step 9: In order to complete the proof of Theorem 6.1, it remains to show that $*:V(X)\to V(Y)$ is a nonstandard framework over $X$, i.e. that there is a countable subset $A$ of $X$ such that ${}^{\sigma}A:=\{{}^{*}a:a\in A\}$ is a proper subset of ${}^{*}A$. This is the only place where we need the fact that the ultrafilter $\mathcal{U}$ is countably incomplete. In fact, we can directly prove:
If $A\in V(X)$ is an infinite set, then ${}^{\sigma}A\subsetneqq{}^{*}A$.
Since $\mathcal{U}$ is countably incomplete, we can partition $I$ into a countable sequence $I_{n}$ of sets, none of whom belong to $\mathcal{U}$. Suppose now that $A\in V(X)$ is an infinite set, with distinct elements $a_{0},a_{1},a_{2},\dots$. There is $m\in\mathbb{N}$ such that $A\in V_{m}(X)$. Define $f\in W_{m}$ by $f(i)=a_{n}$ whenever $i\in I_{n}$. Since ${}^{*}A=c_{A}/\mathcal{U}=\{g/\mathcal{U}:g\in W\land g\in_{\mathcal{U}}c_{A}\}$, we have that $f/\mathcal{U}\in{}^{*}A$. However, if $f/\mathcal{U}={}^{*}b$ for some $b\in A$, then $f/\mathcal{U}=c_{b}/\mathcal{U}$, so $\{i\in I:f(i)=b\}\in\mathcal{U}$, by Step 4. Since $f$ only takes the values $a_{n}$, it follows that $b=a_{n_{0}}$ for some $n_{0}\in\mathbb{N}$, But then $I_{n_{0}}=\{i\in I:f(i)=a_{n_{0}}\}\in\mathcal{U}$ — contradicting the fact that $I_{n}\not\in\mathcal{U}$ for all $n$. Hence $f/\mathcal{U}\in{}^{*}A-{}^{\sigma}A$.
$\dashv$
6.1.2 Compactness Theorem Proof
For this section, first recall the Compactness Theorem of first order logic — cf. Theorem A.17.
Definition 6.2
Let $\mathfrak{B}:=(B,E)$ be a model for $\mathcal{L}_{\in}$. A submodel $\mathfrak{A}$ of $\mathfrak{B}$ is said to be a transitive submodel if whenever $a\in A,b\in B$ and $bEa$, then $b\in A$.
$\Box$
Lemma 6.3
Suppose that $\mathfrak{A}=(A,E)$ is a transitive submodel of $\mathfrak{B}=(B,E)$. Then $\mathfrak{A}$ is a bounded elementary submodel of $\mathfrak{B}$.
Proof: We show, by induction on the complexity of $\varphi$, that for all $a_{1},\dots,a_{n}\in A$ we have
$\mathfrak{B}\vDash\varphi[a_{1},\dots,a_{n}]\quad\Leftrightarrow\quad\mathfrak{A}\vDash\varphi[a_{1},\dots,a_{n}].$
This is obvious for atomic formulas, and then easy to verify for the propositional connectives. Suppose therefore that $\mathfrak{B}\vDash\exists x\in a_{1}\psi[x,a_{1},\dots,a_{n}]$, where $a_{1},\dots,a_{n}\in A$, and $\psi$ is a bounded formula. Then there is $b\in B$ such that $bEa_{1}$ and $\mathfrak{B}\vDash\psi[b,a_{1},\dots,a_{n}]$. As $\mathfrak{A}$ is a transitive submodel, we have $b\in A$, so by induction we have $\mathfrak{A}\vDash\psi[b,a_{1},\dots,a_{n}]$, and hence $\mathfrak{A}\vDash\exists x\in a_{1}\psi[x,a_{1},\dots,a_{n}]$. As $\mathfrak{A}$ is a submodel of $\mathfrak{B}$, it is obvious that $\mathfrak{A}\vDash\exists x\in a_{1}\psi[x,a_{1},\dots,a_{n}]$ implies $\mathfrak{B}\vDash\exists x\in a_{1}\psi[x,a_{1},\dots,a_{n}]$.
$\dashv$
Given a model $\mathfrak{B}=(B,E)$, and an $X\in B$, we have that — from $\mathfrak{B}$’s point of view — the element $X$ is a base set if and only if
$\mathfrak{B}\vDash\text{BASE}[X],\quad\text{where}\quad\text{BASE}(x)\equiv\forall y\in x\;\forall z\in y\;(z\neq z),$
i.e. there are no $c,b\in B$ such that $cEbEX$: every $b\in B$ such that $bEX$ looks like an atom to $\mathfrak{B}$.
Given that $\mathfrak{B}\vDash\text{BASE}[X]$, we want to truncate $\mathfrak{B}$ by removing all sets which are not at a finite level over $X$. Recall the formulas $\nu_{n}(a,X)\equiv\varphi_{6,n}(X,a)$ which assert that $a\in V_{n}(X)$:
$\nu_{0}(y,x)\equiv y\in x,\qquad\nu_{n+1}(y,x)\equiv\nu_{n}(y,x)\lor\forall z\in y\;\nu_{n}(z,x).$
Thus for a superstructure $V(X)$ over a base set $X$, we have
$V_{n}(X)=\{a\in V(X):(V(X),\in)\vDash\nu_{n}[a,X]\},\qquad(V(X),\in)\vDash\text{BASE}[X].$
To define the truncation $\mathfrak{A}$ of $\mathfrak{B}$ over $X\in B$, we imitate: Define
$A:=\{a\in B:\text{ there is }n<\omega\text{ such that }\mathfrak{B}\vDash\nu_{n}[a,X]\},$
and let $\mathfrak{A}=(A,E)$ be the resulting submodel of $\mathfrak{B}$.
Lemma 6.4
Let $X$ be a base set, let $\mathfrak{B}=(B,E)$ be a bounded elementary extension of $(V(X),\in)$, and let $\mathfrak{A}$ be the truncation of $\mathfrak{B}$ over $X$. Then $\mathfrak{A}$ is a transitive submodel of $\mathfrak{B}$. Hence
$(V(X),\in)\preceq_{b}\mathfrak{A}\preceq_{b}\mathfrak{B}.$
In addition, the truncation of $\mathfrak{A}$ over $X$ is $\mathfrak{A}$ itself.
Proof: Observe that since $\text{BASE}(x)$ is a bounded formula and $V(X)\vDash\text{BASE}[X]$, we also have $\mathfrak{B}\vDash\text{BASE}[X]$. Now if $a\in A$, then there is a least $n<\omega$ such that $\mathfrak{B}\vDash\nu_{n}[a,X]$. If then $b\in B$ is such that $bEa$, then we cannot have $n=0$, because $\mathfrak{B}\vDash\text{BASE}[X]$, and hence necessarily $\mathfrak{B}\vDash\nu_{n-1}[b,X]$. It follows that $b\in A$, and thus that $\mathfrak{A}$ is a transitive submodel of $\mathfrak{B}$.
Furthermore, if $v\in V(X)$, then $V(X)\vDash\nu_{n}[v,X]$ for some $n$, As each $\nu_{n}(y,x)$ is a bounded formula, we have $\mathfrak{B}\vDash\nu_{n}[v,X]$, from which it follows that $v\in A$, and thus $(V(X),\in)$ is a transitive submodel of $\mathfrak{A}$.
By Lemma 6.3, we see that $(V(X),\in)\preceq_{b}(A,E)\preceq_{b}(B,E)$.
Finally if $a\in A$, then $\mathfrak{B}\vDash\nu_{n}[a,X]$ for some $n<\omega$. Since $\mathfrak{A}\preceq_{b}\mathfrak{B}$, we have $\mathfrak{A}\vDash\nu_{n}[a,X]$, from which it follows that $\mathfrak{A}$ is its own truncation over $X$.
$\dashv$
Theorem 6.5
(Mostowski Collapse) Let $\mathfrak{A}=(A,E)$ be a model of $\mathcal{L}_{\in}$ with an element $X\in A$ such that
  1. (i)
    $\mathfrak{A}\vDash\text{\rm BASE}[X]$.
  2. (ii)
    $\mathfrak{A}$ is its own truncation over $X$, i.e. for every $a\in A$ there is $n<\omega$ such that $\mathfrak{A}\vDash\nu_{n}[a,X]$.
  3. (iii)
    $\mathfrak{A}$ is extensional over $X$:
    $\mathfrak{A}\vDash\forall u\;\forall v\;(u\in x\lor v\in x\lor(u=v\leftrightarrow\forall z\;(z\in u\leftrightarrow z\in v))[X],$
    i.e. two sets relative to $\mathfrak{A}$ are equal if and only if ($\mathfrak{A}$ thinks that) they have the same elements.
  4. (iv)
    The set $Y:=\{a\in A:aEX\}$ is a base set.
Then there is a unique bounded elementary embedding $h:\mathfrak{A}\hookrightarrow(V(Y),\in)$ with the properties that:
  1. 1)
    $h(a)=a$ for all $a\in Y$,
  2. 2)
    $h(X)=Y$,
  3. 3)
    $\text{\rm ran}\;h$ is a transitive subset of $V(Y)$.
Proof: For $n<\omega$, let $A_{n}:=\{a\in A:\mathfrak{A}\vDash\nu_{n}[a,X]\}$. Then $(A_{n})_{n}$ is an increasing sequence of sets, and as $\mathfrak{A}$ is its own truncation over $X$, we have that $\bigcup_{n}A_{n}=A$. By definition of $\nu_{0}$, we have that $A_{0}=Y$. We now define $h\restriction A_{n}$ by induction, and then take $h:=\bigcup_{n}h\restriction A_{n}$.
For $a\in A_{0}=Y$, put $h(a)=a$.
Now suppose that $h\restriction A_{n}$ has been defined, and that $a\in A_{n+1}-A_{n}$. If $bEa$, then by definition of $\nu_{n+1}$ we must have $b\in A_{n}$, so that $h(b)$ is already defined. Thus put $h(a):=\{h(b):bEa\}$.
This completes the definition of $h$.
Now by definition of $h$ we have that $h(a)=a$ for all $a\in Y$. Furthermore,
$h(X)=\{h(a):aEX\}=\{h(a):a\in Y\}=\{a:a\in Y\}=Y.$
Next, we show that $\text{\rm ran}\;h$ is transitive. Suppose that $b^{\prime}\in a^{\prime}\in\text{\rm ran}\;h$. We must show that $b^{\prime}\in\text{\rm ran}\;h$. Now as $a^{\prime}\in\text{\rm ran}\;h$, there exists $a\in A$ such that $h(a)=a^{\prime}$, and so $b^{\prime}\in h(a)$. By definition of $h$, therefore, there must be $b\in A$ such that $bEa$ and $h(b)=b^{\prime}$. In particular, we see that $b^{\prime}\in\text{\rm ran}(h)$. This demonstrates that $\text{\rm ran}\;h$ is a transitive set.
For $n<\omega$, consider the statement:
$P_{n}\equiv\qquad\text{The restriction }h\restriction A_{n}\text{ is one-to one, and }h[A_{n}]\subseteq V_{n}(Y).$
It is clear that $P_{0}$ holds. Now suppose that $P_{n}$ holds, and that $a\in A_{n+1}$. Then $b\in A_{n}$ whenever $bEa$, and so $h(a)=\{h(b):bEa\}\subseteq V_{n}(Y)$, from which it follows that $h(a)\in V_{n+1}(Y)$, and thus that $h[A_{n+1}]\subseteq V_{n+1}(Y)$. Moreover, if $a,a^{\prime}\in A_{n+1}$ are such that $h(a)=h(a^{\prime})$. Then for every $bEa$ there is $cEa^{\prime}$ such that $h(b)=h(c)$. But necessarily $b,c\in A_{n}$, and as $P_{n}$ holds, we have that $h(b)=h(c)$ implies $b=c$. Thus $bEa$ if and only if $bEa^{\prime}$, and hence by the extensionality property we have $a=a^{\prime}$. It follows that $h\restriction A_{n+1}$ is one-to one, and thus that $P_{n+1}$ holds. Thus, by induction, $P_{n}$ holds for all $n<\omega$.
It follows that $\text{\rm ran}\;\subseteq V(Y)$ and that $h$ is one-to-one.
Next note that, by definition of $h$, we see that $bEa$ implies $h(b)\in h(a)$. Conversely, if $h(b)\in h(a)$, then $h(b)$ must be equal to $h(c)$ for some $cEa$. But as $h$ is one-to-one, we have that $b=c$. Hence $h(b)\in h(a)$ implies $bEa$. It follows that $h:(A,E)\hookrightarrow(V(Y),\in)$ is an embedding. As $(\text{\rm ran}\;h,\in)$ is a transitive submodel of $(V(Y),\in)$, it follows by Lemma 6.3 that
$(A,E)\cong(\text{\rm ran}\;h,\in)\preceq_{n}(V(Y),\in).$
In particular, $h:\mathfrak{A}\to(V(Y),\in)$ is a bounded elementary embedding.
It remains to show that $h$ is the unique bounded elementary embedding with the properties 1)-3). Suppose that $h^{\prime}$ is another such map. Again, we use induction to show that $h\restriction A_{n}=h^{\prime}\restriction A_{n}$. This is clear if $n=0$. Now suppose that $h\restriction A_{n}=h^{\prime}\restriction A_{n}$ and that $a\in A_{n+1}$. By definition of $h$ we see that $x\in h(a)$ implies $x=h(b)$ for some $bEa$. But then $h^{\prime}(b)\in h^{\prime}(a)$. As necessarily $b\in A_{n}$, we have $h(b)=h^{\prime}(b)$, and thus $x\in h^{\prime}(a)$.It follows that $h(a)\subseteq h^{\prime}(a)$. Conversely, if $x\in h^{\prime}(a)$, then as $\text{\rm ran}\;h^{\prime}$ is transitive, there is $b$ such that $h^{\prime}(b)=x$, i.e. $h^{\prime}(b)\in h^{\prime}(a)$. But as $h^{\prime}$ is a bounded elementary embedding, it follows that $bEa$, so that $b\in A_{n}$, and hence $h(b)=h^{\prime}(b)$. Now $bEa$ implies $h(b)\in h(a)$, and thus $x\in h(a)$. It follows that $h^{\prime}(a)\subseteq h(a)$, i.e. that $h^{\prime}(a)=h(a)$, and hence that $h\restriction A_{n+1}=h^{\prime}\restriction A_{n+1}$.
$\dashv$
We are now in a position to prove Theorem 6.1:
Proof: Suppose that $(V(X),\in)$ is a superstructure over an infinite base set $X$. Consider the language $\mathcal{L}_{V(X)}=\mathcal{L}\cup\{c_{u}:u\in V(X)\}$, and let $\Delta$ be the elementary diagram of $V(X)$ — cf. Definition A.8. For each infinite set $U\subseteq X$ in $V(X)$, let $d_{U}$ be a new constant symbol, and let $\Sigma=\Delta\cup\{\varphi_{U,u}:U\in V(X)\text{ an infinite subset of }X,u\in V(X)\}$, where
$\varphi_{U,u}\equiv d_{U}\in c_{U}\land d_{U}\neq c_{u}.$
If $\Sigma^{\prime}\subseteq\Sigma$ is finite, then it refers to at most finitely many $U,u$, and hence $(V(X),\in)$ can be expanded to a model $\Sigma^{\prime}$, where each $c_{u}$ is interpreted as the element $u\in V(X)$, and $d_{U}$ is interpreted to be a member of the set $U$. Hence $\Sigma$ is consistent, and therefore has a model $\mathfrak{B}=(B,E)_{c_{u},d_{U}}$. As this is a model of the elementary diagram, of $(V(X),\in)$ we see that we have an elementary extension $(V(X),\in)\preceq(B,E)$ — cf. Lemma A.11. Moreover, if $b_{U}\in B$ is the interpretation of the constant $d_{U}$, then we have $b_{U}EU$, but $b_{U}\neq u$ for any $u\in V(X)$, i.e. for every infinite $U\subseteq X$ there is $b\in B-V(X)$ such that $bEU$. By renaming, we may choose the set $B$ so that $Y:=\{b\in B:bEX\}$ is a base set.
Let $\mathfrak{A}$ be the truncation of $\mathfrak{B}$ over $X$. By Lemma 6.4 , we have $(V(X),\in)\preceq_{b}\mathfrak{A}\preceq_{b}\mathfrak{B}$, and $\mathfrak{A}$ is its on truncation over $X$. Furthermore, as $\text{BASE}(x)$ is a bounded formula and $V(X)\vDash\text{BASE}[X]$, it follows that $\mathfrak{A}\vDash\text{BASE}[X]$. In addition, $Y=\{b\in B:\nu_{0}[b,X]\}\subseteq A$, so $Y=\{a\in A:aEX\}$ is a base set.
Next, we show that $\mathfrak{A}$ is extensional over $X$. Suppose that $a,b\in A$ are such that $a,b\not\in X$ and that $a\neq b$. The model $(V(X),\in)$ is certainly extensional over $X$, and as it is an elementary submodel of $\mathfrak{B}$ (i.e. not merely a bounded elementary submodel), it follows that $\mathfrak{B}$ is extensional over $X$. Since $a\neq b$ are members of $B$, we have $\mathfrak{B}\vDash\exists z\>(z\in a\leftrightarrow z\not\in b)$. But this can also be written as a bounded formula: $\mathfrak{B}\vDash\exists z\in a(z\not\in b)\lor\exists z\in b(z\not\in a)$. Hence also $\mathfrak{A}\vDash\exists z\in a(z\not\in b)\lor\exists z\in b(z\not\in a)$, from which it follows that there is $c\in A$ such that $cEa$ if and only if $\lnot cEb$. This shows that $\mathfrak{A}$ is extensional over $X$.
We are now able to apply Theorem 6.5 to deduce that there is a bounded elementary extension $h:\mathfrak{A}\hookrightarrow(V(Y),\in)$. Let $*:=h\restriction V(X)$. Then $*:(V(X),\in)\hookrightarrow(V(Y),\in)$ is a composition of two bounded elementary embeddings, and thus a bounded elementary embedding, with ${}^{*}x=x$ for all $x\in X$, and ${}^{*}X=Y$. Finally, if $U\subseteq X$ is infinite, there is $b\in B-V(X)$ such that $bEU$. As $V(X)\vDash\forall u\in c_{U}\;(u\in c_{X})[U,X]$, we also have that $bEU$ implies $bEX$ for all $b\in B$, and hence $bEU$ implies $b\in A$, by definition of truncation. Thus $h(b)$ is defined, and $h(b)\in h(U)={}^{*}U$. The fact that $b\not\in X$ means that $h(b)\not\in\{h(x):x\in X\}=\{^{*}x:x\in X\}$, so $\{^{*}x:x\in U\}$ is a proper subset of ${}^{*}U$.
$\dashv$
6.2 Existence of Enlargements
Lemma 6.6
Suppose that $*:\mathbb{U}\to\mathbb{V}$ is a $\kappa$-enlargement. Then $|\mathbb{N}^{*}|\geq\sup\{|A|:A\in\mathbb{U},|A|<\kappa\}$.
Proof: Suppose that $A\in\mathbb{U}$ has cardinality $<\kappa$. By Theorem 5.10, there is a hyperfinite set $B$ such that ${}^{\sigma}A\subseteq B\subseteq{}^{*}A$. As $B$ is hyperfinite, there is a $n\in{}^{*}\mathbb{N}$ and a bijection $f:\{0,1,\dots,n-1\}\to B$, and thus an injection $h:B\hookrightarrow{}^{*}\mathbb{N}$. Thus ${}^{*}\mathbb{N}$ has a subset of cardinality $|B|$, and thus one of cardinality $|A|=|{}^{\sigma}A|\leq|B|$. Hence $|{}^{*}\mathbb{N}|\geq|A|$ for any $A\in\mathbb{U}$ with $|A|<\kappa$.
$\dashv$
Using the above lemma, it can be seen that the ultrapower construction does not automatically provide enlargements. For example, consider the ultrapower construction $*:V(X)\to V(Y)$ over an infinite base set $X$, with $Y=X^{I}/\sim_{\mathcal{U}}$ and $I=\mathbb{N}$. we may assume that $\mathbb{N}\subseteq X$. Then ${}^{*}\mathbb{N}$ is of the form $c_{\mathbb{N}}/\mathcal{U}=\{g/\mathcal{U}:g\in_{\mathcal{U}}\mathbb{N}\}$, and so $|{}^{*}\mathbb{N}|\leq|\mathbb{N}^{\mathbb{N}}|=2^{\aleph_{0}}$. Now $\mathbb{N},\mathcal{P}(\mathbb{N}),\mathcal{P}\mathcal{P}(\mathbb{N})\dots$ are all members of $\mathbb{U}$, so if $\kappa>(2^{\aleph_{0}})^{+}$, then $\sup\{|A|:A\in\mathbb{U},|A|<\kappa\}>2^{\aleph_{0}}$. Hence if $\kappa>(2^{\aleph_{0}})^{+}$, then * cannot be a $\kappa$-enlargement.
In order to use the ultrapower construction to obtain an enlargement, we have to be a bit more careful about the set $I$.
Theorem 6.7
[…]
Proof: Observe that if $a_{1},\dots,a_{n}\in I$, then $a_{1}\cup\dots\cup a_{n}$, being finite, is a member of $I_{a_{1}}\cap\dots\cap I_{a_{n}}$, which shows that the family $\{I_{a}:a\in I\}$ has the f.i.p. There is therefore an ultrafilter $\mathcal{U}$ over $I$ such that $I_{a}\in\mathcal{U}$ for every finite subset $a$ of $V(X)$.
Method 1: Suppose that $\mathcal{B}\in V(X)$ is a family of sets with the f.i.p. Define $f:I\to V(X)$ as follows: If $a\in I$, then $a\cap\mathcal{B}\in I$ also. If $a\cap\mathcal{B}\neq\varnothing$, choose $f(a)\in\bigcap(a\cap\mathcal{B})$; else, put $f(a)=\varnothing$. Observe that if $\mathcal{B}\in V_{n}(X)$, then $f(a)\in V_{n}(X)$, so that $f\in W_{n}$. We claim that $f/\mathcal{U}\in\bigcap{}^{\sigma}\mathcal{B}$.
For if $B\in\mathcal{B}$, then $\{B\}\in I$, and hence $I_{\{B\}}=\{a\in I:B\in a\}\in\mathcal{U}$. Now
$a\in I_{\{B\}}\Rightarrow B\in a\Rightarrow a\cap\mathcal{B}\neq\varnothing\Rightarrow f(a)\in\bigcap(a\cap\mathcal{B})\subseteq B,$
and hence $I_{\{B\}}\subseteq\{a\in I:f(a)\in B\}$. It follows that $f/\mathcal{U}\in c_{B}/\mathcal{U}={}^{*}B$. As $B\in\mathcal{B}$ is arbitrary, it follows that $f/\mathcal{U}\in\bigcap_{B\in\mathcal{B}}{}^{*}B=\bigcap{}^{\sigma}\mathcal{B}$. As $\mathcal{B}$ is arbitrary, it follows that $\bigcap{}^{\sigma}\mathcal{B}\neq\varnothing$ for all $\mathcal{B}\in V(X)$ with the f.i.p, and thus by Theorem 5.3, the extension $*:V(X)\to V(Y)$ is an enlargement.
Method 2: We use the hyperfinite approximation property: Let $B\in V(X)$, and define $g:I\to\mathcal{P}^{<\omega}(B)$ by $g(a):=a\cap B$. Then $\{a\in I:g(a)\in\mathcal{P}^{<\omega}(B)\}=I\in\mathcal{U}$, and thus $g/\mathcal{U}\in{}^{*}\mathcal{P}^{<\omega}(B)$, i.e. $g/\mathcal{U}$ is a hyperfinite subset of ${}^{*}B$. Now if $b\in B$ and $b\in a$, then $b\in a\cap B=g(a)$, and hence for $b\in B$ we have
$I_{\{b\}}:=\{a\in I:b\in a\}=\{a\in I:b\in g(a)\}.$
As $I_{\{b\}}\in\mathcal{U}$, we see that ${}^{*}b=c_{b}/\mathcal{U}\in g/\mathcal{U}$. Thus with $A:=g/\mathcal{U}\in{}^{*}\mathcal{P}^{<\omega}(B)$, we have ${}^{\sigma}B\subseteq A\subseteq{}^{*}B$. It follows that for every $B\in V(X)$ there is a hyperfinite set $A$ such that ${}^{\sigma}B\subseteq A\subseteq{}^{*}B$, and this is equivalent to * being an enlargement, by Theorem 5.10 .
$\dashv$
6.3 Existence of Comprehensive Transfer Maps
Every nonstandard embedding obtained from an ultrapower construction is comprehensive:
Theorem 6.8
Let $V(X)$ be a superstructure over $X$, and let $*:V(X)\to V({}^{*}X)$ be the transfer map provided by an ultrapower construction, as in Section 6.1.1. Then * is comprehensive.
Proof: Suppose that $V({}^{*}X)$ is obtained from $V(X)$ via an ultrafilter $\mathcal{U}$ over a set $I$. Suppose further that $A,B\in V(X)$, and that $f:A\to{}^{*}B$. We must show that there is an internal map ${}^{+}\!f:{}*A\to{}^{*}B$ with the property that ${}^{+}\!f({}^{a})=f(a)$ for all $a\in A$.
For $a\in V(X)$, let $c_{a}$ denote the constant map $c_{a}:I\to V(X):i\mapsto a$. Then ${}^{*}a:=c_{a}/\mathcal{U}$ — see Step 5 of the ultrapower proof of Theorem 6.1. Also, let $\rho_{f(a)}:I\to V(X)$ be such that $f(a)=\rho_{f(a)}/\mathcal{U}\in{}^{*}B$. As $V({}^{*}X)\vDash\rho_{f(a)}/\mathcal{U}\in c_{B}/\mathcal{U}$ we may, via Łos’ Theorem, and without loss of generality, assume that $\rho_{f(a)}(i)\in B$ for all $i\in I$. For $i\in I$, define $f_{i}:A\to B:a\mapsto\rho_{f(a)}(i)$. Now let $F:I\to V(X):i\mapsto f_{i}$.
Observe first that if $A,B\in V_{n}(X)$, then each $f_{i}\in V_{n+2}(X)$, so that $F\in V_{n+2}(X)^{I}$, i.e. $F\in W_{n+2}$ has finite rank, and so $F/\mathcal{U}\in V({}^{*}X)$. Since $V(X)\vDash F(i):c_{A}(i)\to c_{B}(i)$ holds for all $i\in I$, we see by Łos’ Theorem that $V({}^{*}X)\vDash F/\mathcal{U}:c_{A}/\mathcal{U}\to c_{B}/\mathcal{U}$, where we use Lemma 2.1(f). Moreover, $F/\mathcal{U}\in c_{V_{n+2}(X)}/\mathcal{U}={}^{*}V_{n+2}(X)$.
Thus if we define ${}^{+}\!f:=F/\mathcal{U}$, then we immediately see that ${}^{+}\!f:{}^{*}A\to{}^{*}B$ is internal.
Finally, ${}^{+}\!f({}^{*}a)=F/\mathcal{U}(c_{a}/\mathcal{U})$. Now in $V(X)$, we have that $F(i)(c_{a}(i))=f_{i}(a)=\rho_{f(a)}(i)$ for all $i\in I$. By Łos’ Theorem, therefore, we have that $F/\mathcal{U}(c_{a}/\mathcal{U})=\rho_{f(a)}/\mathcal{U}$ holds in $V({}^{*}X)$, i.e. that ${}^{+}\!f({}^{*}a)=f(a)$.
$\dashv$
Since only special types of ultrapower constructions provide enlargements, not every comprehensive extension is an enlargement.
In Section 5.3, we showed that countably comprehensive transfer maps are $\omega_{1}$-saturated, and vice versa. Example 6.19 will show that every ultrafilter is $\omega_{1}$–good, which, combined with Theorem 6.20, shows that every transfer map induced by an ultrapower construction is $\omega_{1}$-saturated, and thus countably comprehensive.
6.4 Existence of Polysaturated Extensions via Ultrapowers
Recall the following definitions from basic model theory:
Definition 6.9
  1. (a)
    Let $\mathcal{L}$ be a first-order language. Given a model $\mathfrak{A}$ of the language $\mathcal{L}$, we denote its universe $A$. If $X\subseteq A$, we denote by $\mathcal{L}_{X}$ the language $\mathcal{L}$ augmented with a set of new constant symbols $\{c_{a}:a\in X\}$. We expand the $\mathcal{L}$–structure $\mathfrak{A}$ to a $\mathcal{L}_{X}$-structure $(\mathfrak{A},a)_{a\in X}$, where the constant symbol $c_{a}$ is interpreted as the element $a$ in $(A,a)_{a\in X}$.
  2. (b)
    A set of formulas $\Sigma(x)$ in one free variable $x$ is said to be satisfiable in a model $\mathfrak{A}$ if and only if there is $b\in A$ such that $\mathfrak{A}\vDash\Sigma[b]$.
    $\Sigma(x)$ is said to be finitely satisfiable in $\mathfrak{A}$ if every finite subset of $\Sigma$ is satisfiable.
  3. (c)
    Suppose that $\mathfrak{A}$ is a model of the language $\mathcal{L}$, and that $\kappa$ be an infinite cardinal. $\mathfrak{A}$ is said to be $\kappa$-saturated if and only if and only if the following condition holds: Given a subset $X\subseteq A$ with $|X|<\kappa$ and a set of $\mathcal{L}_{X}$-formulas $\Sigma(x)$ in one free variable $x$, then $\Sigma(x)$ is satisfiable in $(\mathfrak{A},a)_{a\in X}$ whenever it is finitely satisfiable.
$\Box$
We first deal with a simple case:
Theorem 6.10
Suppose that $\mathcal{L}$ is a countable language, and that $\mathcal{U}$ is a countably incomplete ultrafilter over a set $I$. Then any ultraproduct $\prod_{I}\mathfrak{A}_{i}/\mathcal{U}$ is $\omega_{1}$-saturated
Proof: As $\mathcal{U}$ is countably incomplete, there is a decreasing chain $(I_{n})_{n\in\mathbb{N}}$ of elements of $\mathcal{U}$ such that $\bigcap_{n}I_{n}=\varnothing$. Without loss of generality, we may assume $I_{0}=I$.
We first show that the following claim holds: Claim: If $\mathcal{L}$ is a countable language, and if $\Sigma(x)$ is a set of $\mathcal{L}$-formulas which is finitely satisfiable in an ultraproduct $\prod_{I}\mathfrak{A}_{i}/\mathcal{U}$, then $\Sigma(x)$ is satisfiable. So suppose that $\Sigma(x)$ is finitely satisfiable in $\prod_{I}\mathfrak{A}_{i}/\mathcal{U}$. Since $\mathcal{L}$ is countable, so is $\Sigma(x)$, and hence we can enumerate it:
$\Sigma(x)=\{\sigma_{n}(x):n\geq 1\}.$
Define
$U_{0}:=I,\qquad U_{n}=I_{n}\cap\{i\in I:\mathfrak{A}_{i}\vDash\exists x\bigwedge_{1\leq m\leq n}\sigma_{m}(x)\},$
so that each $U_{n}\in\mathcal{U}$, by Łos’ Theorem. Now define $N(i):=\max\{n:i\in U_{n}\}$, and choose $a\in\prod_{I}A_{i}$ as follows: If $N(i)=0$, let $a(i)\in A_{i}$ be arbitrary. Else, choose $a(i)$ so that
$\mathfrak{A}_{i}\vDash\bigwedge_{1\leq m\leq N(i)}\sigma_{m}[a(i)].$
Now note that if $n\geq 1$ and $i\in U_{n}$, then $N(i)\geq n$, and hence $\mathfrak{A}_{i}\vDash\sigma_{n}[a(i)]$. It follows that
$U_{n}\subseteq\{i:\mathfrak{A}_{i}\vDash\sigma_{n}[a(i)]\},$
As $U_{n}\in\mathcal{U}$, it follows that $\prod_{I}\mathfrak{A}_{i}/\mathcal{U}\vDash\sigma_{n}[a/\mathcal{U}]$. As $n\geq 1$ is arbitrary, the element $a/\mathcal{U}$ satisfies $\Sigma(x)$ in $\prod_{I}\mathfrak{A}_{i}/\mathcal{U}$. This proves the Claim.
Now to prove $\omega_{1}$-saturation: Suppose that $X=\{a_{n}/\mathcal{U}:n\in\mathbb{N}\}$ is a countable set of elements of $\prod_{I}\mathfrak{A}_{i}/\mathcal{U}$. Let $\Sigma(x)$ be a set of formulas of $\mathcal{L}(X)$ which is finitely satisfiable in $(\prod_{I}\mathfrak{A}_{i}/\mathcal{U},(a_{n}/\mathcal{U})_{n})$. We must show that $\Sigma(x)$ is satisfiable in $(\prod_{I}\mathfrak{A}_{i}/\mathcal{U},(a_{n}/\mathcal{U})_{n})$. Now the expanded language $\mathcal{L}(X)$ is still countable, and it is easy to verify that
$(\prod_{I}\mathfrak{A}_{i}/\mathcal{U},(a_{n}/\mathcal{U})_{n})=\prod_{I}(\mathfrak{A}_{i},(a_{n}(i))_{n})/\mathcal{U},$
i.e. $(\prod_{I}\mathfrak{A}_{i}/\mathcal{U},(a_{n}/\mathcal{U})_{n})$ is an ultraproduct. By the Claim, in any ultraproduct modulo $\mathcal{U}$ of structures that interpret a countable language, every finitely satisfiable $\Sigma(x)$ is satisfiable. The result now follows by applying the claim to the ultraproduct $\prod_{I}(\mathfrak{A}_{i},(a_{n}(i))_{n})/\mathcal{U}$ of structures interpreting the countable language $\mathcal{L}(X)$.
$\dashv$
Suppose now that we have a family$\{\mathfrak{A}_{i}:i\in I\}$ of $\mathcal{L}$-structures, indexed by a set $I$, and an ultrafilter $\mathcal{U}$ on $I$. We seek conditions on $\mathcal{U}$ which will ensure that the ultraproduct $\prod_{I}A_{i}/\mathcal{U}$ is $\kappa$-saturated for $\kappa\geq\omega_{1}$. Moreover, as in Theorem 6.10, we seek a condition on $\mathcal{U}$, i.e. one which is independent of the models $\mathfrak{A}_{i}$.
Condition S: Whenever $\Sigma(x)$ with $|\Sigma(x)|<\kappa$ is finitely satisfiable in an ultraproduct modulo $\mathcal{U}$, then it is satisfiable.
This condition is independent of the language $\mathcal{L}$ or the models that make up the ultraproduct.
Observe that if $\mathcal{U}$ satisfies (S) then any ultraproduct modulo $\mathcal{U}$ interpreting a language $\mathcal{L}$ of cardinality $<\kappa$ is $\kappa$-saturated. To see this, suppose that $X\subseteq\prod_{I}A_{i}/\mathcal{U}$ has $|X|<\kappa$. Then $\mathcal{L}_{X}$ is also a language of cardinality $<\kappa$. Automatically, therefore, any set of $\mathcal{L}_{X}$-formulas $\Sigma(x)$ has $|\Sigma(x)|<\kappa$. Condition (S) then immediately yields that any set of $\mathcal{L}_{X}$-formulas $\Sigma(x)$ which is finitely satisfiable in $\prod_{I}\mathfrak{A}_{i}/\mathcal{U}$ is satisfiable in $\prod_{I}\mathfrak{A}_{i}/\mathcal{U}$.
We seek a property of $\mathcal{U}$ which guarantees that (S) holds.
Thus let $\Sigma(x)$ be a set of formulas of cardinality $<\kappa$ which is finitely satisfiable in $\prod_{I}\mathfrak{A}_{i}/\mathcal{U}$. Then if $\Theta$ is a finite subset of $\Sigma$, the set $\{i\in I:\mathfrak{A}_{i}\vDash\exists x(\bigwedge\Theta)\}$ belongs to $\mathcal{U}$, by Łos’ Theorem. We thus have a map
$p:\mathcal{P}^{<\omega}(\Sigma)\to\mathcal{U}:\Theta\mapsto\{i\in I:\mathfrak{A}_{i}\vDash\exists x(\bigwedge\Theta)\}.$
Thus $i\,\in p(\Theta)$ if and only if $\mathfrak{A}_{i}\vDash\exists x\bigwedge\Theta$, i.e. if and only if $\Theta$ is satisfiable in $\mathfrak{A}_{i}$.
It should be clear that $\Theta\subseteq\Theta^{\prime}\Rightarrow p(\Theta)\supseteq p(\Theta^{\prime})$. Equivalently
$p(\Theta\cup\Theta^{\prime})\subseteq p(\Theta)\cap(\Theta^{\prime})$
Lemma 6.11
If there exists a sequence $(\Phi_{i})_{i\in I}$ in $\mathcal{P}^{<\omega}(\Sigma)$ such that
  1. (i)
    $i\in p(\Phi_{i})$ for all $i\in I$, and
  2. (ii)
    for all $\theta\in\Sigma$, $\{i\in I:\theta\in\Phi_{i}\}\in\mathcal{U}$,
then $\Sigma$ is satisfiable in $\prod_{I}\mathfrak{A}_{i}/\mathcal{U}$.
Proof: $p(\Theta)$ is just the set of all $i\in I$ for which $\Theta$ is satisfiable in $\mathfrak{A}_{i}$. Hence by if $i\in p(\Phi_{i})$, then there is $a_{i}\in A_{i}$ such that $\mathfrak{A}_{i}\vDash\bigwedge\Phi_{i}(a_{i})$. By (i), there is such an $a_{i}\in A_{i}$ for every $i\in I$. Thus by (ii),
$\{i\in I:\mathfrak{A}_{i}\vDash\theta(a_{i})\}\supseteq\{i\in I:\theta\in\Phi_{i}\}\in\mathcal{U}\qquad\text{for every }\theta\in\Sigma,$
so that by Łos’ Theorem, $\prod_{I}\mathfrak{A}_{i}/\mathcal{U}\vDash\theta[(a_{i})_{i\in I}/\mathcal{U}]$ for every $\theta\in\Sigma$. Thus the element $(a_{i})_{i\in I}/\mathcal{U}$ satisfies $\Sigma$ in $\prod_{I}\mathfrak{A}_{i}/\mathcal{U}$.
$\dashv$
We say that a sequence $(\Phi_{i})_{i\in I}$ of members of $\mathcal{P}^{<\omega}(\Sigma)$ supports $p$ if it satisfies (i), (ii) of Lemma 6.11. Thus if there is a sequence supporting the relation $p$ defined in ()($\star$)( ⋆ ) above, then $\Sigma$ is satisfiable in $\prod_{i\in I}\mathfrak{A}_{i}/\mathcal{U}$.
Now the above observations do not in any way depend on the linguistic aspects of $\mathcal{L},\Sigma$. We can therefore fruitfully move to a slightly more abstract realm:
Definition 6.12
  1. (a)
    Let $X$ be a set. An order-reversal is a map $p:\mathcal{P}^{<\omega}(X)\to\mathcal{U}$ such that $\Theta\subseteq\Theta^{\prime}\Rightarrow p(\Theta)\supseteq p(\Theta^{\prime})$, or equivalently
    $p(\Theta\cup\Theta^{\prime})\subseteq p(\Theta)\cap p(\Theta^{\prime}).$
    (Recall that $\mathcal{P}^{<\omega}(X)$ denotes the family of finite subsets of $X$.)
  2. (b)
    A sequence $(\Phi_{i})_{i\in I}$ of members of $\mathcal{P}^{<\omega}(X)$ supports the order-reversal $p:\mathcal{P}^{<\omega}(X)\to\mathcal{U}$ if and only if
    1. (i)
      $i\in p(\Phi_{i})$ for all $i\in I$.
    2. (ii)
      For all $\theta\in X$, $\{i\in I:\theta\in\Phi_{i}\}\in\mathcal{U}$, or equivalently, for all $\Theta\in\mathcal{P}^{<\omega}(X)$, $\{i\in I:\Theta\subseteq\Phi_{i}\}\in\mathcal{U}$.
  3. (c)
    An order-reversal $p$ is anti-additive if also
    $p(\Theta\cup\Theta^{\prime})=p(\Theta)\cap p(\Theta^{\prime}).$
  4. (d)
    An order-reversal $p$ is locally finite if for all $i\in I$
    $\sup\{|\Theta|:\Theta\in\mathcal{P}^{<\omega}(X),i\in p(\Theta)\}<\infty.$
    This means that for all $i\in I$ there is $N_{i}\in\mathbb{N}$ such that $i\not\in\Theta$ whenever $|\Theta|>N_{i}$.
$\Box$
Observe that the order-reversal defined in ()($\star$)( ⋆ ) need not be either anti-additive or locally finite. However, if an abstract order-reversal $p$ has a support $\Phi:=(\Phi_{i})_{i\in I}$, then we can define another order-reversal $p_{\Phi}$ which has those properties:
Lemma 6.13
Suppose that $p:\mathcal{P}^{<\omega}(X)\to\mathcal{U}$ is an order-reversal possessing a support $\Phi:=(\Phi_{i})_{i\in I}$. Define a map
$p_{\Phi}:\mathcal{P}^{<\omega}(X)\to\mathcal{U}:\Theta\mapsto\{i\in I:\Theta\subseteq\Phi_{i}\}.$
Then $p_{\Phi}$ is a anti-additive locally finite order-reversal with the properties that (i) $p_{\Phi}\leq p$, and (ii) $\Phi$ is a support for $p_{\Phi}$ also.
Proof: Clearly if $\Theta\subseteq\Theta^{\prime}$ belong to $\mathcal{P}^{<\omega}(X)$, then $p_{\Phi}(\Theta)=\{i:\Theta\subseteq\Phi_{i}\}\supseteq\{i:\Theta^{\prime}\subseteq\Phi_{i}\}=p_{\Phi}(\Theta^{\prime})$. Hence $p_{\Phi}$ is an order-reversal. Moreover, since $\Phi_{i}\subseteq\Phi_{i}$, we have $i\in p_{\Phi}(\Phi_{i})$ for all $i\in I$, and hence $p_{\Phi}$ is supported by $(\Phi_{i})_{i\in I}$.
Note also that if $i\in p_{\Phi}(\Theta)$, then $\Theta\subseteq\Phi_{i}$, and hence $p(\Theta)\supseteq p(\Phi_{i})$. Since also $i\in p(\Phi_{i})$, we see that $i\in p(\Theta)$ whenever $i\in p_{\Phi}(\Theta)$, i.e. $p_{\Phi}(\Theta)\subseteq p(\Theta)$ for all $\Theta\in\mathcal{P}^{<\omega}(X)$.
It remains to show that $p_{\Phi}$ is anti-additive and locally finite. To prove that it is anti-additive, observe that
$p_{\Phi}(\Theta\cup\Theta^{\prime})=\{i:\Theta\cup\Theta^{\prime}\subseteq\Phi_{i}\}=\{i:\Theta\subseteq\Phi_{i}\}\cap\{i:\Theta^{\prime}\subseteq\Phi_{i}\}=p_{\Phi}(\Theta)\cap p_{\Phi}(\Theta^{\prime}).$
Next, to prove that $p_{\Phi}$ is locally finite, observe that if $i\in p_{\Phi}(\Theta)$, then $\Theta\subseteq\Phi_{i}$. Hence $\sup\{|\Theta|:i\in p_{\Phi}(\Theta)\}=|\Phi_{i}|<\infty$.
$\dashv$
Lemma 6.14
Suppose that $p:\mathcal{P}^{<\omega}(X)\to\mathcal{U}$ is a anti-additive and locally finite order-reversal. Then $p$ has a support $\Phi$ such that $p=p_{\Phi}$.
Proof: Since $p$ is anti-additive, the set $\{\Theta\in\mathcal{P}^{<\omega}(X):i\in p(\Theta)\}$ is closed under finite unions, for each $i\in I$. Thus, as $p$ is locally finite, the set $\{\Theta\in\mathcal{P}^{<\omega}(X):i\in p(\Theta)\}$ has a maximum element, namely $\Phi_{i}:=\bigcup\{\Theta:i\in p(\Theta)\}$. Then certainly $i\in p(\Phi_{i})$. Now observe that
$i\in p(\Theta)\Rightarrow\Theta\subseteq\Phi_{i}\Rightarrow p(\Theta)\supseteq p(\Phi_{i})\Rightarrow i\in p(\Theta),$
i.e. all the above are equivalent. In particular, for any $\Theta\in\mathcal{P}^{<\omega}(X)$ we have
$\{i:\Theta\subseteq\Phi_{i}\}=\{i:i\in p(\Theta)\}=p(\Theta)\in\mathcal{U},$
and hence $(\Phi_{i})_{i\in I}$ is a support of $p$.
Finally,
$p_{\Phi}(\Theta):=\{i:\Theta\subseteq\Phi_{i}\}=p(\Theta),$
i.e. $p_{\Phi}=p$.
$\dashv$
We now seek conditions that will ensure that any order-reversal on any set $X$ of cardinality $<\kappa$ has a support. Observe that if $p,p^{\prime}$ are order-reversals so that $p^{\prime}\leq p$, then $p^{\prime}$ is locally finite if $p$ is, since $\{\Theta:i\in p^{\prime}(\Theta)\}\subseteq\{\Theta:i\in p(\Theta)\}$ for all $i\in I$. Furthermore, if $\Phi$ is a support of $p^{\prime}$, then it is a support of $p$ also, since $I=\{i:i\in p^{\prime}(\Phi_{i})\}\subseteq\{i:i\in p(\Phi_{i})\}$. Hence if we can define for every order-reversal $p$ two smaller order -eversals $Lp\leq p,Cp\leq p$ so that $Lp$ is locally finite, and $Cp$ is anti-additive, then $CLp\leq Lp\leq p$ is anti-additive and locally finite. Hence $CLp$ has a support, and this will be a support of $p$ also.
According to Lemma 6.11, if the ultrafilter $\mathcal{U}$ has the property that any order-reversal $p:\mathcal{P}^{<\omega}(X)\to\mathcal{U}$ (where $|X|<\kappa$) has a support, then any ultraproduct modulo $\mathcal{U}$ interpreting a language with $|\mathcal{L}|<\kappa$ is $\kappa$-saturated.
So our aim is to find conditions on an ultrafilter which guarantee the existence of operators $C,L$.
We can easily deal with the operator $L$:
Lemma 6.15
If $\mathcal{U}$ is a countably incomplete ultrafilter on a set $I$, then $L$ exists, i.e. for every order-reversal $p:\mathcal{P}^{<\omega}(X)\to\mathcal{U}$, there is a locally finite order-reversal $q\leq p$.
Proof: Since $\mathcal{U}$ is countably incomplete, there exists a decreasing sequence $(I_{n})_{n\in\mathbb{N}}$ of members of $\mathcal{U}$ such that $\bigcap_{n}I_{n}=\varnothing$. Without loss of generality, we may take $I_{0}=I$.
For each $i\in I$, define $N(i):=\min\{n:i\not\in I_{n}\}$. As the $I_{n}$ form a decreasing sequence, we have $I_{n}=\{i:N(i)>n\}$.
Now suppose that we have an order-reversal $p:\mathcal{P}^{<\omega}(X)\to\mathcal{U}$. Define $Lp$ by
[…]
$i\in(Lp)(\Theta)\quad\text{iff}\quad i\in p(\Theta)\land|\Theta|<N(i).$
It is then easy to see that $Lp$ is an order-reversal and that $Lp\leq p$. Furthermore, $\sup\{|\Theta|:i\in(Lp)(\Theta)\}<N(i)<\infty$ for each $i\in I$, so $Lp$ is locally finite.
$\dashv$
To tackle the existence of $C$, we introduce the following definition:
Definition 6.16
Let $\kappa$ be a cardinal. An ultrafilter $\mathcal{U}$ is $\kappa$-good if for every cardinal $\alpha<\kappa$ and every order-reversal $p:\mathcal{P}^{<\omega}(\alpha)\to\mathcal{U}$, $Cp$ exists, i.e. there is a anti-additive order-reversal $q:\mathcal{P}^{<\omega}(\alpha)\to\mathcal{U}$ such that $q\leq p$.
$\Box$
To recapitulate:
  • Suppose that $\mathcal{U}$ is a countably incomplete $\kappa$-good ultrafilter on a set $I$, and that $(\mathfrak{A}_{i})_{i\in I}$ is a family of models interpreting a language $\mathcal{L}$ of cardinality $<\kappa$.
  • Let $X\subseteq\prod_{I}\mathfrak{A}_{i}/\mathcal{U}$ be such that $|X|<\kappa$, and suppose that $\Sigma(x)$ is a family of formulas of $\mathcal{L}_{X}$ that is finitely satisfiable in $\prod_{I}\mathfrak{A}_{i}/\mathcal{U}$.
  • Then $|\Sigma(x)|<\kappa$.
  • If $\Theta\in\mathcal{P}^{<\omega}(\Sigma)$, then $\{i\in I:\mathfrak{A}_{i}\vDash\exists x\bigwedge\Theta\}\in\mathcal{U}$, since $\Theta$ is satisfiable in $\prod_{I}\mathfrak{A}_{i}/\mathcal{U}$. Thus the map
    $p:\mathcal{P}^{<\omega}(\Sigma)\to\mathcal{U}:\Theta\mapsto\{i\in I:\mathfrak{A}_{i}\vDash\exists x\bigwedge\Theta\}$
    is an order-reversal.
  • By Lemma 6.11, the set $\Sigma$ is satisfiable in $\prod_{I}\mathfrak{A}_{i}/\mathcal{U}$ when $p$ has a support.
  • Lemma 6.14 shows that every locally finite consistent order-reversal has a support.
  • Since $\mathcal{U}$ is countably incomplete, Lemma 6.15 shows that there is a locally finite $p^{\prime}\leq p$.
  • The fact that $\mathcal{U}$ is $\kappa$-good then yields the existence of a $p^{\prime\prime}\leq p^{\prime}$ which is consistent. That $p^{\prime\prime}$ will therefore also be locally finite, because $p^{\prime}$ is. Hence $p^{\prime\prime}$ has a support $\Phi$. Then as $p^{\prime\prime}\leq p$, $\Phi$ will also be a support for $p$.
  • Hence $p$ has a support, and thus $\Sigma$ is satisfiable in $\prod_{I}\mathfrak{A}_{i}/\mathcal{U}$.
  • Since $X$, $\Sigma(x)$ were arbitrary, every finitely satisfiable family of formulas $\Sigma(x)$ in the expanded language $\mathcal{L}_{X}$ is satisfiable in $\prod_{I}\mathfrak{A}_{i}/\mathcal{U}$ (for $|X|<\kappa$). Thus $\prod_{I}\mathfrak{A}_{i}/\mathcal{U}$ is $\kappa$-saturated.
Thus we have shown:
Theorem 6.17
If $\mathcal{U}$ is a countably incomplete $\kappa$-good ultrafilter, then every ultraproduct modulo $\mathcal{U}$ interpreting a language of cardinality $<\kappa$ is $\kappa$-saturated.
$\Box$
It remains to address the existence of good ultrafilters. The following theorem requires some heavy-duty combinatorics, so its proof has been relegated to the appendix.
Theorem 6.18
Suppose that $\kappa$ is an infinite cardinal. Then there exists a countably incomplete $\kappa^{+}$-good ultrafilter over $\kappa$.
$\Box$
Example 6.19
For the case $\kappa^{+}=\omega_{1}$, is quite easy to show that there are $\omega_{1}$–good countably incomplete ultrafilters, for the simple reason that every ultrafilter is $\omega_{1}$-good. Indeed, suppose that $\mathcal{U}$ is an ultrafilter over a set $I$, and that $p:\mathcal{P}^{<\omega}(\omega)\to\mathcal{U}$ is an order-reversal. For $s\in\mathcal{P}^{<\omega}(\omega)$, define
$q(s):=p\{m\in\omega:m\leq\max s\}.$
Then as $s\subseteq\{m\in\omega:m\leq\max s\}$, we have $q(s)\subseteq p(s)$. Moreover, since for $s,t\in\mathcal{P}^{<\omega}(\omega)$ we have
$\{m:m\leq\max(s\cup t)\}=\left\{\begin{aligned} \{m:m\leq\max s\}\quad&\text{if }\max s\geq\max t,\\ \{m:m\leq\max t\}\quad&\text{else },\end{aligned}\right.$
it follows that $q(s\cup t)=q(s)\cap q(t)$, i.e. that $q$ is anti-additive.
Hence Theorem 6.10 also follows from Theorem 6.18.
$\Box$
6.4.1 Construction of Polysaturated Extensions via Good Ultrapowers
Recall the ultrapower construction of a nonstandard extension $V(X)\overset{*}{\hookrightarrow}V(^{*}X)$: We start with a base set $X$ and an ultrafilter $\mathcal{U}$ over a set $I$, where $I$ is chosen so that the ultrapower $Y:=X^{I}/\mathcal{U}$ is another base set. We define relations $=_{\mathcal{U}},\in_{\mathcal{U}}$ on $V(X)^{I}$ by
$f=_{\mathcal{U}}g\Leftrightarrow\{f=g\}\in\mathcal{U},\qquad f\in_{\mathcal{U}}g\Leftrightarrow\{f\in g\}\in\mathcal{U}.$
For $n\in\mathbb{N}$, we define
$W_{n}:=\{f\in V(X)^{I}:f\in_{\mathcal{U}}c_{V_{n}(X)}\},\quad\text{ and then }W:=\bigcup_{n\in\mathbb{N}}W_{n},$
where for $a\in V(X)$ the map $c_{a}:I\to V(X)$ is the constant map with value $a$. Observe that if $a\in V(X)$, then $a\in V_{n}(X)$ for some $n\in\mathbb{N}$, and hence $c_{a}\in W_{n}$. Thus there is a natural inclusion
$\iota:V(X)\hookrightarrow W:a\mapsto c_{a}.$
By induction, we construct a map $\cdot/\mathcal{U}:W\to V(Y)$, as follows: For $f\in W_{0}$, define
$f/\mathcal{U}:=\{g\in X^{I}:f=_{\mathcal{U}}g\},$
and for $f\in W-W_{0}$, define
$f/\mathcal{U}:=\{g/\mathcal{U}:g\in W\land g\in_{\mathcal{U}}f\}.$
The map $\cdot/\mathcal{U}$ has the property that $f/\mathcal{U}\in V_{n}(Y)$ whenever $f\in W_{n}$.
[…]
${}^{*}V(X)=\{f/\mathcal{U}:f\in W\}.$
Theorem 6.20
Let $\kappa$ be a cardinal, and suppose that a nonstandard framework $V(X)\overset{*}{\hookrightarrow}V({}^{*}X)$ is obtained as an ultrapower construction via a $\kappa$-good countably incomplete ultrafilter $\mathcal{U}$ over a set $I$. Then $V({}^{*}X)$ is $\kappa$-saturated for then language $\mathcal{L}_{{}^{*}V(X)}$, i.e. if $\mathcal{A}$ is a family of internal sets with the f.i.p. such that $|\mathcal{A}|<\kappa$, then $\bigcap\mathcal{A}\neq\varnothing$.
Proof: Let $\Gamma<\kappa$ be an ordinal, and suppose that $\mathcal{A}:=\{A_{\gamma}:\gamma<\Gamma\}$ is a family of internal sets with the f.i.p. We must show that $\bigcap\mathcal{A}\neq\varnothing$. Without loss of generality, by replacing $A_{\gamma}$ by $A_{\gamma}\cap A_{0}$, we may assume that $A_{\gamma}\subseteq A_{0}$ for all $\gamma<\Gamma$: This affects neither the f.i.p. nor the value of $\bigcap\mathcal{A}$.
As each $A_{\gamma}$ is internal, there is a function $a_{\gamma}\in W$ such that $A_{\gamma}=a_{\gamma}/\mathcal{U}$.
As $\mathcal{U}$ is countably incomplete, there is a sequence $I=I_{1}\supseteq I_{2}\supseteq I_{3}\supseteq\dots$ of members of $\mathcal{U}$ such that $\bigcap_{n=1}^{\infty}I_{n}=\varnothing.$ Now define a map
$f:\mathcal{P}^{<\omega}(\Gamma)\to\mathcal{U}:\Delta\mapsto I_{n}\cap\Big{\{}i\in I:\bigcap_{\gamma\in\Delta}a_{\gamma}(i)\neq\varnothing\Big{\}}.$
(Note that $\bigcap_{\gamma\in\Delta}A_{\gamma}\neq\varnothing$ for any finite $\Delta\subseteq\Gamma$, by the f.i.p., so the set $\Big{\{}i\in I:\bigcap_{\gamma\in\Delta}a_{\gamma}(i)\neq\varnothing\Big{\}}$ belongs to $\mathcal{U}$.) Observe that $f$ is a reversal. Since $\Gamma<\kappa$ and $\mathcal{U}$ is $\kappa$-good, there is a strict reversal $g:\mathcal{P}^{<\omega}(\Gamma)\to\mathcal{U}$ such that $g\leq f$.
For $i\in I$, define
$\Gamma_{i}:=\{\gamma\in\Gamma:i\in g(\{\gamma\})\},$
so that $\gamma\in\Gamma_{i}$ if and only if $i\in g(\{\gamma\})$. We first show that each $\gamma_{i}$ is a finite set. Fix $i\in I$. Since $\bigcap_{n=1}^{\infty}I_{n}=\varnothing$, there is $n\in\mathbb{N}$ such that $i\not\in I_{n}$. We claim that $|\Gamma_{i}|<n$. Indeed, if $\gamma_{1},\dots,\gamma_{n}$ are distinct elements of $\Gamma_{i}$, then
$i\in\bigcap_{m=1}^{n}g(\{\gamma_{m}\})=g(\{\gamma_{1},\dots,\gamma_{n}\})\subseteq f(\{\gamma_{1},\dots,\gamma_{n}\})\subseteq I_{n},$
which is impossible, as $i\not\in I_{n}$. Thus each $\Gamma_{i}\in\mathcal{P}^{<\omega}(\Gamma)$, so that $g(\Gamma_{i})$ is defined.
Now observe that $g(\Gamma_{i})=g(\bigcup_{\gamma\in\Gamma_{i}}\{\gamma\})=\bigcap_{\gamma\in\Gamma_{i}}g(\{\gamma\})$. As $i\in g(\{\gamma\})$ whenever $\gamma\in\Gamma_{i}$, we see that $i\in g(\Gamma_{i})$. As $f(\Gamma_{i})\supseteq g(\Gamma_{i})$, we have $i\in f(\Gamma_{i})$, so that $\bigcap_{\gamma\in\Gamma_{i}}a_{\gamma}(i)\neq\varnothing$, by definition of $f$.
Now choose a map $x\in W$ so that $x(i)\in\bigcap_{\gamma\in\Gamma_{i}}a_{\gamma}(i)$. (Recall that $A_{\gamma}\subseteq A_{0}$, and that $A_{0}=a_{0}/\mathcal{U}$ for some $a_{0}\in W$. Then $a_{0}\in W_{n}$ for some $n$, and hence we may take $x\in W_{n-1}\subseteq W$.)
Now if $\gamma\in\Gamma_{i}$, then $x(i)\in a_{\gamma}(i)$, by definition of $x$. Hence if $\gamma\in\Gamma$, then
$\{i\in I:x(i)\in a_{\gamma}(i)\}\supseteq\{i\in I:\gamma\in\Gamma_{i}\}=g(\{\gamma\})\in\mathcal{U},$
from which we see that $x/\mathcal{U}\in a_{\gamma}/\mathcal{U}=A_{\gamma}$. As $\gamma\in\Gamma$ was arbitrary, we have $x/\mathcal{U}\in\bigcap\mathcal{A}$.
$\dashv$
6.5 Existence of Polysaturated Extensions via Ultralimits
6.5.1 Limits of Chains of Superstructures
Suppose that $\lambda$ is a limit ordinal, and that $\{X^{\alpha}:\alpha<\lambda\}$ is a collection of base sets. Suppose further that, for $\alpha\leq\beta<\lambda$, we have a chain of superstructures $(V(X^{\alpha})))_{\alpha<\lambda}$ linked by bounded elementary embeddings $V(X^{\alpha})\overset{\iota_{\alpha\beta}}{\to}V(X^{\beta})$ with the following properties:
  1. (i)
    $\iota_{\alpha\alpha}=\text{id}_{V(X^{\alpha})}$.
  2. (ii)
    If $\alpha\leq\beta\leq\gamma<\lambda$, then $\iota_{\beta\gamma}\circ\iota_{\alpha\beta}=\iota_{\alpha\gamma}$.
  3. (iii)
    If $\alpha\leq\beta<\lambda$, then $\iota_{\alpha\beta}(X^{\alpha})=X^{\beta}$.
We now want to construct a limit model $V(X^{\lambda})$ and bounded elementary embeddings
$V(X^{\alpha})\overset{\iota_{\alpha\lambda}}{\to}V(X^{\lambda})$ so that properties (i)-(iii) hold for $\alpha\leq\beta\leq\lambda$, i.e. for $\lambda$ as well.
Lemma 6.21
The bounded elementary embeddings $\iota_{\alpha\beta}$ are rank-preserving, i.e. if $\alpha\leq\beta$ and $n\in\mathbb{N}$, then $x\in V_{n}(X^{\alpha})$ if and only if $\iota_{\alpha\beta}(x)\in V_{n}(X^{\beta})$.
Proof: By Lemma 2.1 there is a bounded formula $\varphi_{6,n}(X,x)$ such that $V(X)\vDash\varphi_{6,n}(X,x)$ if and only if $x\in V_{n}(X)$. Thus $x\in V_{n}(X^{\alpha})$ if and only if $V(X^{\alpha})\vDash\varphi_{6,n}(X^{\alpha},x)$ if and only if $V(X^{\beta})\vDash\varphi_{6,n}(\iota_{\alpha\beta}(X^{\alpha}),\iota_{\alpha\beta}(x))$ if and only if $\iota_{\alpha\beta}(x)\in V_{n}(X^{\beta})$, using the fact that $\iota_{\alpha\beta}(X^{\alpha})=X^{\beta}$.
$\dashv$
We now proceed to make a limit model out of the models $V(X^{\alpha})$. For $n\in\mathbb{N}$, define
$P_{n}:=\{(a,\alpha):\alpha<\lambda\land a\in V_{n}(X^{\alpha})\},\qquad P:=\bigcup_{n\in\mathbb{N}}P_{n}.$
Observe that $P_{0}\subseteq P_{1}\subseteq P_{2}\subseteq\dots$.
Define binary relations $\sim,E$ on $P$ as follows:
$(a,\alpha)\sim(b,\beta)\Leftrightarrow\exists\gamma\geq\alpha,\beta\;\Big{(}\iota_{\alpha\gamma}(a)=\iota_{\beta\gamma}(b)\Big{)},\qquad(a,\alpha)E(b,\beta)\Leftrightarrow\exists\gamma\geq\alpha,\beta\;\Big{(}\iota_{\alpha\gamma}(a)\in\iota_{\beta\gamma}(b)\Big{)}.$
Lemma 6.22
  1. (a)
    $\sim$ is an equivalence relation.
  2. (b)
    have that
    $(a,\alpha)\sim(b,\beta)\Leftrightarrow\forall\delta\geq\alpha,\beta\;\Big{(}\iota_{\alpha\delta}(a)=\iota_{\beta\delta}(b)\Big{)},$
    and that
    $(a,\alpha)E(b,\beta)\Leftrightarrow\forall\delta\geq\alpha,\beta\;\Big{(}\iota_{\alpha\delta}(a)\in\iota_{\beta\delta}(b)\Big{)}.$
  3. (c)
    If $\alpha\leq\beta$, then $(a,\alpha)\sim(b,\beta)$ if and only if $\iota_{\alpha\beta}(a)=b$, and $(a,\alpha)E(b,\beta)$ if and only if $\iota_{\alpha\beta}(a)\in b$.
  4. (d)
    $(a,\alpha)\sim(b,\beta)$ if and only if $\forall(c,\gamma)\in P\;\Big{(}(c,\gamma)E(a,\alpha)\leftrightarrow(c,\gamma)E(b,\beta)\Big{)}$.
Proof: (a) It is clear that $\sim$ is reflexive and symmetric. If $(a,\alpha)\sim(b,\beta)\sim(c,\gamma)$, then there are $\eta\geq\alpha,\beta$ and $\xi\geq\beta,\gamma$ such that $\iota_{\alpha\eta}(a)=\iota_{\beta\eta}(b)$ and $\iota_{\beta\xi}(b)=\iota_{\gamma\xi}(c)$. Let $\delta\geq\eta,\xi$. Then
$\iota_{\alpha\delta}(a)=\iota_{\eta\delta}\circ\iota_{\alpha\eta}(a)=\iota_{\eta\delta}\circ\iota_{\beta\eta}(b)=\iota_{\beta\delta}(b)=\iota_{\xi\delta}\circ\iota_{\beta\xi}(b)=\iota_{\xi\delta}\circ\iota_{\gamma\xi}(c)=\iota_{\gamma\delta}(c),$
which shows that $(a,\alpha)\sim(c,\gamma)$, establishing transitivity.
(b) Suppose that $(a,\alpha)\sim(b,\beta)$ and that $\gamma\geq\alpha,\beta$ is such that $\iota_{\alpha\gamma}(a)=\iota_{\beta\gamma}(b)$. Let $\delta\geq\alpha,\beta$. If $\delta\geq\gamma$, then
$\iota_{\alpha\delta}(a)=\iota_{\gamma\delta}\circ\iota_{\alpha\gamma}(a)=\iota_{\gamma\delta}\circ\iota_{\beta\gamma}(b)=\iota_{\beta\delta}(b).$
If, on the other hand, $\alpha,\beta\leq\delta\leq\gamma$, then
$i_{\delta\gamma}\circ\iota_{\alpha\delta}(a)=\iota_{\alpha\gamma}(a)=\iota_{\beta\gamma}(b)=\iota_{\delta\gamma}\circ\iota_{\beta\delta}(b).$
Then since $\iota_{\delta\gamma}$ is one-to-one, we conclude that $\iota_{\alpha\delta}(a)=\iota_{\beta\delta}(b)$ in this case also.
The proof for $E$ is similar.
(c) follows directly from (b): Since $\beta\geq\alpha,\beta$, it follows that $\iota_{\alpha\beta}(a)=\iota_{\beta\beta}(b)=b$.
The proof for $E$ is similar.
(d) Suppose that $(a,\alpha)\sim(b,\beta)$ and that $(c,\gamma)E(a,\alpha)$. Applying (b), we see that if $\delta\geq\alpha,\beta,\gamma$, then $\iota_{\gamma\delta}(c)\in\iota_{\alpha\delta}(a)=\iota_{\beta\delta}(b)$, so that $(c,\gamma)E(b,\beta)$. By symmetry, we see that if $(a,\alpha)\sim(b,\beta)$, then $(c,\gamma)E(a,\alpha)$ if and only if $(c,\gamma)E(b,\beta)$.
Conversely, suppose that for all $(c,\gamma)\in P$ we have $(c,\gamma)E(a,\alpha)$ if and only if $(c,\gamma)E(b,\beta)$. Let $\gamma\geq\alpha,\beta$. If $(a,\alpha)\not\sim(b,\beta)$, then $\iota_{\alpha\gamma}(a)\neq\iota_{\beta\gamma}(b)$, so there is $c\in(\iota_{\alpha\gamma}(a)-\iota_{\beta\gamma}(b))\cup(\iota_{\beta\gamma}(b)-\iota_{\alpha\gamma}(a))$. Without loss of generality, suppose that $c\in(\iota_{\alpha\gamma}(a)-\iota_{\beta\gamma}(b))$. Then by (c),
$c\in\iota_{\alpha\gamma}(a)\Rightarrow(c,\gamma)E(a,\alpha)\Rightarrow(c,\gamma)E(b,\beta)\Rightarrow c\in\iota_{\beta\gamma}(b),$
contradiction. Hence $(a,\alpha)\sim(b,\beta)$.
$\dashv$
Lemma 6.23
  1. (a)
    If $(a,\alpha)\in P_{n}$ and $(b,\beta)\sim(a,\alpha)$, then $(b,\beta)\in P_{n}$.
  2. (b)
    If $(a,\alpha)\in P_{n}$ and $(b,\beta)E(a,\alpha)$, then $(b,\beta)\in P_{n-1}$.
Proof: (a) Suppose that $(a,\alpha)\in P_{n}$ and $(b,\beta)\sim(a,\alpha)$. If $\gamma\geq\alpha,\beta$, then $\iota_{\alpha\gamma}(a)=\iota_{\beta\gamma}(b)$. But by the rank-preserving property of the $\iota_{\alpha\beta}$, we have
$(a,\alpha)\in P_{n}\Rightarrow a\in V_{n}(X^{\alpha})\Rightarrow\iota_{\alpha\gamma}(a)\in V(X^{\gamma})\Rightarrow\iota_{\beta\gamma}(b)\in V(X^{\gamma})\Rightarrow b\in V_{n}(X^{\beta})\Rightarrow(b,\beta)\in P_{n}.$
(b) First observe that if $(a,\alpha)\in P_{0}$, and $(b,\beta)E(a,\alpha)$, then $b\in\iota_{0\beta}(a)\in V_{0}(X^{\beta})=X^{\beta}$, i.e. $\iota_{0\beta}(a)$ is a member of the base set $X^{\beta}$ which has an element $b\in V(X^{\beta})$ — contradicting the definition of base set. Hence if $(a,\alpha)\in P_{0}$ there can be no $(b,\beta)$ such that $(b,\beta)E(a,\alpha)$. Thus if $(b,\beta)E(a,\alpha)$, then $n\geq 1$. Then if $\gamma\geq\alpha,\beta$, it follows by the rank-preserving properties of the $\iota_{\alpha\beta}$ that
$\iota_{\beta\gamma}(b)\in\iota_{\alpha\gamma}(a)\in V_{n}(X^{\gamma})\Rightarrow\iota_{\beta\gamma}(b)\in V_{n-1}(X^{\gamma})\Rightarrow b\in V_{n-1}(X^{\beta})\Rightarrow(b,\beta)\in P_{n-1}.$
$\dashv$
Now define a sequence of sets $(W_{n})_{n\in\mathbb{N}}$ by induction, as follows: First, if $(a,\alpha)\in P_{0}$, let $[a,\alpha]:=\{(b,\beta):(b,\beta)\sim(a,\alpha)\}$. By Lemma 6.23, $[a,\alpha]\subseteq P_{0}$. Then let
$X^{\lambda}:=W_{0}:=\{[a,\alpha]:(a,\alpha)\in P_{0}\}.$
If necessary, modify $X^{\lambda}$ so that it is a base set.
Now assume that $[b,\beta]$ and $W_{k}$ have already been defined for all $k<n$ and $(b,\beta)\in P_{k}$. Now define $[a,\alpha]$ for $(a,\alpha)\in P_{n}$, and define $W_{n}$ by
$[a,\alpha]:=\{[b,\beta]:(b,\beta)E(a,\alpha)\},\qquad W_{n}:=\{[a,\alpha]:(a,\alpha)\in P_{n}\}.$
(Observe that if $(b,\beta)E(a,\alpha)$, then $(b,\beta)\in P_{n-1}$ by Lemma 6.23, so that $[b,\beta]$ has already been defined.) Then define
$W:=\bigcup_{n\in\mathbb{N}}W_{n}.$
The elements of $W_{0}=X^{\lambda}$ act as atoms. The elements of $W-W_{0}$ act as sets. Observe that $W$ is transitive over sets: If $x\in[a,\alpha]$, where $[a,\alpha]\in W-W_{0}$ is a set, then $x=[b,\beta]$ for some $(b,\beta)E(a,\alpha)$ and hence $x\in W$ also. Note also that $W_{n}\subseteq V_{n}(X^{\lambda})$, for all $n\in\mathbb{N}$.
Lemma 6.24
  1. (a)
    If $\alpha\leq\beta<\lambda$ and $a\in V(X^{\alpha})$, then $[a,\alpha]=[\iota_{\alpha\beta}(a),\beta]$.
  2. (b)
    If $\alpha,\beta<\lambda$, then $[b,\beta]\in[X^{\alpha},\alpha]$ if and only if $b\in X^{\beta}$.
Proof: (a) If $(a,\alpha)\in P_{0}$, then
$[\iota_{\alpha\beta}(a),\beta]$ =$=\{(c,\gamma):(c,\gamma)\sim(\iota_{\alpha\beta}(a),\beta)\}$
=$=\{(c,\gamma):\exists\delta\geq\beta,\gamma\;\Big{(}\iota_{\gamma\delta}(c)=\iota_{\beta\delta}(\iota_{\alpha\beta}(a))=\iota_{\alpha\delta}(a)\Big{)}\}$
=$=\{(c,\gamma):(c,\gamma)\sim(a,\alpha)\}$
$=[a,\alpha]$
Similarly, if $(a,\alpha)\in P-P_{0}$, then
$[\iota_{\alpha\beta}(a),\beta]$ =$=\{[c,\gamma]:(c,\gamma)E(\iota_{\alpha\beta}(a),\beta)\}$
=$=\{[c,\gamma]:\exists\delta\geq\beta,\gamma\;\Big{(}\iota_{\gamma\delta}(c)\in\iota_{\beta\delta}(\iota_{\alpha\beta}(a))=\iota_{\alpha\delta}(a)\Big{)}\},$
=$=\{[c,\gamma]:(c,\gamma)E(a,\alpha)\},$
$=[a,\alpha].$
(b) We have
$[b,\beta]\in[X^{\alpha},\alpha]$ $\Leftrightarrow(b,\beta)E(X^{\alpha},\alpha),$
$\Leftrightarrow\iota_{\beta\gamma}(b)\in\iota_{\alpha\gamma}(X^{\alpha})\quad\text{for $\gamma\geq\alpha,\beta$},$
$\Leftrightarrow\iota_{\beta\gamma}(b)\in X^{\gamma},$
$\Leftrightarrow b\in X^{\beta},\quad\text{by the rank-preserving property of $\iota_{\beta\gamma}$.}$
$\dashv$
For $\alpha\leq\lambda$, define maps $V(X^{\alpha})\overset{\iota_{\alpha\lambda}}{\to}V(X^{\lambda})$ as follows:
$\iota_{\alpha\lambda}(a)=[a,\alpha]\quad\text{for $\alpha<\lambda$},\qquad\iota_{\lambda\lambda}=\text{id}_{V(X^{\lambda})}.$
We claim that each $\iota_{\alpha\lambda}$ is a bounded elementary embedding with the desired properties:
  1. (i)
    $\iota_{\lambda\lambda}=\text{id}_{V(X^{\lambda})}$.
  2. (ii)
    If $\alpha\leq\beta\leq\lambda$, then $\iota_{\beta\lambda}\circ\iota_{\alpha\beta}=\iota_{\alpha\lambda}$.
  3. (iii)
    If $\alpha\leq\lambda$, then $\iota_{\alpha\lambda}(X^{\alpha})=X^{\lambda}$.
(i) holds by definition. Observe that if $\alpha\leq\beta\leq\lambda$, and $(a,\alpha)\in P$, then
$\iota_{\beta\lambda}\circ\iota_{\alpha\beta}(a)=[\iota_{\alpha\beta}(a),\beta]=[a,\alpha]=\iota_{\alpha\lambda}(a),$
using Lemma 6.24(a). This proves (ii).
To prove (iii), note that
$\iota_{\alpha\lambda}(X^{\alpha})$ $=[X^{\alpha},\alpha],$
=$=\{[b,\beta]:[b,\beta]\in[X^{\alpha},\alpha]\},$
=$=\{[b,\beta]:b\in X^{\beta}=V_{0}(X^{\beta})\},$
=$=\{[b,\beta]:(b,\beta)\in P_{0}\},$
$=X^{\lambda},$
using Lemma 6.24(b).
It remains to show that each $\iota_{\alpha\lambda}$ is a bounded elementary embedding. This follows by induction on the complexity of formulas. For atomic formulas, we have
$V(X^{\alpha})\vDash a\in b\quad\Leftrightarrow\quad[a,\alpha]\in[b,\alpha]\quad\Leftrightarrow\quad V(X^{\lambda})\vDash\iota_{\alpha\lambda}(a)\in\iota_{\alpha\lambda}(b).$
Since $\iota_{\alpha\beta}$ is one-to-one, also $a=b$ if and only if $\iota_{\alpha\lambda}(a)=\iota_{\alpha\lambda}(b)$.
The propositional connectives ,$\land$,$\lnot$ , ¬ are dealt with very easily.
Suppose now that $\varphi(x_{1},\dots,x_{n})$ is of the form $\exists y\in x_{1}\;\psi(y,x_{1},\dots,x_{n})$. If $V(X^{\alpha})\vDash\exists y\in a_{1}\;\psi(y,a_{1},\dots,a_{n})$, then there is $b\in a_{1}$ such that $V(X^{\alpha})\vDash\psi(b,a_{1},\dots,a_{n})$. By induction hypothesis, $V(X^{\lambda})\vDash\psi(\iota_{\alpha\lambda}(b),\iota_{\alpha\lambda}(a_{1}),\dots,\iota_{\alpha\lambda}(a_{n}))$, where $\iota_{\alpha\lambda}(b)\in\iota_{\alpha\lambda}(a_{1})$, and hence $V(X^{\lambda})\vDash\exists y\in\iota_{\alpha\lambda}(a_{1})\;\psi(y,\iota_{\alpha\lambda}(a_{1}),\dots,\iota_{\alpha\lambda}(a_{n}))$.
Conversely, suppose that $V(X^{\lambda})\vDash\exists y\in\iota_{\alpha\lambda}(a_{1})\;\psi(y,\iota_{\alpha\lambda}(a_{1}),\dots,\iota_{\alpha\lambda}(a_{n}))$. Then there is $[b,\beta]\in[a_{1},\alpha]$ such that $V(X^{\lambda})\vDash\psi([b,\beta],[a_{1},\alpha],\dots,[a_{n},\alpha])$. Let $\gamma\geq\alpha,\beta$. Then $[b,\beta]=[\iota_{\beta\gamma}(b),\gamma]=\iota_{\gamma\lambda}(\iota_{\beta\gamma}(b))$ and $[a_{i},\alpha]=[\iota_{\alpha\gamma}(a_{i}),\gamma]=\iota_{\gamma\lambda}(\iota_{\alpha\gamma}(a_{i}))$, with $\iota_{\beta\gamma}(b)\in\iota_{\alpha\gamma}(a_{1})$. Thus
$V(X^{\lambda})\vDash\psi\Big{(}\iota_{\gamma\lambda}(\iota_{\beta\gamma}(b)),\iota_{\gamma\lambda}(\iota_{\alpha\gamma}(a_{1})),\dots,\iota_{\gamma\lambda}(\iota_{\alpha\gamma}(a_{n}))\Big{)},$
where $\iota_{\gamma\lambda}(\iota_{\beta\gamma}(b))\in\iota_{\gamma\lambda}(\iota_{\alpha\gamma}(a_{1}))$. By induction hypothesis, we obtain
$V(X^{\gamma})\vDash\psi(\iota_{\beta\gamma}(b),\iota_{\alpha\gamma}(a_{1}),\dots,\iota_{\alpha\gamma}(a_{n})),$
where $\iota_{\beta\gamma}(b)\in\iota_{\alpha\gamma}(a_{1})$, from which we obtain
$V(X^{\gamma})\vDash\exists y\in\iota_{\alpha\gamma}(a_{1})\;\psi(y,\iota_{\alpha\gamma}(a_{1}),\dots,\iota_{\alpha\gamma}(a_{n})).$
But as $V(X^{\alpha})\overset{\iota_{\alpha\gamma}}{\to}V(X^{\gamma})$ is a bounded elementary embedding, it follows also that
$V(X^{\alpha})\vDash\exists y\in a_{1}\psi(y,a_{1},\dots,a_{n}).$
This completes the induction, and the proof that each map $V(X^{\alpha})\overset{\iota_{\alpha\lambda}}{\to}V(X^{\lambda})$ is a bounded elementary embedding.
6.5.2 Construction of Polysaturated Extensions via Ultralimits
Recall Theorem 6.7, which states that for every superstructure $V(X)$ there is a set $I$ and an ultrafilter over $I$ such that the induced bounded elementary embedding $*:V(X)\to V(^{*}X)$ is an enlargement, where ${}^{*}X=X^{I}/\mathcal{U}$. The idea behind the proof of the existence of a polysaturated extension is to iterate this construction.
Let $V(X)$ be a superstructure over a base set $X$, and let $\kappa=|V(X)|$ be its cardinality. For $\alpha\leq\beta\leq\kappa^{+}$, we construct superstructures $V(X^{\alpha})$ and bounded elementary embeddings $*_{\alpha\beta}:V(X^{\alpha})\to V(X^{\beta})$ such that
  1. (i)
    If $\alpha\leq\kappa^{+}$, then $*_{\alpha\alpha}=\text{id}_{V(X^{\alpha})}$.
  2. (ii)
    If $\alpha\leq\beta\leq\gamma\leq\kappa^{+}$, then $*_{\beta\gamma}\circ*_{\alpha\beta}=*_{\alpha\gamma}$.
  3. (iii)
    If $\alpha\leq\beta\leq\kappa^{+}$, then $*_{\alpha\beta}(X^{\alpha})=X^{\beta}$.
We proceed by transfinite induction:
We define $V(X^{0})=V(X)$, and $*_{00}=\text{id}_{V(X)}$.
Suppose now that superstructures $V(X^{\alpha})$ and bounded elementary embeddings $*_{\alpha\beta}:V(X^{\alpha})\to V(X^{\beta})$ have already been constructed for $\alpha\leq\beta<\lambda$, such that (i)-(iii) are satisfied. We now consider two cases:
Case 1: $\lambda$ is a limit ordinal. In that case construct $V(X^{\lambda})$ and $*_{\alpha\lambda}$ as a limit of the $V(X^{\alpha})$ and $*_{\alpha\beta}$ for $\alpha,\beta<\lambda$. We have just seen that such a construction yields bounded elementary embeddings which preserve properties (i)-(iii).
Case 2: $\lambda=\gamma+1$ is a successor ordinal. In that case, let $V(X^{\gamma})\overset{*_{\gamma\lambda}}{\to}V(X^{\lambda})$ be an enlargement, as provided by Theorem 6.7. Observe that $X^{\lambda}={}^{*_{\gamma\lambda}}X^{\gamma}$ by construction. For $\alpha<\gamma$, define $*_{\alpha\lambda}=*_{\gamma\lambda}\circ*_{\alpha\gamma}$, and define $*_{\lambda\lambda}=\text{id}_{V(X^{\lambda})}$. It is then straightforward to show that that properties (i)-(iii) are satisfied.
Let’s briefly recall the construction of the final step in the transfinite induction: $V(X^{\kappa^{+}})$: This is a limit step, so we have
$P_{n}:=\{(a,\alpha):\alpha<\kappa^{+},a\in V_{n}(X^{\alpha})\},\qquad P=\bigcup_{n}P_{n}.$
The binary relations $\sim,E$ are given by
$(a,\alpha)\sim(b,\beta)\Leftrightarrow\exists\gamma\geq\alpha,\beta\;(*_{\alpha\gamma}(a)=*_{\beta\gamma}(b)),\qquad(a,\alpha)E(b,\beta)\Leftrightarrow\exists\gamma\geq\alpha,\beta\;(*_{\alpha\gamma}(a)\in*_{\beta\gamma}(b)).$
For $(a,\alpha)\in P_{0}$, we define $[a,\alpha]:=\{(b,\beta):(b,\beta)\sim(a,\alpha)\}$ and for $(a,\alpha)\in P-P_{0}$ we put $[a,\alpha]:=\{[b,\beta]:(b,\beta)E(a,\alpha)$. Then we define
$X^{\kappa^{+}}:=W_{0}:=\{[a,\alpha]:(a,\alpha)\in P_{0}\},\qquad W_{n}:=\{[a,\alpha]:(a,\alpha)\in P_{n}\},\qquad W:=\bigcup_{n}W_{n}.$
It then turns out that each $W_{n}\subseteq V_{n}(X^{\kappa^{+}})$, and that $W$ is a submodel of $V(X^{\kappa^{+}})$ which is transitive over sets. In addition,
$[a,\alpha]=[*_{\alpha\beta}(a),\beta]\quad\text{ for }\alpha\leq\beta<\kappa^{+}.$
Finally we define
$*_{\alpha\kappa^{+}}:V(X^{\alpha})\to V(X^{\kappa^{+}}):a\mapsto[a,\alpha].$
Let $\mathcal{A}\subseteq W_{n}$ be a family of cardinality $\kappa$ with the f.i.p. We will show that $\bigcap\mathcal{A}\neq\varnothing$. Note that each $A\in\mathcal{A}$ is of the form $[a,\alpha]$ for some $a$ in some $V_{n}(X^{\alpha})$. Suppose that
$\mathcal{A}=\{A_{\xi}:\xi<\kappa\}=\{[a_{\xi},\alpha_{\xi}]:\xi<\kappa\}\quad\text{is an enumeration of }\mathcal{A}.$
Let $\beta:=\sup_{\xi<\kappa}\alpha_{\xi}$. Since $\kappa^{+}$ is a regular cardinal and each $\alpha_{\xi}<\kappa^{+}$, we have that $\beta<\kappa^{+}$. Define $a_{\xi}^{\prime}=*_{\alpha_{\xi}\beta}(a_{\xi})\in V_{n}(X^{\beta})$, so that $A_{\xi}=[a_{\xi},\alpha_{\xi}]=[a^{\prime}_{\xi},\beta]$, and let $\mathcal{A}^{\prime}:=\{a^{\prime}_{\xi}:\xi<\kappa\}$ Observe that $*_{\beta\kappa^{+}}(a^{\prime}_{\xi})=[a^{\prime}_{\xi},\beta]=A_{\xi}$. As $\mathcal{A}$ has the f.i.p. and $*_{\beta\kappa^{+}}$ is a bounded elementary embedding, it follows easily that $\mathcal{A}^{\prime}\subseteq V_{n}(X^{\beta})$ has the f.i.p. as well. As $*_{\beta(\beta+1)}:V(X^{\beta})\to V(X^{\beta+1})$ is an enlargement, it follows that
$\bigcap*_{\beta(\beta+1)}[\mathcal{A}^{\prime}]\neq\varnothing.$
Now if $x\in\bigcap*_{\alpha(\alpha+1}[\mathcal{A}^{\prime}]$, then $x\in*_{\beta(\beta+1)}a_{\xi}^{\prime}$ for all $\xi<\kappa$. It follows that $*_{(\beta+1)\kappa^{+}}(x)\in*_{(\beta+1)\kappa^{+}}\circ*_{\beta(\beta+1)}(a_{\xi}^{\prime})=*_{\beta\kappa^{+}}(a^{\prime}_{\xi})=A_{\xi}$ for all $\xi<\kappa$, and hence that .$*_{(\beta+1)\kappa^{+}}(x)\in\bigcap\mathcal{A}$. Thus $\bigcap\mathcal{A}\neq\varnothing$ whenever $\mathcal{A}\subseteq W_{n}$ is a family of $\leq\kappa$-many sets with the f.i.p.
Observe that $X^{\kappa^{+}}=*_{\alpha\kappa^{+}}(X_{\alpha})$ for all $\alpha<\kappa^{+}$. We now show that each $*_{\alpha\kappa^{+}}:V(X^{\alpha})\to V(X^{\kappa^{+}})$ is $\kappa^{+}$-saturated.
By Theorem 5.3, we need only show that every family $\mathcal{A}\subseteq{}^{*}V_{n}(X^{\alpha})$ of cardinality $\kappa$ with the f.i.p. has non-empty intersection. Let $\mathcal{A}$ be such a family. Observe that if $A\in\mathcal{A}$, then $A\in{}^{*}V_{n}(X^{\alpha})=*_{\alpha\kappa^{+}}(V_{n}(X^{\alpha}))=[V_{n}(X^{\alpha}),\alpha]\in W_{n+1}$, and so $A\in W_{n}$, by transitivity of $W$ over sets and Lemma 6.23(b). It follows that $\mathcal{A}\subseteq W_{n}$. By what we have just seen, $\bigcap\mathcal{A}\neq\varnothing$.
Since $\kappa=|V(X)|$, it follows that if we define $*:=*_{0\kappa^{+}}$ and ${}^{*}X:=X^{\kappa^{+}}$, then the map $*:V(X)\to V({}^{*}X)$ is polysaturated.
Appendix A A Refresher on Basic First-Order Logic and Model Theory
A.1 First-Order Languages, Models and Satisfaction
A first-order language $\mathcal{L}=(\mathcal{R},\mathcal{F})$ consists of a collection of relation symbols (predicate symbols) $\mathcal{R}$ and function symbols $\mathcal{F}$. If =$\mathcal{R}=\{R_{1},\dots,R_{n}\}$ and =$\mathcal{F}=\{F_{1},\dots,F_{m}\}$ are finite sets, we may write $\mathcal{L}=(R_{1},\dots,R_{n},F_{1},\dots,F_{m})$.
A first-order structure (or model) for $\mathcal{L}$ is a set equipped with relations and functions that interpret these symbols. We will define what this means shortly. With every $R\in\mathcal{R}$ is associated an arity $n\in\mathbb{N}$, which indicates that $R$ is to be interpreted as an $n$-ary relation. Similarly, with every $F\in\mathcal{F}$ is associated an arity $n\in\mathbb{N}$, which indicates that $F$ is to be interpreted as an $n$-ary function. The function symbols of arity 0 are to be interpreted as constants.
Definition A.1
[…]
$\mathcal{R}^{\mathfrak{A}}=\{R^{\mathfrak{A}}:R\in\mathcal{R}\},\qquad\mathcal{F}^{\mathfrak{A}}=\{F^{\mathfrak{A}}:F\in\mathcal{F}\}.$
In particular if $c\in\mathcal{F}$ is a nullary function, then $c^{\mathfrak{A}}$ is a constant element of $A$.
The set $A$ is called the universe of $\mathfrak{A}$, and $\mathfrak{A}$ is said to be a model of $\mathcal{L}$.
If =$\mathcal{R}=\{R_{1},\dots,R_{n}\}$ and =$\mathcal{F}=\{F_{1},\dots,F_{m}\}$ are finite sets, we may write $\mathfrak{A}=(A,R_{1}^{\mathfrak{A}},\dots R_{n}^{\mathfrak{A}},F_{1}^{\mathfrak{A}},\dots,F_{m}^{\mathfrak{A}})$. When the interpretation is clear, we may dispense with the $\mathfrak{A}$-superscripts entirely, and simply write $(A,R_{1},\dots,R_{n},F_{1},\dots,F_{m})$.
$\Box$
For example, if $X$ is a base set, then the superstructure $\mathbb{U}:=(V(X),\in)$ is a $\mathcal{L}_{\in}$-structure.
Apart from the relation– and function symbols which define it, the first-order languages that we consider also come with various other symbols including:
  • Countably many variable symbols $x_{n}$ ($n\in\mathbb{N}$) — But we will often use $x,y,z,\dots$ instead.
  • The equality symbol $=$, which is always to be interpreted as equality.
  • Logical connectives $\lnot$ (not) and $\land$ (and).
  • The universal quantifier $\forall$ (for all).
  • Punctuation symbols such as parentheses and commas.
Definition A.2
(Terms and Formulas) Consider a first-order language $\mathcal{L}=(\mathcal{R},\mathcal{L})$.
  1. (a)
    The terms of $\mathcal{L}$ are defined inductively:
    1. (i)
      Every variable and every constant symbol is a term.
    2. (ii)
      If $F\in\mathcal{F}$ is a $n$-ary function symbol and $t_{1},\dots,t_{n}$ are terms, then the string $F(t_{1},\dots,t_{n})$ is a term.
    3. (iii)
      A string is a term if and only if it can be obtained via a finite number of applications of (i), (ii).
  2. (b)
    The atomic formulas of $\mathcal{L}$ are the expressions of the following type:
    1. (i)
      $s=t$, where $s,t$ are terms.
    2. (ii)
      $R(t_{1},\dots,t_{n})$, where $R$ is an $n$-ary relation symbol and $t_{1},\dots,t_{n}$ are terms.
  3. (c)
    The formulas of a language $\mathcal{L}=(\mathcal{R},\mathcal{L})$ are defined inductively:
    1. (i)
      Every atomic formula is a formula.
    2. (ii)
      If $\varphi,\psi$ are formulas and $x$ is a variable, then $\lnot\varphi,(\varphi\land\psi)$ and $(\forall x)\;\varphi$ are formulas.
    3. (iii)
      A string is a formula if and only if it can be obtained via a finite number of applications of (i), (ii).
$\Box$
We will also introduce a few other symbols, to simplify notation, namely $\lor$ (or), $\to$ (then, implies), $\leftrightarrow$ (if and only if) and $\exists$ (there exists). Suppose that $\varphi,\psi$ are formulas and that $x$ is a variable.
  • $(\varphi\lor\psi)$ is an abbreviation for ¬($\lnot(\lnot\varphi\land\lnot\psi)$.
  • $(\varphi\to\psi)$ is an abbreviation for ($(\lnot\varphi\lor\psi)$.
  • $(\varphi\leftrightarrow\psi)$ is an abbreviation for $((\varphi\to\psi)\land(\psi\to\varphi))$.
  • $(\exists x)\;\varphi$ is an abbreviation for $\lnot(\forall x)\;\lnot\varphi$.
In an effort to make formulas more readable, we may omit parentheses, or replace parentheses with brackets, etc.
A formula $\psi$ is a subformula of a formula $\varphi$ if $\psi$ is a consecutive string of symbols within the formula $\varphi$.
If $\varphi$ is a formula, then a variable $x$ is said to be within the scope of a quantifier $\forall x$ (or $\exists x$) occurring in $\varphi$ if there is a subformula of the form $\forall x\;\psi$ (or $\exists x\;\psi)$ such that $x$ occurs in $\psi$. A variable $x$ may occur a number of times within a formula $\varphi$. An occurrence of a variable $x$ in formula is said to be bound if it occurs within the scope of a quantifier; otherwise, the occurrence is said to be free.
A formula is said to be a sentence if it has no free variables, i.e. if every occurrence of a variable is bound.
Remarks A.3
For nonstandard universes, the appropriate language is $\mathcal{L}_{\in}$, consisting of just one binary relation symbol $\in$. In addition, we typically work with a modification of the first order language, where the quantifiers are bounded, i.e. of the form $\forall y\in x$ and $\exists y\in x$.
$\Box$
Definition A.4
[…]
  1. (a)
    Terms: The value $t[a_{1},\dots,a_{n}]\in A$ of a term $t(x_{1},\dots,x_{n})$ at $a_{1},\dots,a_{n}\in A$ is defined inductively as follows:
    1. (i)
      If $t\equiv x_{i}$ is a variable, then $t[a_{1},\dots,a_{n}]:=a_{i}$.
      If $t\equiv c$ is a constant (i.e. a nullary function symbol), then $t[a_{1},\dots,a_{n}]:=c^{\mathfrak{A}}$.
    2. (ii)
      If $t\equiv F(t_{1},\dots,t_{m})$, where $F$ is an $m$-ary function symbol and $t_{1},\dots,t_{m}$ are terms, then
      $t[a_{1},\dots,a_{n}]:=F^{\mathfrak{A}}(t_{1}[a_{1},\dots,a_{n}],\dots,t_{m}[a_{1},\dots,a_{n}]).$
  2. (b)
    Formulas: For a formula $\varphi(x_{1},\dots,x_{n})$, the satisfaction relation $\mathfrak{A}\vDash\varphi[a_{1},\dots,a_{n}]$ is defined inductively, as follows.
    1. (i)
      If $\varphi$ is the atomic formula $t_{1}=t_{2}$, then $\mathfrak{A}\vDash\varphi[a_{1},\dots,a_{n}]$ if and only if $t_{1}[a_{1},\dots,a_{n}]=t_{2}[a_{1},\dots,a_{n}]$.
      Similarly, if $\varphi$ is the atomic formula $R(t_{1},\dots,t_{m})$, then $\mathfrak{A}\vDash\varphi[a_{1},\dots,a_{n}]$ if and only if the relation $R^{\mathfrak{A}}(t_{1}[a_{1},\dots,a_{n}],\dots,t_{m}[a_{1},\dots,a_{n}])$ holds in $\mathfrak{A}$.
    2. (ii)
      If $\varphi$ is the formula $\lnot\psi$, then $\mathfrak{A}\vDash\varphi[a_{1},\dots,a_{n}]$ if and only if not $\mathfrak{A}\vDash\psi[a_{1},\dots,a_{n}]$.
    3. (iii)
      If $\varphi$ is the formula $\psi\land\chi$, then $\mathfrak{A}\vDash\varphi[a_{1},\dots,a_{n}]$ if and only if both $\mathfrak{A}\vDash\psi[a_{1},\dots,a_{n}]$ and $\mathfrak{A}\vDash\chi[a_{1},\dots,a_{n}]$
    4. (iv)
      If $\varphi$ is the formula $\forall x_{i}\;\psi$ (where $x_{i}\in\{x_{1},\dots,x_{n}\}$), then
      $\mathfrak{A}\vDash\varphi[a_{1},\dots,a_{n}]\quad\text{if and only if for every $a\in A$,}\quad\mathfrak{A}\vDash\psi[a_{1},\dots,a_{i-1},a,a_{i+1},\dots,a_{n}].$
$\Box$
It is easy to verify that:
  • $\mathfrak{A}\vDash(\varphi\lor\psi)[a_{1},\dots,a_{n}]$ if and only if $\mathfrak{A}\vDash\varphi[a_{1},\dots,a_{n}]$ or $\mathfrak{A}\vDash\psi[a_{1},\dots,a_{n}]$.
  • $\mathfrak{A}\vDash(\varphi\to\psi)[a_{1},\dots,a_{n}]$ if and only if whenever $\mathfrak{A}\vDash\varphi[a_{1},\dots,a_{n}]$, then also $\mathfrak{A}\vDash\psi[a_{1},\dots,a_{n}]$.
  • $\mathfrak{A}\vDash(\exists x_{i}\;\varphi)[a_{1},\dots,a_{n}]$ if and only if there is $a\in A$ such that $\mathfrak{A}\vDash\psi[a_{1},\dots,a_{i-1},a,a_{i+1},\dots,a_{n}]$.
The next lemma show that whether or not a $\mathfrak{A}\vDash\varphi[a_{1},\dots,a_{n}]$ depends only on those $a_{i}$ which correspond to variables $x_{i}$ that have a free occurrence in $\varphi$.
Lemma A.5
  1. (a)
    If $t$ is a term with variables among $x_{1},\dots,x_{n}$, then $t[a_{1},\dots,a_{n}]$ depends only on the values $a_{i}$ corresponding to variables $x_{i}$ which actually occur in $t$.
    More precisely, suppose that the variables occurring in a term $t$ are among $x_{1},\dots,x_{n}$. Suppose further that $a_{1},\dots,a_{p}$ and $b_{1},\dots,b_{q}$ are elements of $A$, where $p,q\geq n$, and that $a_{i}=b_{i}$ whenever $x_{i}$ actually occurs in $t$. Then $t[a_{1},\dots,a_{p}]=t[b_{1},\dots,b_{q}]$.
  2. (b)
    If $\varphi$ is a formula with variables among $x_{1},\dots,x_{n}$, then whether or not $\mathfrak{A}\vDash\varphi[a_{1},\dots,a_{n}]$ depends only on those $a_{i}$ for which $x_{i}$ has a free occurrence in $\varphi$. More precisely, suppose that the variables occurring in a formula $\varphi$ are among $x_{1},\dots,x_{n}$, where some occur freely, and the others bound. Suppose further that $a_{1},\dots,a_{p}$ and $b_{1},\dots,b_{q}$ are elements of $A$, where $p,q\geq n$, and $a_{i}=b_{i}$ whenever $i\leq n$ and $x_{i}$ has a free occurrence in $\varphi(x_{1},\dots,x_{n})$.
    Then
    $\mathfrak{A}\vDash\varphi[a_{1},\dots,a_{p}]\Longleftrightarrow\mathfrak{A}\vDash\varphi[b_{1},\dots,b_{q}].$
Proof: These facts are is easily proved via induction on the length of $t,\varphi$.
(a) If $t$ is the variable $x_{i}$, then $t[a_{1},\dots,a_{p}]=t[b_{1},\dots,b_{q}]$ whenever $a_{i}=b_{i}$. If $t$ is a constant symbol, the result is obvious. If $t$ is the term $F(t_{1},\dots,t_{m})$, then
$t[a_{1},\dots,a_{p}]=F^{\mathfrak{A}}(t_{1}[a_{1},\dots,a_{p}],\dots,t_{m}[a_{1},\dots,a_{p}])$. But the length of $t_{1},\dots,t_{n}$ is clearly less than the length of $t$, so by induction we have that $t_{i}[a_{1},\dots,a_{p}]=t_{i}[b_{1},\dots,b_{q}]$ for $i\leq m$. Clearly, therefore $t[a_{1},\dots,a_{p}]=t[b_{1},\dots,b_{q}]$.
(b) If $\varphi$ is an atomic formula, then all variables that occur in $\varphi$ are free, and hence the result follows by (a). If $\varphi$ is of the form $\lnot\psi$, then the free variables of $\varphi$ are the same as the free variables of $\psi$. Now $\mathfrak{A}\vDash\varphi[a_{1},\dots,a_{p}]$ if and only if not $\mathfrak{A}\vDash\psi[a_{1},\dots,a_{p}]$. But as the length of the formula $\psi$ is shorter than that of $\varphi$, we have $\mathfrak{A}\vDash\psi[a_{1},\dots,a_{n}]\Leftrightarrow\mathfrak{A}\vDash\psi[b_{1},\dots,b_{q}]$, and hence $\mathfrak{A}\vDash\varphi[a_{1},\dots,a_{p}]\Leftrightarrow\mathfrak{A}\vDash\varphi[b_{1},\dots,b_{q}]$. The case where $\varphi$ is of the form $\psi\land\chi$ is dealt with in a similar fashion. Finally, if $\varphi$ is of the form $\forall x_{i}\;\psi$, then the free variables of $\psi$ are just the free variables of $\varphi$, plus (possibly) the variable $x_{i}$. Now
$\phantom{aaaaa}\mathfrak{A}\vDash\varphi[a_{1},\dots,a_{p}],$
for every $\Leftrightarrow\quad\text{for every $a\in A$, }\mathfrak{A}\vDash\psi[a_{1},\dots,a_{i-1},a,a_{i+1},\dots,a_{p}],$
[…]
$\Leftrightarrow\quad\mathfrak{A}\vDash\varphi[b_{1},\dots,b_{q}].$
$\dashv$
By the above lemma, we may henceforth write $\varphi(x_{1},\dots,x_{n})$ to indicate a formula whose free variables are among $x_{1},\dots,x_{n}$.
Corollary A.6
Truth-values of sentences are fixed: If $\varphi$ is a sentence, then either $\mathfrak{A}\vDash\varphi[a_{1},\dots,a_{n}]$ for all sequences $a_{1},\dots,a_{n}\in A$, or for none of them.
$\Box$
Suppose that $\Sigma$ is a a set of $\mathcal{L}$-sentences, and that $\mathfrak{A}$ is an $\mathcal{L}$-model. We say that $\mathfrak{A}$ is a model of $\Sigma$, or that $\mathfrak{A}$ satisfies $\Sigma$ — and write $\mathfrak{A}\vDash\Sigma$ — if and only if $\mathfrak{A}\vDash\varphi$ for every $\varphi\in\Sigma$.
A.2 Elementary Embeddings and Elementary Equivalence
Definition A.7
[…]
  1. (a)
    $A\subseteq B$.
  2. (b)
    If $R$ as an $n$-ary relation symbol of $\mathcal{L}$, then $R^{\mathfrak{A}}=R^{\mathfrak{B}}\restriction A$, i.e. for any $a_{1},\dots,a_{n}\in A$, we have that $R^{\mathfrak{A}}(a_{1},\dots,a_{n})$ holds in $\mathfrak{A}$ if and only if $R^{\mathfrak{B}}(a_{1},\dots,a_{n})$ holds in $\mathfrak{B}$.
  3. (c)
    Similarly, if $F$ as an $n$-ary function symbol of $\mathcal{L}$, then $F^{\mathfrak{A}}=F^{\mathfrak{B}}\restriction A$.
    In particular, if $c$ is a constant symbol, then $c^{\mathfrak{A}}=c^{\mathfrak{B}}$.
$\Box$
Definition A.8
[…]
  1. (a)
    The theory of $\mathfrak{A}$ is the set ${Th}(\mathfrak{A})$ of all $\mathcal{L}$-sentences that are satisfied by $\mathfrak{A}$.
  2. (b)
    Let $X\subseteq A$. By $\mathcal{L}_{X}$, we mean the language $\mathcal{L}$ augmented with additional constant symbols $\{c_{a}:a\in X\}$. The model $\mathfrak{A}_{X}:=(\mathfrak{A},a)_{a\in X}$ denotes the expansion of $\mathfrak{A}$ to a model of $\mathcal{L}_{X}$, where, for $a\in X$, the new constant $c_{a}$ is interpreted to be the element $a\in X$.
  3. (c)
    The elementary diagram $\Gamma_{\mathfrak{A}}$ of $\mathfrak{A}$ is the theory $\text{Th}(\mathfrak{A}_{A})$ of the expansion $\mathfrak{A}_{A}:=(\mathfrak{A},a)_{a\in A}$, i.e. it is the set of all sentences of $\mathcal{L}_{A}$ which hold in $\mathfrak{A}_{A}$.
$\Box$
[…]
$\mathfrak{A}_{A}\vDash\varphi(c_{a_{1}},\dots,c_{a_{n}})\quad\text{if and onlyif }\quad\mathfrak{A}\vDash\varphi[a_{1},\dots,a_{n}].$
To simplify notation, we will write $\varphi(a_{1},\dots,a_{n})$ instead of $\varphi(c_{a_{1}},\dots,c_{a_{n}})$. Then the elementary diagram of $\mathfrak{A}$ is
$\Gamma_{\mathfrak{A}}=\{\varphi(a_{1},\dots,a_{n}):\mathfrak{A}\vDash\varphi[a_{1},\dots,a_{n}]\}.$
Definition A.9
(Elementary Equivalence, Elementary Embedding) Suppose that $\mathfrak{A},\mathfrak{B}$ are two models of a first-order language $\mathcal{L}$.
  1. (a)
    We say that $\mathfrak{A},\mathfrak{B}$ are elementarily equivalent — and write $\mathfrak{A}\equiv\mathfrak{B}$ — if and only if $\mathfrak{A},\mathfrak{B}$ satisfy the same $\mathcal{L}$-sentences, i.e. $\text{Th}(\mathfrak{A})=\text{Th}(\mathfrak{B})$.
  2. (b)
    We say that $\mathfrak{A}$ is an elementary submodel of $\mathfrak{B}$ — and write $\mathfrak{A}\preceq\mathfrak{B}$ — if and only if $\mathfrak{A}\subseteq\mathfrak{B}$ and for all $\mathcal{L}$-formulas $\varphi(x_{1},\dots,x_{n})$ and all $a_{1},\dots,a_{n}\in A$,
    $\mathfrak{A}\vDash\varphi[a_{1},\dots,a_{n}]\quad\text{if and only if}\quad\mathfrak{B}\vDash\varphi[a_{1},\dots,a_{n}].$
  3. (c)
    We say that a map $f:A\to B$ is an elementary embedding — and write $f:\mathfrak{A}\overset{\equiv}{\hookrightarrow}\mathfrak{B}$ — if and only if for all $\mathcal{L}$-formulas $\varphi(x_{1},\dots,x_{n})$ and all $a_{1},\dots,a_{n}\in A$,
    $\mathfrak{A}\vDash\varphi[a_{1},\dots,a_{n}]\quad\text{if and only if}\quad\mathfrak{B}\vDash\varphi[f(a_{1}),\dots,f(a_{n})].$
$\Box$
Observe the following trivial fact:
Lemma A.10
If there is an elementary embedding $\mathfrak{A}\overset{\equiv}{\hookrightarrow}\mathfrak{B}$, then $\mathfrak{A}\equiv\mathfrak{B}$.
$\Box$
The following lemma follows by chasing through the above definitions:
Lemma A.11
Suppose that $\mathfrak{A},\mathfrak{B}$ are two models of a first-order language $\mathcal{L}$.
  1. (a)
    If $\mathfrak{A}\subseteq\mathfrak{B}$, then $\mathfrak{A}\preceq\mathfrak{B}$ if and only if $(\mathfrak{B},a)_{a\in A}\vDash\Gamma_{\mathfrak{A}}$.
  2. (b)
    Similarly, there is an elementary embedding $\mathfrak{A}\overset{\equiv}{\hookrightarrow}\mathfrak{B}$ if and only if there is an expansion $(\mathfrak{B},b_{a})_{a\in A}$ of $\mathfrak{B}$ such that $(\mathfrak{B},b_{a})_{a\in A}\vDash\Gamma_{\mathfrak{A}}$ (and then $a\mapsto b_{a}$ supplies the required elementary embedding).
$\Box$
A.3 Ultrafilters
Definition A.12
(Ultrafilter) Let $I$ be a set.
  1. (a)
    A family $\mathcal{A}\subseteq\mathcal{P}(I)$ is said to satisfy the finite intersection property (f.i.p.) if and only whenever :=$\mathcal{B}:=\{A_{1},\dots,A_{n}\}$ is any finite subcollection of $\mathcal{A}$, then $\bigcap\mathcal{B}$ is non-empty, i.e. $A_{1}\cap\dots\cap A_{n}\neq\varnothing$.
  2. (b)
    A filter over $I$ is a non-empty family $\mathcal{F}\subseteq\mathcal{P}(I)$ with the following properties:
    1. (i)
      $\varnothing\not\in\mathcal{F}$.
    2. (ii)
      $F,G\in\mathcal{F}$ implies $F\cap G\in\mathcal{F}$.
    3. (iii)
      $F\in\mathcal{F}$ and $G\supseteq F$ implies $G\in\mathcal{F}$.
  3. (c)
    A filter $\mathcal{U}$ over $I$ is said to be an ultrafilter if and only if for every $A\subseteq I$ we have either $A\in\mathcal{U}$ or its complement $A^{c}\in\mathcal{U}$.
  4. (d)
    A filter $\mathcal{F}$ is said to be countably incomplete if and only if there is a countable subfamily $\{F_{n}:n\in\mathcal{F}\}$ of members of $\mathcal{F}$ such that $\bigcap_{n}F_{n}\not\in\mathcal{F}$.
In Section 6.1 it will be shown that any ultrafilter will induce a transfer map $*:V(X)\to V(Y)$ between superstructures. In order for that transfer map to give rise to a nonstandard framework — i.e. so that ${}^{\sigma}C\subsetneq{}^{*}C$ for some countable $C\in V(X)$ — we will require that the ultrafilter is countably incomplete.
Here follow some basic facts:
Theorem A.13
Let $I$ be a set.
  1. (a)
    Every filter on $I$ has the f.i.p.
  2. (b)
    if $\mathcal{A}$ is a family of subsets of a set $I$ with the f.i.p., then
    :=$\mathcal{F}:=\{F\subseteq I:\text{there are }A_{1},\dots,A_{n}\in\mathcal{A}\text{ such that }A_{1}\cap\dots\cap A_{n}\subseteq F\}$
    is a filter containing $\mathcal{A}$.
  3. (c)
    If $\mathcal{A}$ is a family of subsets of a set $I$ with the f.i.p. and $B\subseteq I$, then either $\mathcal{A}\cup\{B\}$ has the f.i.p., or else $\mathcal{A}\cup\{B^{c}\}$ has the f.i.p.
  4. (d)
    A filter over $I$ is an ultrafilter if and only if it is a maximal filter, i.e. if and only if it is not contained in any strictly larger filter.
  5. (e)
    If $\mathcal{A}$ is a family of subsets of a set $I$ with the f.i.p., then there is an ultrafilter $\mathcal{U}$ such that $\mathcal{A}\subseteq\mathcal{U}$.
  6. (f)
    $\mathcal{U}$ is an ultrafilter over $I$ if and only if whenever $A_{1},\dots,A_{n}\subseteq I$ are such that $A_{1}\cup\dots\cup A_{n}\in\mathcal{U}$, then there is $i\leq n$ such that $A_{i}\in\mathcal{U}$.
  7. (g)
    $\mathcal{U}$ is a countably incomplete ultrafilter if and only if there is a partition $I$ into a sequence $I_{n}$ $(n\in\mathbb{N})$ of disjoint non-empty sets with the property that $\bigcap_{n}I_{n}=I$, and $I_{n}\not\in\mathcal{U}$ for any $n$.
Proof: (a) is obvious.
(b) is straightforward.
(c) Suppose that $\mathcal{A}\cup\{B\}$ does not have the f.i.p., then there exists $A_{1},\dots,A_{n}\in\mathfrak{A}$ such that $A_{1}\cap\dots\cap A_{n}\cap B=\varnothing$, so that $A_{1}\cap\dots\cap A_{n}\subseteq B^{c}$. Then if $A^{\prime}_{1},\dots,A^{\prime}_{m}\in\mathcal{A}$, we have $A^{\prime}_{1}\cap\dots\cap A^{\prime}_{m}\cap B^{c}\supseteq A^{\prime}_{1}\cap\dots\cap A^{\prime}_{m}\cap A_{1}\cap\dots\cap A_{n}\neq\varnothing$, as $\mathcal{A}$ has the f.i.p. and $A^{\prime}_{i},A_{j}\in\mathcal{A}$. Thus $\mathcal{A}\cup\{B^{c}\}$ has the f.i.p.
(d) Suppose that $\mathcal{U}$ is an ultrafilter over $I$. If $\mathcal{F}$ is a strictly larger filter and $F\in\mathcal{F}-\mathcal{U}$, then $F^{c}\in\mathcal{U}$, and hence $F\cap F^{c}=\varnothing\in\mathcal{F}$ — contradicting the definition of filter. Conversely, an easy application of Zorn’s Lemma shows that any filter can be extended to a maximal filter. If $\mathcal{F}$ is a maximal filter and $B\subseteq I$, then either $\mathcal{F}\cup\{B\}$ or $\mathcal{F}\cup\{B^{c}\}$ has the f.i.p. Hence there is a filter $\mathcal{G}$ such that $\mathcal{G}\supseteq\mathcal{F}\cup\{B\}$ or $\mathcal{G}\supseteq\mathcal{F}\cup\{B^{c}\}$. But as $\mathcal{F}$ is maximal, we must have $\mathcal{G}=\mathcal{F}$. Thus either $B\in\mathcal{F}$ or $B^{c}\in\mathcal{F}$, proving that $\mathcal{F}$ is an ultrafilter. Thus the ultrafilters are precisely the maximal filters.
(e) By (b), there is a filter $\mathcal{F}$ such that $\mathcal{A}\subseteq\mathcal{F}$. By an easy application of Zorn’s Lemma, there is a maximal filter $\mathcal{U}$ over $I$ such that $\mathcal{F}\subseteq\mathcal{U}$. By (d), $\mathcal{U}$ is an ultrafilter.
(f) Suppose that $\mathcal{U}$ is an ultrafilter with $A_{1}\cup\dots\cup A_{n}\in\mathcal{U}$. If $A_{i}\not\in\mathcal{U}$ for any $i\leq n$, then also $A_{1}^{c}\cap\dots\cap A_{n}^{c}=(A_{1}\cup\dots\cup A_{n})^{c}\in\mathcal{U}$, and hence $\varnothing\in\mathcal{U}$ — contradicting the definition of filter. Conversely, if $\mathcal{F}$ is a filter over $I$ with the stated property and $B\subseteq I$, then $I=B\cup B^{c}\in\mathcal{F}$, and hence either $B\in\mathcal{F}$ or $B^{c}\in\mathcal{F}$. Hence $\mathcal{F}$ is an ultrafilter.
(g) Suppose that $\mathcal{U}$ is countably incomplete. Then there are $U_{n}\in\mathcal{U}$ (for $n\in\mathbb{N}$) such that $\bigcap_{n}U_{n}\not\in\mathcal{U}$. Since $\mathcal{U}$ is closed under finite intersections, we may assume that $U_{1}\supseteq U_{2}\supseteq U_{3}\supseteq\dots$, and that $U_{1}=I$. By removing duplicates, we may assume that all the $U_{n}$ are distinct. Now define
$I_{0}:=\bigcap_{n}U_{n},\quad\text{and}\quad I_{n}:=U_{n}-U_{n+1}\text{ for }n>0.$
Then $I=\bigcup_{n}I_{n}$, and clearly the sets $I_{n}$ are non-empty and partition $I$. By assumption, $I_{0}\not\in\mathcal{U}$. Furthermore, since $U_{n+1}^{c}\supseteq U_{n}-U_{n+1}=I_{n}$, we cannot have $I_{n}\in\mathcal{U}$.
Conversely, if $\{I_{n}:n\in\mathbb{N}\}$ is a partition of $I$ into disjoint non-empty sets, then $\bigcap_{m\neq n}I_{m}^{c}=I_{n}\neq\varnothing$, from which it follows that $\{I_{n}^{c}:n\in\mathbb{N}\}$ has the f.i.p. It follows that there is an ultrafilter $\mathcal{U}$ such that $I_{n}^{c}\in\mathcal{U}$ for each $n\in\mathbb{N}$. As $\bigcap_{n}I_{n}^{c}=\varnothing\not\in\mathcal{U}$, the ultrafilter $\mathcal{U}$ is countably incomplete.
$\dashv$
A.4 Ultraproducts and Ultrapowers
Suppose that $I$ is a set, and that $\mathcal{F}$ is a a filter over $I$. Suppose further that $\mathfrak{A}_{i}=(A_{i},\mathcal{L}^{\mathfrak{A}_{i}})$ $(i\in I)$ are non-empty models of a first-order language $\mathcal{L}$. Recall that the product $\prod_{i\in I}A_{i}$ is the set of all choice functions $f:I\to\bigcup_{i\in I}A_{i}$, i.e. all those functions with the property that $f(i)\in A_{i}$ for all $i\in I$. Define a binary relation $\sim_{\mathcal{F}}$ on the set $\prod_{i\in I}A_{i}$ by
$f\sim_{\mathcal{F}}g\quad\text{if and only if }\quad\{i\in I:f(i)=g(i)\}\in\mathcal{F}.$
It is straightforward to verify that $\sim_{\mathcal{F}}$ is an equivalence relation. For example, if $f\sim_{\mathcal{F}}g$ and $g_{\mathcal{F}}H$, then
$\{i\in I:f(i)=h(i)\}\supseteq\{i\in I:f(i)=g(i)\}\cap\{i\in I:g(i)=h(i)\}\in\mathcal{F},$
from which it follows that $f\sim_{\mathcal{F}}h$, i.e. that $\sim_{\mathcal{F}}$ is transitive.
Denote the equivalence relation corresponding to $f\in\prod_{I}A_{i}$ by $f/\mathcal{F}$, and let
$\prod_{I}A_{i}/\mathcal{F}:=\{f/\mathcal{F}:f\in\prod_{I}A_{i}\}$
denote the corresponding quotient set.
We now show how to equip $\prod_{I}A_{i}/\mathcal{F}$ with relations and functions which turn it into a $\mathcal{L}$-structure $\mathfrak{B}=\prod_{I}\mathfrak{A}_{i}/\mathcal{F}$. For $n$-ary relation– and function symbols $R,F$ of $\mathcal{L}$, let $R^{\mathfrak{A}_{i}},F^{\mathfrak{A}_{i}}$ denote their corresponding interpretations in the model $\mathfrak{A}_{i}$. Define the relation $R^{\mathfrak{B}}$ and function $F^{\mathfrak{B}}$ on $B:=\prod_{I}A_{i}/\mathcal{F}$ by
$R^{\mathfrak{B}}(f_{1}/\mathcal{F},\dots,f_{n}/\mathcal{F})\quad\text{if and only if}\quad\{i\in I:R^{\mathfrak{A}_{i}}(f_{1}(i),\dots,f_{n}(i))\}\in\mathcal{F},$
and
$F^{\mathfrak{B}}(f_{1}/\mathcal{F},\dots,f_{n}/\mathcal{F}):=g/\mathcal{F},\quad\text{where }g(i):=F^{\mathfrak{A}_{i}}(f_{1}(i),\dots,f_{n}(i)).$
It is easy to verify that these notions are well-defined. For example, if $f_{i}/\mathcal{F}=g_{i}/\mathcal{F}$ for $i\leq n$ and $\{i\in I:R^{i}(f_{1}(i),\dots,f_{n}(i))\}\in\mathcal{F}$, then also
$\{i:R^{i}(g_{1}(i),\dots,g_{n}(i))\}\supseteq\{i:R^{i}(f_{1}(i),\dots,f_{n}(i))\}\cap\{i:f_{1}(i)=g_{1}(i)\}\cap\dots\cap\{i\in I:f_{n}(i)=g_{n}(i)\}\in\mathcal{F}.$
The $\mathcal{L}$-structure with base set $\prod_{I}A_{i}$ and corresponding relations and functions $R^{\mathfrak{B}},F^{\mathfrak{B}}$ (for $R,F\in\mathcal{L}$) is called the reduced product of the models $\mathfrak{A}_{i}$, and denoted by $\prod_{I}\mathfrak{A}_{i}/\mathcal{F}$. If the $\mathfrak{A}_{i}$ are all identical, then we have a reduced power, and denote it by $\mathfrak{A}^{I}/\mathcal{F}$. If $\mathcal{F}$ is an ultrafilter over $I$, then a reduced product is called an ultraproduct, and a reduced power is called an ultrapower.
Lemma A.14
Suppose that $\mathfrak{A}_{i}$ $(i\in I)$ are models of a first-order language $\mathcal{L}$, and that $\mathcal{F}$ is a filter over $I$. If $t(x_{1},\dots,x_{n})$ is an $\mathcal{L}$-term, then the interpretation $t^{\mathcal{F}}$ of $t$ in $\prod_{I}\mathfrak{A}_{i}/\mathcal{F}$ is given by
$t^{\mathfrak{B}}(f_{1}/\mathcal{F},\dots,f_{n}/\mathcal{F})=g/\mathcal{F},\quad\text{where}\quad g(i):=t^{\mathfrak{A}_{i}}(f_{1}(i),\dots,f_{n}(i)).$
Similarly, if $R$ is an $m$-ary relation symbol of $\mathcal{L}$, and $t_{1},\dots,t_{m}$ are $\mathcal{L}$-terms, then
$R^{\mathfrak{B}}\Big{(}t_{1}^{\mathfrak{B}}(f_{1}/\mathcal{F},\dots,f_{n}/\mathcal{F}),\dots,t_{m}^{\mathfrak{B}}(f_{1}/\mathcal{F},\dots,f_{n}/\mathcal{F})\Big{)}$ holds in $\prod_{I}\mathfrak{A}_{i}/\mathcal{F}$ if and only if
$\Big{\{}i\in I:R^{\mathfrak{A}_{i}}\Big{(}t_{1}^{\mathfrak{A}_{i}}(f_{1}(i),\dots,f_{n}(i)),\dots,t_{m}^{\mathfrak{A}_{i}}(f_{1}(i),\dots,f_{n}(i))\Big{)}\Big{\}}\in\mathcal{F}.$
Proof: Induction: If $t\equiv F(x_{1},\dots,x_{n})$ for some $n$-ary function symbol $F\in\mathcal{L}$, the result follows by definition of $F^{\mathfrak{B}}$. If $t\equiv F(t_{1}(x_{1},\dots,x_{n}),\dots,t_{m}(x_{1},\dots,x_{n}))$ for some $m$-ary function symbol $F$ and terms $t_{1},\dots,t_{n}$, then by the inductive definition of $t^{\mathfrak{B}}$ (cf. Definition A.4), we have
$t^{\mathfrak{B}}(f_{1}/\mathcal{F},\dots,f_{n}/\mathcal{F}):=F^{\mathfrak{B}}\Big{(}t_{1}^{\mathfrak{B}}(f_{1}/\mathcal{F},\dots,f_{n}/\mathcal{F}),\dots,t_{m}^{\mathfrak{B}}(f_{1}/\mathcal{F},\dots,f_{n}/\mathcal{F})\Big{)}.$
By induction hypothesis, we have, for $k\leq m$,
$t_{k}^{\mathfrak{B}}(f_{1}/\mathcal{F},\dots,f_{n}/\mathcal{F})=g_{k}/\mathcal{F},\quad\text{where}\quad g_{k}(i):=t_{k}^{\mathfrak{A}_{i}}(f_{1}(i),\dots,f_{n}(i)).$
Thus
$t^{\mathfrak{B}}(f_{1}/\mathcal{F},\dots,f_{n}/\mathcal{F})=F^{\mathfrak{B}}(g_{1}/\mathcal{F},\dots,g_{m}/\mathcal{F})=g/\mathcal{F}\quad\text{where}\quad g(i):=F^{\mathfrak{A}_{i}}(g_{1}(i),\dots,g_{m}(i)).$
But then by the inductive definition of $t^{\mathfrak{A}_{i}}$ we have
$g(i)=F^{\mathfrak{A}_{i}}\Big{(}t_{1}^{\mathfrak{A}_{i}}(f_{1}(i),\dots,f_{n}(i)),\dots,t_{m}^{\mathfrak{A}_{i}}(f_{1}(i),\dots,f_{n}(i))\Big{)}=t^{\mathfrak{A}_{i}}(f_{1}(i),\dots,f_{n}(i)),$
from which the result follows.
Now consider the case $R(t_{1},\dots,t_{m})$. For $k\leq m$, let $t_{k}^{\mathfrak{B}}(f_{1}/\mathcal{F},\dots,f_{n}/\mathcal{F})=g_{k}/\mathcal{F}$, where $g_{k}(i):=t^{\mathfrak{A}_{i}}(f_{1}(i),\dots,f_{n}(i))$. Then we have $R^{\mathfrak{B}}\Big{(}t_{1}^{\mathfrak{B}}(f_{1}/\mathcal{F},\dots,f_{n}/\mathcal{F}),\dots,t_{m}^{\mathfrak{B}}(f_{1}/\mathcal{F},\dots,f_{n}\mathcal{F})\Big{)}$ if and only if $\mathcal{R}^{\mathfrak{B}}(g_{1}/\mathcal{F},\dots,g_{m}/\mathcal{F})$ if and only if $\{i\in I:R^{\mathfrak{A}_{i}}(g_{1}(i),\dots,g_{n}(i))\}\in\mathcal{F}$ if and only if $\Big{\{}i\in I:R^{\mathfrak{A}_{i}}\Big{(}t_{1}^{\mathfrak{A_{i}}}(f_{1}(i),\dots,f_{n}(i)),\dots,t_{m}^{\mathfrak{A_{i}}}(f_{1}(i),\dots,f_{n}(i))\Big{)}\Big{\}}\in\mathcal{F}$.
$\dashv$
Theorem A.15
(Łos) Suppose that $\mathfrak{A}_{i}$ $(i\in I)$ are models of a first-order language $\mathcal{L}$, and that $\mathcal{U}$ is an ultrafilter over $I$. If $\varphi(x_{1},\dots,x_{n})$ is an $\mathcal{L}$-formula whose free variables are among $x_{1},\dots,x_{n}$, then
$\prod_{I}\mathfrak{A}_{i}\Big{/}\mathcal{U}\vDash\varphi[f_{1}/\mathcal{U},\dots,f_{n}/\mathcal{U}]\quad\Longleftrightarrow\quad\{i\in I:\mathfrak{A}_{i}\vDash\varphi[f_{1}(i),\dots,f_{n}(i)]\}\in\mathcal{U}.$
Proof: Induction on the complexity of $\varphi$. If $\varphi$ is an atomic sentence, then the result follows from the previous lemma. If $\varphi$ is of the form $\psi\land\chi$, then by the recursive definition of the satisfaction relation (cf. Definition A.4) and the induction hypothesis, we have
$\prod_{I}\mathfrak{A}_{i}/\mathcal{U}\vDash\varphi[f_{1}/\mathcal{U},\dots,f_{n}/\mathcal{U}]$
$\Longleftrightarrow\quad\prod_{I}\mathfrak{A}_{i}/\mathcal{U}\vDash\psi[f_{1}/\mathcal{U},\dots,f_{n}/\mathcal{U}]\quad\text{and}\quad\prod_{I}\mathfrak{A}_{i}/\mathcal{U}\vDash\chi[f_{1}/\mathcal{U},\dots,f_{n}/\mathcal{U}],$
$\Longleftrightarrow\quad\{i:\mathfrak{A}_{i}\vDash\psi[f_{1}(i),\dots,f_{n}(i)]\}\in\mathcal{U}\quad\text{and}\quad\{i:\mathfrak{A}_{i}\vDash\chi[f_{1}(i),\dots,f_{n}(i)]\}\in\mathcal{U},$
$\Longleftrightarrow\quad\{i:\mathfrak{A}_{i}\vDash\varphi[f_{1}(i),\dots,f_{n}(i)]\}\in\mathcal{U},$
using the fact that $\{i:\mathfrak{A}_{i}\vDash\varphi\}=\{i:\mathfrak{A}_{i}\vDash\psi\}\cap\{i:\mathfrak{A}_{i}\vDash\chi\}$.
If $\varphi$ is of the form $\lnot\psi$, then, using the fact that $\mathcal{U}$ is an ultrafilter,
$\prod_{I}\mathfrak{A}_{i}/\mathcal{U}\vDash\varphi[f_{1}/\mathcal{U},\dots,f_{n}/\mathcal{U}],$
$\Longleftrightarrow\quad\prod_{I}\mathfrak{A}_{i}/\mathcal{U}\not\vDash\psi[f_{1}/\mathcal{U},\dots,f_{n}/\mathcal{U}],$
$\Longleftrightarrow\quad\{i:\mathfrak{A}_{i}\vDash\psi[f_{1}(i),\dots,f_{n}(i)]\}\not\in\mathcal{U},$
$\Longleftrightarrow\quad\{i:\mathfrak{A}_{i}\vDash\lnot\psi[f_{1}(i),\dots,f_{n}(i)]\}\in\mathcal{U},$
$\Longleftrightarrow\quad\{i:\mathfrak{A}_{i}\vDash\varphi[f_{1}(i),\dots,f_{n}(i)]\}\in\mathcal{U}.$
Finally, if $\varphi$ is of the form $\forall y\;\psi(y,x_{1},\dots,x_{n})$, then
$\prod_{I}\mathfrak{A}_{i}/\mathcal{U}\vDash\varphi[f_{1}/\mathcal{U},\dots,f_{n}/\mathcal{U}],$
$\Longleftrightarrow\quad\prod_{I}\mathfrak{A}_{i}/\mathcal{U}\vDash\psi[g/\mathcal{U},f_{1}/\mathcal{U},\dots,f_{n}/\mathcal{U}]\quad\text{for all }g/\mathcal{U}\in\prod_{I}A_{i}/\mathcal{U},$
$\Longleftrightarrow\quad\{i\in I:\mathfrak{A}_{i}\vDash\psi(g(i),f_{1}(i),\dots,f_{n}(i))\}\in\mathcal{U}\quad\text{for all }g\in\prod_{I}A_{i},$
$\Longleftrightarrow\quad\{i\in I:\mathfrak{A}_{i}\vDash\forall y\;\psi(y,f_{1}(i),\dots,f_{n}(i))\}\in\mathcal{U}\quad\text{for all }g\in\prod_{I}A_{i},$
(since $g(i)$ can be chosen to be any member of $A_{i}$ whatsoever)
$\Longleftrightarrow\quad\{i:\mathfrak{A}_{i}\vDash\varphi[f_{1}(i),\dots,f_{n}(i)]\}\in\mathcal{U}.$
$\dashv$
Corollary A.16
[…]
Hence $\mathfrak{A}^{I}/\mathcal{U}\equiv\mathfrak{A}$.
Proof: If $\varphi(_{1},\dots,x_{n})$ is an $\mathcal{L}$-formula, then by Łos’ Theorem,
$\mathfrak{A}^{I}/\mathcal{U}\vDash\varphi[c_{a_{1}}/\mathcal{U},\dots,c_{a_{n}}/\mathcal{U}]\quad\text{if and only if}\quad\{i\in I:\mathfrak{A}\vDash\varphi[a_{1},\dots,a_{n}]\}\in\mathcal{U}.$
But $\{i\in I:\mathfrak{A}\vDash\varphi[a_{1},\dots,a_{n}]\}$ is either $I$ (if $\mathfrak{A}\vDash\varphi[a_{1},\dots,a_{n}]$), or $\varnothing$ (if $\mathfrak{A}\not\vDash\varphi[a_{1},\dots,a_{n}]$).
$\dashv$
Theorem A.17
Let $\Sigma$ be a set of sentences of a first order language. If every finite subset of $\Sigma$ has a model, then $\Sigma$ has a model.
Proof: Let $I:=\mathcal{P}^{<\omega}(\Sigma)-\{\varnothing\}$ be the set of all non-empty finite subsets of $\Sigma$. For each $i\in I$, let $\mathfrak{A}_{i}$ be an $\mathcal{L}$-model so that $\mathfrak{A}_{i}\vDash i$. For $i\in I$, define $I_{i}\subseteq I$ by
$I_{i}:=\{j\in I:i\subseteq j\}.$
Observe that $I_{i}\cap I_{j}=I_{i\cup j}$, so that the family $\{I_{i}:i\in I\}$ has the f.i.p. Let $\mathcal{U}$ be an ultrafilter over $I$ such that $\{I_{i}:i\in I\}\subseteq\mathcal{U}$. We claim that $\prod_{I}\mathfrak{A}_{i}/\mathcal{U}\vDash\Sigma$.
For suppose that $\varphi\in\Sigma$, and let $i_{0}:=\{\varphi\}\in I$. Observe that $I_{i_{0}}=\{j\in I:\varphi\in j\}$. Now $\varphi\in j$ implies $\mathfrak{A}_{j}\vDash\varphi$, since $\mathfrak{A}_{j}\vDash j$. Thus
$\{j\in I:\mathfrak{A}_{j}\vDash\varphi\}\supseteq\{j\in I:\varphi\in j\}=I_{i_{0}}\in\mathcal{U}.$
Hence $\{j\in I:\mathfrak{A}_{j}\vDash\varphi\}\in\mathcal{U}$, so that by Łos’ Theorem $\prod_{I}\mathfrak{A}_{i}/\mathcal{U}\vDash\varphi$. As this is true for any $\varphi\in\Sigma$, it follows that $\prod_{I}\mathfrak{A}_{i}/\mathcal{U}\vDash\Sigma$.
$\dashv$
Appendix B Existence of Good Ultrafilters
Recall the following definition:
Definition B.1
Let $\kappa$ be a cardinal and let $\mathcal{E}\subseteq\mathcal{P}(\kappa)$.
  1. (a)
    An order-reversal is a map $p:\mathcal{P}^{<\omega}(\kappa)\to\mathcal{E}$ such that whenever $s,t\in\mathcal{P}^{<\omega}(\kappa)$ and $s\subseteq t$, then $p(s)\supseteq p(t)$.
  2. (b)
    An anti-additive map is a map $p:\mathcal{P}^{<\omega}(\kappa)\to\mathcal{E}$ such that if $s,t\in\mathcal{P}^{<\omega}(\kappa)$, then $p(s\cup t)=p(s)\cap p(t)$. Clearly every anti-additive map is an order-reversal.
  3. (c)
    An ultrafilter $\mathcal{U}$ over a set $I$ is $\kappa$-good if for every cardinal $\alpha<\kappa$ and every order-reversal $p:\mathcal{P}^{<\omega}(\alpha)\to\mathcal{U}$ there is an anti-additive map $q:\mathcal{P}^{<\omega}(\alpha)\to\mathcal{U}$ such that $q\leq p$, i.e. such that $q(s)\subseteq p(s)$ for all $s\in\mathcal{P}^{<\omega}(\kappa)$.
$\Box$
The following lemma slightly simplifies the verification of the goodness condition for successor cardinals:
Lemma B.2
Let $\kappa$ be a cardinal, and let $\mathcal{U}$ be an ultrafilter over a set $I$. Then $\mathcal{U}$ is $\kappa^{+}$–good if and only if for every order-reversal $p:\mathcal{P}^{<\omega}(\kappa)\to\mathcal{U}$ there is an anti-additive $q:\mathcal{P}^{<\omega}(\alpha)\to\mathcal{U}$ such that $q\leq p$.
Proof: The ()($\Rightarrow$)( ⇒ )-direction is obvious.
For the ()($\Leftarrow$)( ⇐ )-direction, suppose that $\alpha\leq\kappa$ and that $p:\mathcal{P}^{<\omega}(\alpha)\to\mathcal{U}$ is an order-reversal. Note that if $s$ is a finite subset of $\kappa$, then $s\cap\alpha$ is a finite subset of $\alpha$. Thus we can define an order-reversing map
$\bar{p}:\mathcal{P}^{<\omega}(\kappa)\to\mathcal{U}:a\mapsto p(s\cap\alpha).$
By assumption there is an anti-additive $\bar{q}:\mathcal{P}^{<\omega}(\kappa)\to\mathcal{U}$ such that $\bar{q}\leq\bar{p}$.Define $q$ to be the restriction $q:=\bar{q}\restriction\mathcal{P}^{<\omega}(\alpha)$. Then $q:\mathcal{P}^{<\omega}(\alpha)\to\mathcal{U}$ is an clearly anti-additive map. In addition, if $s\in\mathcal{P}^{<\omega}(\alpha)$, then
$q(s)=\bar{q}(s)\subseteq\bar{p}(s)=p(s\cap\alpha)=p(s),$
so $q\leq p$.
$\dashv$
For an infinite cardinal $\kappa$, let $\mathcal{P}^{*}(\kappa)$ be the set of all subsets of $\kappa$ with cardinality $\kappa$:
$\mathcal{P}^{*}(\kappa)=\{X\subseteq\kappa:|X|=\kappa\}.$
If $h:X\to\mathcal{Y}$ is a function whose range is a family of sets, then we say that $h$ is a disjoint function if the sets $h(x),x\in X$ are disjoint.
Lemma B.3
Suppose that $\kappa$ is an infinite cardinal, and that $f:\kappa\to\mathcal{P}^{*}(\kappa)$. Then there is a disjoint function $h:\kappa\to\mathcal{P}^{*}(\kappa)$ such that $h\leq f$.
Proof: Since $|\kappa\kappa|=\kappa$, there is a bijection $\iota:\kappa\to\kappa\times\kappa$. Define $\mathcal{G}$ to be the set of all one-to-one functions $g$ with the following properties:
  1. (i)
    $\text{\rm dom}(g)\in\kappa$.
  2. (ii)
    If $\xi\in\text{\rm dom}(g)$, then $g(\xi)\in f(\pi_{0}\circ\iota(\xi))$, where $\pi_{0}:\kappa\times\kappa\to\kappa:(\zeta,\eta)\mapsto\zeta$.
    Thus if $\xi\in\iota^{-1}[\{\zeta\}\times\kappa]$, then $g(\xi)\in f(\zeta)$.
Let $G\subseteq\mathcal{G}$ be a maximal chain, ordered by inclusion, and define $g^{*}:=\bigcup G$. Then clearly $g^{*}$ is a one-to-one function satisfying (ii), with $\text{\rm dom}(g^{*})\leq\kappa$. We shall show that $\text{\rm dom}(g^{*})=\kappa$. For suppose that $\text{\rm dom}(g^{*})=\alpha<\kappa$. Then $\iota(\alpha)=(\zeta,\eta)$ for some $\zeta,\eta<\kappa$, and $|\text{\rm ran}(g^{*})|=|\alpha|<\kappa$. As $|f(\zeta)|=\kappa$, we may choose $\gamma\in f(\zeta)-\text{\rm ran}(g^{*})$, and define $g^{\prime}\in\mathcal{G}$ by $g^{\prime}=g^{*}\cup\{(\alpha,\gamma)\}$. Then $g^{\prime}$ is a one-to-one function satisfying (ii) with $\text{\rm dom}(g^{\prime})=\alpha+1$, and hence $G\cup\{g^{\prime}\}$ is a chain in $\mathcal{G}$ which extends $G$ — contradicting the maximality of $G$.
Now define $h:\kappa\to\mathcal{P}^{*}(\kappa)$ by
$h(\zeta):=\{g^{*}(\xi):\xi\in\iota^{-1}[\{\zeta\}\times\kappa]\}=\{g^{*}(\xi):\pi_{0}\circ\iota(\xi)=\zeta\}.$
Then clearly $|h(\zeta)|=\kappa$ as $|\iota^{-1}[\{\zeta\}\times\kappa]|=\kappa$. In addition, if $x\in h(\zeta)$, then $x=g^{*}(\xi)\in f(\pi_{0}\circ\iota(\xi))=f(\zeta)$ (by (ii)), and hence $h(\zeta)\subseteq f(\zeta)$.Moreover, the sets $h(\zeta),\zeta<\kappa$ are disjoint, as $g^{*}$ is one-to-one and the sets $\iota^{-1}[\{\zeta\}\times\kappa]$ (for $\zeta<\kappa$) are disjoint.
$\dashv$
The following corollary is merely a restatement of the previous lemma:
Corollary B.4
Suppose that $\kappa$ is an infinite cardinal, and that $\{X_{\alpha}:\alpha<\kappa\}$ is a family of sets, each of cardinality $\kappa$. Then there exist sets $Y_{\alpha},\alpha<\kappa$ such that
  1. (i)
    Each $Y_{\alpha}\subseteq X_{\alpha}$.
  2. (ii)
    Each $|Y_{\alpha}|=\kappa$,
  3. (iii)
    The sets $Y_{\alpha},\alpha<\kappa$ are mutually disjoint.
Proof: Since $|\bigcup_{\alpha<\kappa}X_{\alpha}|=\kappa$, we may assume without loss of generality that each $X_{\alpha}\in\mathcal{P}^{*}(\kappa)$. Define $f:\kappa\to\mathcal{P}^{*}(\kappa):\alpha\mapsto X_{\alpha}$, and apply Lemma B.3 to obtain $h\leq f$. Now define $Y_{\alpha}:=h(\alpha)$.
$\dashv$
Definition B.5
  1. (a)
    A partition $P$ of a set $X$ is a collection of disjoint subsets of $X$ whose union is $X$. These subsets are called the cells of the partition.
  2. (b)
    A partition is $P$ of $X$ is said to be large if $|P|=X$.
  3. (c)
    If $\mathcal{E}\subseteq\mathcal{P}(X)$ and $\Pi$ a family of partitions of $X$, then the pair $(\Pi,\mathcal{E})$ is consistent if whenever $E\in\mathcal{E}$ and $C_{1},\dots,C_{n}$ are cells chosen from distinct partitions $P_{1},\dots,P_{n}\in\Pi$, we have $E\cap\bigcap_{i\leq n}C_{i}\neq\varnothing$.
  4. (d)
    A family $\mathcal{E}\subseteq\mathcal{P}(X)$ is called a $\pi$-system if it is closed under finite intersections.
$\Box$
Lemma B.6
Suppose that $\kappa$ is an infinite cardinal. Let $\mathcal{E}\subseteq\mathcal{P}^{*}(\kappa)$ be a family such that $|\mathcal{E}|\leq\kappa$. Then there exists a family $\Pi$ of large partitions of $\kappa$ such that $|\Pi|=2^{\kappa}$, and such that $(\Pi,\mathcal{E})$ is consistent.
Proof: Suppose that $\{E_{\alpha}:\alpha<\kappa\}$ is an enumeration of $\mathcal{E}$. By Corollary B.4, there are mutually disjoint sets $I_{\alpha},\alpha<\kappa$ such that $I_{\alpha}\subseteq E_{\alpha}$ and such that $|I_{\alpha}|=\kappa$.
Let $\Gamma:=\{(s,r):s\in\mathcal{P}^{<\omega}(\kappa),r:\mathcal{P}(s)\to\kappa\}$, and observe that $|\Gamma|=\kappa$. Since $|I_{\alpha}|=\kappa$, we can enumerate $\Gamma$ along each $I_{\alpha}$, i.e. there is an enumeration $\{(s_{\xi},r_{\xi}):\xi<\kappa\}$ (with repetitions!) of $\Gamma$ such that
$\Gamma=\{(s_{\xi},r_{\xi}):\xi\in I_{\alpha}\}\qquad\text{ for each }\alpha<\kappa.$
For each $J\in\mathcal{P}(\kappa)$, define
$f_{J}:\kappa\to\kappa:\xi\mapsto\left\{\begin{aligned} r_{\xi}(s_{\xi}\cap J)\quad&\text{if }\xi\in\bigcup_{\alpha<\kappa}I_{\alpha},\\ 0\qquad&\text{otherwise}.\end{aligned}\right..$
We now show that for each sequence $\alpha,\gamma_{1},\dots,\gamma_{n}\in\kappa$, and each sequence $J_{1},\dots,J_{n}$ of distinct subsets of $\kappa$
$\text{there exists $\xi\in I_{\alpha}$ such that}\quad f_{J_{i}}(\xi)=\gamma_{i},\qquad i=1,\dots,n.$
For since $J_{1},\dots,J_{n}$ are distinct subsets of $\kappa$, the symmetric differences $J_{i}\Delta J_{j}$ are non-empty, and thus we may choose an element $x_{ij}\in J_{i}\Delta J_{j}$ for each pair $i,j$ with $1\leq 1<j\leq n$. Define $s:=\{x_{ij};1\leq i<j\leq n\}$, and observe that the sets $s\cap J_{i}$ are distinct (for $i=1,\dots,n$). Now let $r:\mathcal{P}(s)\to\kappa$ be any map such that $r(s\cap J_{i})=\gamma_{i}$. Then $(s,r)\in\Gamma$, so there is $\xi\in I_{\alpha}$ so that $(s_{\xi},r_{\xi})=(s,r)$ Then $f_{J_{i}}(\xi)=\gamma_{i}$ for all $i=1,\dots,n$, as required.
Observe that each $f_{J}:\kappa\to\kappa$ is surjective, as can be seen from ()($\star$)( ⋆ ) with $n=1$. Thus if we define $P_{J}$ to be the partition of $\kappa$ given by
$P_{J}:=\{f_{J}^{-1}\{\gamma\}:\gamma<\kappa\},$
then $|P_{J}|=\kappa$, i.e $P_{J}$ is large.
Let $\Pi:=\{P_{J}:J\subseteq\kappa\}$. To show that $(\Pi,\mathcal{E})$ is consistent, suppose that $E\in\mathcal{E}$ and that $C_{1},\dots,C_{n}$ belong to $P_{J_{1}},\dots,P_{J_{n}}$, where the $J_{i}$ are distinct subsets of $\kappa$. Then there exists $\alpha<\kappa$ such that $E=E_{\alpha}$, and there exist $\gamma_{1},\dots,\gamma_{n}<\kappa$ so that $C_{i}=f_{J_{i}}^{-1}\{\gamma_{i}\}$ for $i=1,\dots,n$. But then by ()($\star$)( ⋆ ) there exists $\xi\in I_{\alpha}$ such that $f_{J_{i}}(\xi)=\gamma_{i}$, from which it follows that
$\xi\in I_{\alpha}\cap\bigcap_{i\leq n}C_{i}.$
Since $I_{\alpha}\subseteq E_{\alpha}$, it follows that $E_{\alpha}\cap\bigcap_{i\leq n}C_{i}\neq\varnothing$. Hence $(\Pi,\mathcal{E})$ is consistent.
In particular, if $J_{1},J_{2}$ are distinct subsets of $\kappa$, then $C_{1}\cap C_{2}\neq\varnothing$ for any $C_{1}\in P_{J_{1}},C_{2}\in P_{J_{2}}$. It follows that $P_{J_{1}},P_{J_{2}}$ have no cells in common, and thus that $P_{J_{1}}\neq P_{J_{2}}$, i.e. all the $P_{J}$ are distinct. Hence $|\Pi|=|\mathcal{P}(\kappa)|=2^{\kappa}$.
$\dashv$
Lemma B.7
Suppose that $\Pi$ is a family of partitions of a cardinal $\kappa$ and that $\mathcal{E}\subseteq\mathcal{P}(\kappa)$ is a $\pi$-system. Suppose further that $(\Pi,\mathcal{E})$ is consistent. Let $J\subseteq\kappa$. Then either $(\Pi,\mathcal{E}\cap\{J\})$ is consistent, or else there is a cofinite $\Pi^{\prime}\subseteq\Pi$ so that $(\Pi^{\prime},\mathcal{E}\cap\{\kappa-J\})$ is consistent.
Proof: If $(\Pi,\mathcal{E}\cap\{J\})$ is not consistent, there is a set $E\in\mathcal{E}$, and distinct partitions $P_{1},\dots,P_{n}\in\Pi$ and corresponding cells $C_{i}\in P_{i}$ such that
$J\cap E\cap\bigcap_{i\leq n}C_{i}=\varnothing.$
It follows that $E\cap\bigcap_{i\leq n}C_{i}\subseteq\kappa-J$. Let $\Pi^{\prime}:=\Pi-\{P_{1},\dots,P_{n}\}$. To prove that $(\Pi^{\prime},\mathcal{E}\cup\{\kappa-J\})$ is consistent, take $E^{\prime}\in\mathcal{E}$, distinct partitions $P_{1}^{\prime},\dots,P_{m}^{\prime}\in\Pi^{\prime}$, and cells $C_{1}^{\prime},\dots,C_{m}^{\prime}$ in $P_{1}^{\prime},\dots,P_{m}^{\prime}$. We need only show $(\kappa-J)\cap E^{\prime}\cap\bigcap_{j=1}C_{j}^{\prime}\neq\varnothing$. Now as $(\Pi,\mathcal{E})$ is consistent and $\mathcal{E}$ is closed under finite intersections, we have
$E\cap E^{\prime}\cap\bigcap_{i\leq n}C_{i}\cap\bigcap_{j\leq m}C_{j}^{\prime}\neq\varnothing.$
As $E\cap\bigcap_{i\leq n}C_{i}\subseteq\kappa-J$, we immediately see that also $(\kappa-J)\cap E^{\prime}\cap\bigcap_{j\leq m}C_{j}^{\prime},\neq\varnothing$.
$\dashv$
Lemma B.8
Suppose that $\Pi$ is a family of large partitions of a cardinal $\kappa$ and that $\mathcal{E}\subseteq\mathcal{P}(\kappa)$ is a $\pi$-system. Suppose further that $(\Pi,\mathcal{E})$ is consistent. Let $p:\mathcal{P}^{<\omega}(\kappa)\to\mathcal{E}$ be an order-reversal, and let $P\in\Pi$. Then there is a $\pi$-system $\mathcal{E}^{\prime}\supseteq\mathcal{E}$ and an anti-additive map $q:\mathcal{P}^{<\omega}(\kappa)\to\mathcal{E}^{\prime}$ such that $q\leq p$ and $(\Pi-\{P\},\mathcal{E}^{\prime})$ is consistent.
In addition, we can take $\mathcal{E}^{\prime}$ to be the $\pi$-system generated by $\mathcal{E}\cup\text{\rm ran}(q)$.
Proof: Let $\{C_{\alpha}:\alpha<\kappa\}$ be an enumeration of the cells of $P$, without repetition, and let $\{t_{\alpha}:\alpha<\kappa\}$ be an enumeration of $\mathcal{P}^{<\omega}(\kappa)$. For each $\alpha<\kappa$ define a map $q_{\alpha}:\mathcal{P}^{<\omega}(\kappa)\to\mathcal{P}(\kappa)$ by
$q_{\alpha}(s):=\left\{\begin{aligned} p(t_{\alpha})\cap C_{\alpha}\quad&\text{if }s\subseteq t_{\alpha},\\ \varnothing\qquad&\text{else.}\end{aligned}\right.$
Since $p$ is an order-reversal, $q_{\alpha}(s)\subseteq p(t_{\alpha})\subseteq p(s)$. Furthermore, since $(\Pi,\mathcal{E})$ is consistent, $p(t_{\alpha})\in\mathcal{E}$ and $P\in\Pi$, we always have $p(t_{\alpha})\cap C_{\alpha}\neq\varnothing$. Hence if $s\subseteq t_{\alpha}$, then $q_{\alpha}(s)\neq\varnothing$.
Next observe that each $q_{\alpha}$ is anti-additive: Indeed, if both $s,s^{\prime}\subseteq t_{\alpha}$, then clearly $q_{\alpha}(s\cup s^{\prime})=p(t_{\alpha})\cap C_{\alpha}=q_{\alpha}(s)\cap q_{\alpha}(s^{\prime})$. On the other hand, if one of $s,s^{\prime}$ is not contained in $t_{\alpha}$, then $q_{\alpha}(s\cap s^{\prime})=\varnothing=q_{\alpha}(s)\cap q_{\alpha}(s^{\prime})$.
Define a function $q$ on $\mathcal{P}^{<\omega}(\kappa)$ by
$q(s):=\bigcup_{\alpha<\kappa}q_{\alpha}(s).$
Then since each $q_{\alpha}(s)\subseteq p(s)$ we have $q\leq p$. Further note that if $s,s^{\prime}\in\mathcal{P}^{<\omega}(\kappa)$ (not necessarily distinct), and $\alpha\neq\beta$, then $q_{\alpha}(s)\cap q_{\beta}(s^{\prime})=\varnothing$, since $C_{\alpha}\cap C_{\beta}=\varnothing$. In particular, $q(s)$ is a disjoint union of subsets of the cells $C_{\alpha}$, and hence $q$ is also anti-additive:
$q(s)\cap q(s^{\prime})$ $=\bigcup_{\alpha<\kappa}q_{\alpha}(s)\cap\bigcup_{\beta<\kappa}q_{\beta}(s^{\prime}),$
$=\bigcup_{\alpha<\kappa}\bigcup_{\beta<\kappa}\big{(}q_{\alpha}(s)\cap q_{\beta}(s^{\prime})\big{)},$
$=\bigcup_{\alpha<\kappa}\big{(}q_{\alpha}(s)\cap q_{\alpha}(s^{\prime})\big{)},$
$=\bigcup_{\alpha<\kappa}q_{\alpha}(s\cup s^{\prime}),$
$=q(s\cup s^{\prime}).$
Hence $\text{\rm ran}(q)$ is a $\pi$-system. Define $\mathcal{E}^{\prime}$ to be the $\pi$-system generated by $\mathcal{E}\cup\text{\rm ran}(q)$:
:=$\mathcal{E}^{\prime}:=\{E\cap q(s):E\in\mathcal{E},s\in\mathcal{P}^{<\omega}(\kappa)\},$
and note that $q:\mathcal{P}^{<\omega}(\kappa)\to\mathcal{E}^{\prime}$. We claim that $(\Pi-\{P\},\mathcal{E}^{\prime})$ is consistent. For suppose that, for $i=1,\dots,n$ we have $D_{i}\in P_{i}\in\Pi-\{P\}$, and that $E\in\mathcal{E},s\in\mathcal{P}^{<\omega}(\kappa)$. Then $s=t_{\alpha}$ for some $\alpha<\kappa$, and hence $q_{\alpha}(s)=p(t_{\alpha})\cap C_{\alpha}=p(s)\cap C_{\alpha}$. Now since $p:\mathcal{P}^{<\omega}(\kappa)\to\mathcal{E}$ and $\mathcal{E}$ is a $\pi$-system, we have $p(s)\cap E\in\mathcal{E}$. Since $(\Pi,\mathcal{E})$ is consistent,
$p(s)\cap E\cap C_{\alpha}\cap\bigcap_{i\leq n}D_{i}\neq\varnothing,$
and hence
$\big{(}E\cap q(s)\big{)}\cap\bigcap_{i\leq n}D_{i}\supseteq\big{(}E\cap q_{\alpha}(s)\big{)}\cap\bigcap_{i\leq n}D_{i}=p(s)\cap E\cap C_{\alpha}\cap\bigcap_{i\leq n}D_{i}\varnothing.$
$\dashv$
We are now ready to prove the existence of good ultrafilters under certain conditions:
Theorem B.9
Suppose that $\kappa$ is an infinite cardinal. Then there exists a $\kappa^{+}$-good countably incomplete ultrafilter over $\kappa$.
Proof: Start with a sequence $I_{n}\downarrow\varnothing$ of subsets of $\kappa$, each of cardinality $\kappa$. (This exists since $|\omega\times\kappa|=\kappa$.) Let $\mathcal{F}_{0}$ be the filter over $\kappa$ generated by the sets $I_{n}$. Then any filter $\mathcal{F}$ which extends $\mathcal{F}_{0}$ will be countably incomplete.
By Lemma B.6 there is a family $\Pi_{0}$ of large partitions of $\kappa$ such that $|\Pi_{0}|=2^{\kappa}$, and such that $(\Pi_{0},\mathcal{F}_{0})$ is consistent. We now use transfinite induction to define, for ordinals $\xi<2^{\kappa}$, a sequence $\Pi_{\xi}$ of partitions of $\kappa$, and a sequence $\mathcal{F}_{\xi}$ of filters over $\kappa$ such that
  1. (i)
    If $\eta\leq\xi<2^{\kappa}$, then $\Pi_{\eta}\supseteq\Pi_{\xi}$ and $\mathcal{F}_{\eta}\subseteq\mathcal{F}_{\xi}$.
  2. (ii)
    $|\Pi_{\xi}|=2^{\kappa}$.
  3. (iii)
    $|\Pi_{\xi+1}-\Pi_{\xi}|<\omega$.
  4. (iv)
    $\Pi_{\lambda}=\bigcap_{\xi<\lambda}\Pi_{\xi}$ if $\lambda<2^{\kappa}$ is a limit ordinal.
  5. (v)
    Each pair $(\Pi_{\xi},\mathcal{F}_{\xi})$ is consistent.
Suppose that $\{p_{\xi}:\xi<2^{\kappa}\}$ is an enumeration of all order-reversing maps $\mathcal{P}^{<\omega}(\kappa)\to\mathcal{P}(\kappa)$. Similarly, let $\{J_{\xi}:\xi<2^{\kappa}\}$ enumerate $\mathcal{P}(\kappa)$. Assume now that we have defined $(\Pi_{\eta},\mathcal{F}_{\eta})$ for all $\eta<\xi$, so that (i)-(iv) are satisfied.
There are three cases in the construction of the $\kappa^{+}$-good ultrafilter $\mathcal{U}$ by transfinite induction. Cases 1 and 2 deal with odd and even successor ordinals. Case 1 ensures that the filter we construct is an ultrafilter, whereas Case 2 ensures that it is good. Case 3, which is the simplest, deals with the step at limit ordinals.
Case 1: $\xi$ is a successor ordinal of the form $\lambda+2n+1$, where $\lambda$ is a limit ordinal and $n<\omega$.
Let $J_{\eta}$ be the first subset of $\kappa$ in the enumeration of $\mathcal{P}(\kappa)$ which is not in $\mathcal{F}_{\xi-1}$. By Lemma B.7, we can find a partition $\Pi_{\xi}\subseteq\Pi_{\xi-1}$ such that $|\Pi_{\xi-1}-\Pi_{\xi}|<\omega$ (so that $|\Pi_{\xi}|=2^{\kappa}$ also), such that either $(\Pi_{\xi},\mathcal{F}_{\xi-1}\cup\{J_{\eta}\})$ is consistent, or $(\Pi_{\xi},\mathcal{F}_{\xi-1}\cup\{\kappa-J_{\eta}\})$ is consistent. In the former case, define $\mathcal{F}_{\xi}$ to be the filter generated by $\mathcal{F}_{\xi-1}\cup\{J_{\eta}\}$, and in the latter, the filter generated by $\mathcal{F}_{\xi-1}\cup\{\kappa-J_{\eta}\}$. Thus $(\Pi_{\xi},\mathcal{F}_{\xi})$ is consistent.
Case 2: $\xi$ is a successor ordinal of the form $\lambda+2n+2$, where $\lambda$ is a limit ordinal.
In that case let $p_{\eta}$ be the first function $\mathcal{P}^{<\omega}(\kappa)\to\mathcal{F}_{\xi-1}$ in the enumeration of order-reversing maps $\mathcal{P}^{<\omega}(\kappa)\to\mathcal{P}(\kappa)$ that has not already been dealt with. Pick $P\in\Pi_{\xi-1}$. By Lemma B.8, there is an anti-additive $q:\mathcal{P}^{<\omega}(\kappa)\to\mathcal{F}_{\xi}$ such that $q\leq p_{\eta}$, and such that $(\Pi_{\xi},\mathcal{F}_{\xi})$ is consistent, where $\Pi_{\xi}:=\Pi_{\xi-1}-\{P\}$ and $\mathcal{F}_{\xi}$ is the filter generated by $\mathcal{F}_{\xi-1}\cup\text{\rm ran}(q)$.
Case 3: $\xi$ is a limit ordinal.
In that case, define $\Pi_{\xi}:=\bigcap_{\eta<\xi}\Pi_{\eta}$ and $\mathcal{F}_{\xi}:=\bigcup_{\eta<\xi}\mathcal{F}_{\eta}$. Since at each stage $\eta<\xi$ we remove only finitely many partitions, at most $|\xi\cdot\omega|<2^{\kappa}$–many partitions have been removed from $\Pi_{0}$ to form $\Pi_{\xi}$, and hence $|\Pi_{\xi}|=2^{\kappa}$. It is easy to see that $(\Pi_{\xi},\mathcal{F}_{\xi})$ is consistent: For suppose that $F\in\mathcal{F}_{\xi}$ and $C_{1},\dots,C_{m}$ are cells from distinct partitions $P_{1},\dots,P_{m}\in\Pi_{\xi}$. Then there is $\eta<\xi$ such that $F\in\mathcal{F}_{\eta}$, and moreover $C_{1},\dots,C_{m}$ are cells from distinct partitions in $\Pi_{\eta}$, as $\Pi_{\xi}\subseteq\Pi_{\eta}$. Since $(\Pi_{\eta},\mathcal{F}_{\eta})$ is consistent, $F\cap\bigcap_{i\leq m}C_{i}\neq\varnothing$.
This completes the transfinite induction.
Now define $\mathcal{U}:=\bigcup_{\xi<2^{\kappa}}\mathcal{F}_{\xi}$. By Case 1, for every $J\subseteq\kappa$, either $J\in\mathcal{U}$ or $\kappa-J\in\mathcal{U}$. Hence $\mathcal{U}$ is an ultrafilter. Since $\mathcal{U}\supseteq\mathcal{F}_{0}$ it is a countably incomplete ultrafilter.
Suppose now that $p:\mathcal{P}^{<\omega}(\kappa)\to\mathcal{U}$ is an order-reversal. Observe [^1] Recall König’s Theorem: If $\alpha_{i}<\beta_{i}$ are cardinals, then $\sum_{i}\alpha_{i}<\prod_{i}\beta_{i}$. Now if $\alpha_{i}<2^{\kappa}$, and $\beta_{i}=2^{\kappa}$ for $i<\kappa$, then $\sum_{i<\kappa}\alpha_{i}<\prod_{i<\kappa}2^{\kappa}=2^{\kappa}$, i.e. $\sum_{i<\kappa}\alpha_{i}<2^{\kappa}$. Hence $\text{cf}(2^{\kappa})>\kappa$. that since $\text{cf}(2^{\kappa})>\kappa$ and $|\mathcal{P}^{<\omega}(\kappa)|=\kappa$, there is a least $\eta<\kappa$ so that $\text{\rm ran}(p)\subseteq\mathcal{F}_{\eta}$. Thus $p$ will be dealt with at some stage $\xi\geq\eta$ by Case 2, which guarantees the existence of an anti-additive $q\leq p$ which maps into $\mathcal{F}_{\xi}\subseteq\mathcal{U}$. It follows that $\mathcal{U}$ is $\kappa^{+}$-good.
$\dashv$